A Selected Bibliography of Publications by, and about, George Gamow

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

28 March 2025 Version 1.104

Title word cross-reference

$1.95 [Smi61a]. $16.95 [Hob02]. $2.50 [Ano55a]. $2.75 [Joh54a]. $24.95 [Hob02]. $35.00 [Dys02]. $5.75 [Sit64b]. α [CG30, Gam29d, Gam30b, Gam32a, Gam33b, MP31, Rut27]. αβγ [AWCT09, Tur08]. β [Gam33e, Gam34a, GT36, Gam37b, GT37, HS19]. c [Gam39c]. G [Gam39c]. γ [BG36, Gam33e, Gam75a, MP31]. h [Gam39c]. p [Gam32a].

-and [Gam32a]. -Disintegration [Gam33e, GT36]. -Excitation [Gam33e]. -Feinstruktur [MP31]. -levels [Gam32a]. -Particles [CG30, Gam33b]. -Ray [BG36]. -Rays [Gam30b, Gam75a, Rut27]. -Spektrum [MP31]. -Transformation [GT37]. -Transformations [Gam29d]. -Zerfalls [Gam34a, Gam37b].

0 [Dys02]. 0-521-63009-6 [Per03]. 0-521-63992-1 [Per03]. 0-7382-0532-X [Dys02].

1 [Gam37b, UM86a]. 19 [Ano69, Öpi69]. 1911 [Meh75]. 1930/41 [Fer68, Fer71]. 1933 [CCJ^{+}34, Gam37b]. 1934 [Gam34a]. 1941 [TGF41]. 1960s [Mla98]. 1968 [Huf09, Wei68]. 1982 [MR86]. 1986 [Dys87]. 1993 [Boy93]. 1994 [BCY95]. 1997 [Ano98].

2 [Cas12a, UM86b]. 20 [Gam34a]. 2001 [Ano02]. 2005 [BL09]. 2009 [CBKZ^{+}09]. 2010 [KLR13]. 20th [Sha07]. 22 [CCJ^{+}34].

4th [CBKZ^{+}09].

60th [MF69]. 6th [Rya06].

70th [Ano55a]. 75th [Gam60].

80th [MW88].

9.80 [Uns60]. 90th [Fre94a]. 978 [Cas12a]. 978-0-670-02276-2 [Cas12a]. 9th [CBKZ^{+}09].

år [Gam66g, Gam68e]. Abbildungen [Gam38c]. Abich [Kuh67]. Absorption [Gam47f]. Abundance [Alp48, AH48, LW46, Wat46, TTL07]. Abundances [DW48, Tri10]. Academy [Gor90]. Accelerators [Bet97]. account [Alp73]. Acid [Bre57, Gam54g, GY55, GY56, Gam54f]. Acids [Gam54d, Gam55e, GRY56b, GRY56a, Gam57a]. act [Stu13]. Adult [Cas12a]. Advances [LT56]. Adventures [GY67, Gam94, Sus69]. After [Ano54, Wat02, BKST^{+}07, CBKZ^{+}09, Coc46]. against [Hoy90]. Age [Stu18]. Albert [Ano05, Igg66, Kle05, Wei13]. Albrecht [BL09]. alkaline [HS39]. allowed [HS19]. Alls [Uns60]. Alpha [Lip86, Stu86, Stu97, Rig06, Rig07]. alpha-Centaury [Rig06, Rig07]. Alpha-Decay [Lip86, Stu86]. Alphabet [Gam53c]. Alpher [AH73, Alp12, AM23, Har07, Mar08, Rig06, Rig07]. Amateur [Har01]. amazing [Tur08]. America [Che94a]. American [Ano55b]. Ames [Ske54]. Analysis [BC05]. Analytic [Cri72]. angeregten [DG31]. années [Gam68f, Gam01b]. Anniversary [Fre94a, Che94a, Fre94c]. Annual [GF42, TGF41]. Anomale [PG27]. Anomalous [PG27]. Anregung [Gam31b]. Antonovich [Sco07]. Any [Gam49a]. Apology [Pra93]. Appendix [HPA97a]. apple [Gam62e]. applications [CMT17, Wei72a]. Applied [BP64, BP81]. Appreciation [Sha72, Har01]. Approach [Kra05, GY58]. April [HPA97b, Wei68]. Arbeit [Rac35]. Arbeiten [Hou30]. Archaeology [Fre10]. art [Che94b]. article [Oku02]. Artificial [CG30, Gam34b, Gam33f, Rog62]. Asher [Har07]. aspects [Meh75]. assessment [Hoy90]. astounding [Haw11]. Astrobiology [Rya06]. astronomers [HTB^{+}07]. Astronomical [Ano98, Kra17]. Astronomy [Gam6x, Rya06, Ano55b, FF91, Pen72]. Astrophysics [Bet97, CO07, CBKZ^{+}09, Gin94, Rya06, Sal96, BCY95, BKST^{+}07, CBKZ^{+}09, Ano95].

Atom [Gam44b, Gam47a, Gam61a, Gam63d, Meg61, Smi61a, Smi61b, Gam46c, Gam63a, Gam63b, Rut27, Gam63b, Ano44, GHJ47, Pom44]. Atome [Gam51a, Gam51g]. Atomen [Gam47a, Gam46c]. Atomic [FR13, Gam29c, Gam30a, Gam31a, Gam32f, Gam32c, Gam34j, Gam37a, Gam46a, Gam47b, GHJ47, GC49, Gam4xb, Gam11a, Har32, Pom44, RAC^{+}29, CCJ^{+}34, Gam28b, Gam29e, Gam29a, GH32, Gam32d, Gam32h, Gam33i, Gam34d, Gam34c, Gam34e, Gam35a, GBK48, Gam52b, Gam93b, Hou30, Rac35, vdBS12, Gam35e, Smi61b]. atomiques [CCJ^{+}34, Gam35e]. Atomkernen [Gam34c, Rac35]. Atomkernes [Pau32, Gam28b, Gam29a]. Atomkerns [GH32]. Atomnoe [Gam30a]. Atomnogo [Gam32c]. atomo [Gam63a]. Atoms [Gam50c, GHJ47]. Atomskerns [Hou30]. Atomzertrümmerung [Gam29e]. atoomkern [vdBS12]. attempt [Gam32e]. Aug [Ano69]. August [CBKZ^{+}09, Huf09, Öpi69]. auspices [CCJ^{+}34]. Autobiographical [Pra93]. Autobiography [Gam70, AH71, Ric71, Stu71, Gam66d]. Autoelectric [Opp28]. autres [Kle05, Ano05]. aux [Gam62e].

B [Kuh67, Wil71]. Background [Alp12, Che94b, Nov07, Wil79]. Baitsell [Det55]. Bang [BK21, Cas12a, Kra96b, AH72b, AH90, AH01, Cas12b, Che95, Kra96c, Kra14, Lal14, Seg11, Sin04]. Barbara [Kuh67, Ano50b, GG76]. Basic [Gam40a, GB68]. Basis [GR31, Woe67]. Bau [GH32, Pau32]. be [Ray04, Ray05]. before [Rig95, Tri10]. Being [Bad71]. Beitragen [Kuh67]. Bemerkung [Gam29b, Rac35]. Benedict [Ano47]. Benjamin [Igg66]. Bericht [Gam38c]. Berlin [Gam51b, KLR13]. Bestrahlung [HS39]. Beta [HS19]. Bethe [BL09, Mar08, AH73, Rig06, Rig07]. Between [Stu18, AG68, Fre94b, Gam54f, Gam54g]. Beyond [Inf48, Rya06]. Big [Cas12a, Kra96b, Kra96c, AH72b, AH90, AH01, BK21, Cas12b, Che95, Kra14, Lal14, Seg11, Sin04]. big-bang [AH90]. Biggest [OM18, Wei13]. Biografía [Gam42b, Gam14]. biographical [HTB^{+}07]. Biographie [Gam65a]. Biography [Gam41b, Gam42c, Gam42d, Gam48a, Gam59b, Gam61b, Gam62a, Gam65a, Ano50a, Gam14, Meg62, Van62, C.48, Dix61, Gla49, K.62, Mat42]. Biological [LT56]. Biology [GY67, Gam49f, Sus69]. Birth [Ano40, Ano47, Fre94a, Gam40b, GS42, Gam45a, Gam47d, Gam49b, Gam52b, Gam05, Hun49, M.40a, Mul41, Uns60, AH96, Che94a, Fre94c]. Birthday [Gam60, MF69, MW88]. blackbody [AH90]. Blood [Ske54]. Blunder [OM18, Wei13]. Bohr [Gam60, Gam66c, Kuh67, Gam60, Gam63d]. Bohrs [Kuh67]. Bondi [Gam52d]. Book [AH71, Ano40, Ano44, Ano98, Bar53, Cas12a, Dan65, Dwi66, Dys87, Dys02, Fea62, Fre61, Gam51a, Gam51b, Gam52a, Gre00, Hob02, Joh54a, Jud01, Las62, M.40b, Meg61, Meg62, Pau32, Per03, Sit64b, Sit64a, Smi61a, Stu71, Van62, Wil71, Ano55b, Ano50a, Ano02, Boy93, Uns60]. Books [Azi67, Ber68, Gam42a, Hoo93]. Born [Gam41c]. Boulder [Wei68]. Bound [Gam41a, Gam51b]. boundary [GIL02a]. bright [BBC^{+}07]. brightest [BBC^{+}07]. Broglie [Gam54h, Gam26]. Brussels [CCJ^{+}34]. Bruxelles [CCJ^{+}34, Far01]. Buchbesprechung [Uns60]. bulk [vN96]. Burkholder [Det55]. bursts [BBC^{+}07]. butsurigaku [Gam42].

C [Alp12]. C. [GH45]. ca [Gam55b]. Calculability [Cer05]. calculated [Che94b]. Calculus [Lin00]. Called [Gam63f, Gam64d, Gam67h, Sit64b, Sit64c, Sit64a]. Cambridge [Ano02, Boy93, Dys02, Hob02, Joh54a, Per03, Coc46, Nye02]. Campbell [GHJ47]. Capture [Alp48, SCG08]. Carbon [Hoy54]. Carnegie [HPA97b]. Carpenter [Det55]. Case [Gam67a, Mis08]. cassetête [GS67]. celebrating [MW88]. Celebration [AH96]. Cell [Gam55c]. Centaury [Rig06, Rig07]. Century [Dys02, Sha07]. Cepheid [GL50]. chain [Kav72]. Chains [GM54]. Change [Gam67b]. Changed [Gam66c, Kuh67]. Characteristics [AG68]. Charge [Gam67j]. charged [Lon72]. Charles [Det55]. che [Gam01a, Gam10]. Chemical [ABG48, FB12a, FN12, Fre14, Gam35c, Gam42e, Wat48, Gam66b, Tri10]. Chicago [Wil71]. Christian [Ano47]. Christmas [Gam6x]. Cinétique [Gam36a, Gam36b]. Clarification [The01]. Classic [Gam56b, Sta99]. Classics [Pra93]. Cleveland [Fea62, Fre61]. Cline [Kuh67]. cloth [Hob02]. COBE [MFS93]. Code [Cer05, Hay98, Ric97, Dem07, Nan04, Woe67]. Codes [Bre57]. Coiled [Gam55f]. collaboration [AM23]. Collapse [GS41, GS46]. Colloquium [UM86a, UM86b]. Colorado [Wei68]. Combinatorial [BP64, BP81, Gam64a, Gam68a, Gam81]. Comment [Lip86]. Common [Ano99]. community [Hug93]. Composition [GY55, GY56, SST72]. computers [URR86b]. Computing [Hay98]. Conant [Wil71]. Conceptual [Kuh67]. Concerning [Gam42e, HS39]. Conclusions [Gam50c]. Conditions [AFH53, MF69]. Conference [CBKZ^{+}09, GF42, KLR13, TGF39, TGF41, PD00, BKST^{+}07, Far01]. Conferences [Anoxx, Meh75]. confirmed [Pol72]. confirmée [Pol72]. Conformal [HN72]. Confronts [Nim11]. Congress [Gam34e]. Connecting [Mis08]. conocimiento [Lal14]. Consciousness [Igg66]. conseil [CCJ^{+}34]. consideration [Sab96]. Considérations [Hei34, Hei34]. Constant [Bek86, Mis08, Alp73, Dir72, Tel72, Wei13]. Constants [BF86, Cer05, GIL28, Gam68c, GIL02b, Oku02, Alp73, GIL02a, PG27, Tel72]. Constitution [Gam30c, Gam31a, Gam32c, Gam34g, Har32]. Contracting [Gam44a]. Contraction [Gam61f, Gam67i]. Contractive [Gam43a]. Controversy [Kra96a, Hug93]. Copernicus [Gam42i, Gam42a]. copier [Sch12b]. Core [GT38b]. Correlation [GY55, GY56, Gam67a]. Correspondence [Fre94b]. Cosmic [AGH67, BBC^{+}07, Fre14, Gam33a, Gam46a, Gam47b, Gam11a, AH90, Gam33j, Nov07, SST72, Wil79, Alp12]. Cosmogonies [Kra02]. Cosmogony [Gam49e]. Cosmological [AG68, BF86, Gam52a, Gam52d, Gam66b, Mis08, AM23, Pol72, Wei13]. Cosmologies [Kra02]. cosmologique [Pol72]. Cosmology [Alp73, BF86, BBBM90, Cas12b, Gam52a, Gam54c, Gam55a, Gam67d, Gam67c, Gam90, Kra96a, Kra05, Kra18, Les90, Pee71, Pee93, Pen72, Rei72a, Rei72b, Rin11, Rya06, Wei08, AH72b, AH90, BKST^{+}07, CBKZ^{+}09, HN72, JLA^{+}04, MR86, Rin09, Seg11, Tur08, Wei72a, Cas12a, Dys87, Gam52d]. Cosmonumerology [Kra91b, Kra91a]. cosmos [Gam80, vdBS12]. Course [MR86]. Cover [Nug54]. Creation [Gam52c, Gam58a, Gam61c, Gam04a, Uns60, Stu13, Gam50a]. Criteria [FB12a]. Critique [Ano05]. Crompton [Nug54]. crossings [GIL02a]. cryptographic [GY58]. Culture [Ano47]. current [Gam34a, Gam37b]. Currents [Opp28]. Curve [Gam30c].

D [Det55, Jud01, Oku02, GT56, GT58]. Danish [Gam42g, Gam68e]. dargest [GT58]. Darnton [Ano47]. Davidson [Gam69b]. day.Genes [Jud01]. de-polariz-ed [Luk70]. Death [Ano40, Ano47, Gam40b, GS42, Gam45a, Gam47d, Gam49b, Gam52b, Gam05, Hun49, M.40a, Mul41]. Decay [Gam32f, Lip86, Stu86, Stu97, Gam29b, Gam31b, Gam32e, Gam34a, Gam37b, Gam34d]. Defect [Gam30c]. Defining [Mla98]. degree [GLI26]. del [GS42, Gam63a]. Delbrück [Cas12a, Cas12b, Seg11]. della [Gam01a]. delta [Gam32i]. d’énergie [Gam33d]. Deoxyribonucleic [Gam54g, Gam54f]. destruction [Gam31b]. Determination [GR31]. Development [Gam32f, Kra96a, Kuh67, Gam32d, Gam33i, Gam34d, Gam93b, Meh75]. Dialogue [Kra91b, Kra91a]. diameters [GR33]. diamètres [GR33]. Did [Rub97, Wei13]. died [Ano69, Öpi69]. Different [Gam41d]. Diffusion [SST72]. Dirac [Gam33k, Kra91b, Kra91a]. Dirac/Gamow [Kra91a]. Discovery [BC05, Gra64, GA71, GN00, Nov07]. Discussion [Gam29c, GNF^{+}97, RAC^{+}29, DGS^{+}56]. discussions [CCJ^{+}34]. Disintegration [CG30, Gam28a, Gam32g, Gam33e, GT36, GC28, GC29]. Dispersion [PG27]. Dispersionskonstanten [PG27]. DM [Gam51b, Uns60]. Do [Gam42j, Alp73]. Doctrines [Gam42j]. documentary [GA71]. Does [Gam67b]. Donald [Det55]. done [Sab96]. Double [Wat02]. doublets [PG27]. dreamers [Poh62]. dreams [Haw11]. drei [GT56, GT58]. Drømmeland [Gam42g]. Drop [Ano94, Stu94, Stu97]. drops [CMT17]. Dualism [Gam40c]. Dubletts [PG27]. durch [Gam31b, Gam80]. Dutch [Hun49, vdBS12]. Dynamics [BGK50, BGK51].

Early [AH90, Bet97, SCG08]. Earth [Bin58, Dan65, Dix61, Dwi66, Fie59, Gla49, Mat42, PG66, HS39, Ano50a, Gam41b, Gam42c, Gam42d, Gam42b, Gam48a, Gam54b, Gam58b, Gam59b, Gam59d, Gam63f, Gam65c, Gam69a, C.48, K.62, Mat42, Ske54]. ébranlèrent [Gam68f, Gam01b]. ed [Luk70]. Eddington [Bek86]. Editor [GT37, GT38a, GT39c, Gam67f, Gam67g, Mar08]. Edward [Dys02, MF69, MW88]. effectifs [GR33]. effective [GR33]. Eighth [GF42]. Einelektronige [Gam51a, Gam51g]. Eins [GT56, GT58]. Einstein [Ano05, Gam42a, Gam88a, Kle05, OM18, Rin09, Rin11, Wei13, Gam42i]. electric [Lon72]. Electricity [Gam67d, Gam67c]. electrokinetic [Rig06, Rig07]. electrons [Gam33k, Gam91]. Elektronen [Gam51a, Gam51g]. Elementary [AG68, Gam67j, Gam33c]. Elemente [Gam51a, Gam51g]. Elements [Alp48, ABG48, AH48, BBFH57, Fre14, Gam34i, Gam35c, Gam41d, Gam42e, Gam46b, Gam47c, Gam48e, Gam86, Hoy54, Lew34, SCG08, Wat48, AHG49b, Gam32h, Gam33h, Gam33f, Gam34b, Gam66b, Hoy46, Pen79, TTL07, Tri10]. elle [Pol72]. elliptical [BGK50]. Emilio [Wil71]. Emission [BG36]. Empirical [Rac35, Gam34c]. Empiricism [Kra91b, Kra91a]. Empirische [Rac35, Gam34c]. encyclopedia [HTB^{+}07]. End [Gam4xa, Gam53g]. energía [GS42]. Energie [Gam47d]. Energiequelle [Gam38a, Gam38c]. Energy [Bet39, CGT38, Gam33g, Gam38d, Gam39a, GT39c, Gam40b, GS42, Gam45a, Gam46a, Gam47b, Gam47d, GC49, Gam52b, Gam64c, Gam75a, Gam05, Gam11a, MW88, Gam33d, Gam38a, Gam38c, Tuc72, Mul41]. Energy-Producing [Gam39a]. Energy-sources [GC49]. enquiry [Alp73]. Enrichment [FB12a, FN12]. Enrico [MR86, Wil71]. Enrico-Fermi [MR86]. Entropy [BG61, YvdM72]. entstehenden [HS39]. Epilogue [Gam63c, HPA97a]. Erdalkalimetalle [HS39]. Erde [Gam54b, Gam69a]. Erinnerungen [Gam60]. Ernest [Bad71]. Errata [Ano94, Gam49d, GY56]. Erratum [AHG49a, Gam47c, Gam67d]. Erwin [Pra93]. escale [Gam55d]. Essay [Gam32d, Gam34d]. Essential [GO06]. est [Pol72]. est-elle [Pol72]. estelar [GS42]. était [Ano05, Kle05]. étoiles [Gam38g]. Eugen [Igg66]. Europe [Fer68, Fer71]. Eve [Gam6x]. Ever [Gam51h, Wei13]. Everyday [Nug54]. Everything [Uns60]. evidence [Hoy90]. evolución [GS42]. Evolution [Fre14, Gam38d, Gam38f, Gam39b, Gam39d, Gam39f, Gam40b, Gam40d, GS40, GS42, Gam43a, Gam44a, Gam45a, Gam45b, Gam47d, Gam48b, Gam51e, Gam52b, Gam52e, Gam05, M.40a, Mul41, ZN73, Cla68, Cla83, Gam38g, Gam53f, Nad95, Nug54]. Evolution* [Gam39e]. Evolutionary [Gam56a]. Evolyutsiya [ZN73]. Excitation [Gam33e, Gam31b]. excited [DG31]. Exclusion [Gam59c]. Excursion [Rin11, Rin09]. Exhaustion [Gam51c]. existence [HS39, Gam49c]. Existenz [Gam49c]. Expanding [AHG48, AHG49a, AFH53, GT39a, GT39b, Gam46b, Gam47c, Gam53d, Gam68d, Gam56c, Lal14, Wat72]. expansión [Lal14]. Experimente [Gam65a]. Experiments [Gam65a]. expert [Poh62]. Explained [Gam63d]. Explaining [GHJ47]. Explode [Gam41a, Gam51h]. Explores [Ano44, Gam44b, Pom44, Gam46c]. Exploring [Fre10]. Explosions [Gam51c]. expression [Woe67].

F [Det55, GH45]. facetious [Kle00]. facsimiles [Bey49]. Facts [Atw54, Bar53, Gam53e, GT56, GT58, Gam77, Gam88b, Gam11b, Gla52, Joh54a, Joh54b, R.53]. Factual [Kra05]. Fast [Gam36c]. Feinstruktur [MP31]. Fermi [MR86, Wil71]. festschrift [MW88]. few [Che94b]. FIANa [Gor90]. Field [Opp28]. Fifth [TGF39]. Fifty [Gam46a, Gam47b, Gam11a]. Fine [Bek86, Gam30b, Che94b, MP31]. Fine-Structure [Bek86]. Finnish [Gam64b]. First [ABN02, FB12a, Wei77, Wei93, Stu18]. fisica [Gam01a, Gam14]. Fission [Gra64, Stu94, Gam29e, GA71, Stu13, Ano94]. Flight [Gam48c, Gam04b]. Fluid [GR31]. focusing [Rig06]. Fois [Ano05, Kle05]. force [Gam10]. forgotten [Sch12b]. Formation [ABN02, Alp48, AGH67, Gam54e, Wat48, Gam33h]. forms [Cri72]. Forscher [Gam65a]. forza [Gam10]. Fossils [FB12b]. Foundations [BPP^{+}11, Bey49, Fea62, Fre61, GC60, GC69, GC76]. freedom [GLI26]. Freie [Gam51a, Gam51g]. French [CCJ^{+}34, Gam26, GR33, Gam33d, Gam35e, Gam36a, Gam36b, Gam38g, Gam55d, Gam62e, GS67, Gam68f, Gam01b, Hei34, Kle05, Pol72, SG12]. Freud [Nug54]. Frisch [GHJ47, Pra93, Smi61b, Stu13]. Frontiers [GC60, GC69, GC76, Fea62, Fre61]. fuel [Gam4xb]. function [Cri72]. Fundamental [Gam33b, Oku02, Alp73, Bey49, Gam27]. Funny [BK21]. Further [Jud01]. fusèes [Gam62e]. fusion [Rei72a, Rei72b, Tuc72]. Future [Gam41b, Gam48a, Gam59b, Tuc72, Dix61, Gla49]. fysiken [Gam66g]. fysikken [Gam68e].

G [Ano40, Ano44, M.40b, Pra93, Che94a, Fre94b, Fre94c]. G. [Ano69, Cer05, Gor90, Hun49, Öpi69, Pau32, Sch69]. Galaxies [BGK51, FB12a, Gam48c, Gam48e, Gam53d, Gam86, Gam04b, BGK50, Gam69c, JLA^{+}04]. Galaxy [BC05, SCG08]. Galileo [Gam88a]. Galloping [Gar07]. Game [Kra96c]. gamma [BBC^{+}07]. Gamov [Ano95, BCY95, Che94a, Fre94c, Pol72, Sch69]. Gamow [AH71, Ano69, Ano98, Azi67, Bar53, Ber68, BKST^{+}07, BCY95, Cas12a, CBKZ^{+}09, Dan65, Det55, Dwi66, Fea62, Fre61, Gre00, HPA97b, Hob02, Huf09, Hun49, Igg66, Joh54a, Las62, Meg61, Meg62, Oku02, Öpi69, Pau32, Per03, Pol58, Rac35, Rya06, Sit64b, Smi61a, Sus69, Uns60, Van62, Wil71, AH72a, AH73, Alp73, AH96, AM23, Ano40, Ano44, Ano47, Ano50a, Ano50b, Ano55a, Ano56, Ano68, Ano00, Atw54, BBC^{+}07, Bet97, Bin58, Blo88, C.48, Cas12b, Cer05, Che94a, Che94b, Che95, CMT17, Dem07, Dix61, Fie59, Fre40, Fre94b, Fre94a, Fre94c, Gam55b, GG76, GNF^{+}97, GO06, Gla49, Gla52, Gor90, Gre90, GN00, HPA97a, Har01, Har32, Hen63, Her66, Hun49, Inf48, Joh54b, K.62, Kle66, Kle00, Kra91b, Kra91a, Kra96b, Kra96c, Kra05, Kuh67, Lal14]. Gamow [Lin00, Lip86, M.40b, Mar08, Mat42, Mat66, McC40, MP31, MR86, Mis08, Mul41, Nad95, Nan04, Nov07, Nug54, PD00, Ped12, PG66, Pol72, Pom44, Pra93, Pus96, Pus07, R.53, Rac35, RSJ07, Ray04, Ray05, Rei72a, Rei72b, Ric97, Ric71, Rig06, Rig07, Rin09, Rin11, Rog62, Rub97, Rub02, Rya05, Sab96, Sal96, Sch12a, Sco07, Seg11, Sha72, Sha07, Sha53, Sit64c, Ske54, Smi61b, Sta99, Stu71, Stu86, Stu97, Tel97, TTL07, Ula72, URR86a, Uns60, Van53, Wat01, Wat02, Wei68, Wei13, vdBS12, Hoo93, Jud01, M.40a, Dys87, The01]. Gamows [Gar07]. gamowsche [MP31]. Gas [Gam54e]. Gay [Gam91]. Gaylord [Nug54]. GBP10.95 [Per03]. GBP16.95 [Per03]. Geb [Uns60]. Geburt [Uns60, Gam47d]. Geburtstag [Gam60]. Geistesgeschichte [Kuh67]. gemacht [GT56]. General [Gam35a, GBK48, Ped12, Hei34, KE05, Wei72a]. générales [Hei34]. Generalizations [GT37]. Generator [HS19]. Genes [Wat01, Wat02, The01]. Genesis [AH01, Lew34]. Genetic [Hay98, Ric97, Dem07, Nan04, Woe67]. Genetics [Gam64a, Gam81, Gam68a]. genius [Wil83]. Geniuses [Cas12a, Cas12b, Seg11]. Genomics [Cas12a, Cas12b, Seg11]. genshiryoku [Gam42]. George [Ano47, Ano50a, Ano55a, Ano56, Ano68, Ano98, Ano00, Atw54, Bin58, C.48, Cas12a, Det55, Dix61, Fre40, Gla49, Gla52, Har32, Hen63, Her66, Hoo93, Inf48, Joh54a, Joh54b, K.62, Kle66, Kuh67, M.40a, Mat42, Mat66, McC40, Mul41, Nug54, Ped12, Per03, PG66, Pol58, Pom44, Pra93, R.53, Ric71, Rog62, Sco07, Sha53, Sit64b, Sit64c, Smi61a, Smi61b, Stu71, Sus69, Uns60, Van53, Wil71, AH71, Alp73, AM23, Ano50b, Ano95, Azi67, Bar53, Ber68, BBC^{+}07, BCY95, Cas12b, Che95, Dan65, Dem07, Dwi66, Dys93, Fea62, Fre61, Fre94a, Fre94c, Gam55b, GG76, GNF^{+}97, GO06, Gre00, Gre90, HPA97a, Har01, Hob02, Huf09, Kle00, Kra05, Las62, Meg61, Meg62, Nan04, Nov07, Oku02, Pus96, Pus07, RSJ07, Rei72a, Rei72b, Rub97, Sab96, Sal96]. George [Sch12a, Seg11, Sha72, Sha07, Sta99, Tel97, Uns60, Van62, Wei68, Wei13]. Georges [Kra18]. Georgiy [Sco07]. Gerald [Ske54]. Gerhard [Igg66]. German [DG31, GI26, Gam28b, Gam29b, Gam29e, GH29, Gam29a, Gam31b, GH32, Gam34c, Gam34a, Gam38a, Gam38c, Gam47d, Gam49c, Gam54b, GT56, GT58, Gam60, Gam63b, Gam65a, Gam67h, Gam69a, Gam80, HS39, Hou30, MP31, PG27, Rac35, Uns60, vW35]. Germany [Gam51b]. Gets [GO06]. giant [Sha07, Gam45b, GK45]. Giants [Gam39b, GT39c, GL50]. Gino [Cas12a]. Girls [The01, Jud01, Wat02, Wat01]. Gödel [BPP^{+}11, Rin09, Rin11]. governa [Gam10]. governs [Gam10]. gradients [Rig07]. Gravità [Gam10]. Gravitaatio [Gam64b]. Gravitation [Gam62e, Lin00, Wei72a, Gam62e]. gravitational [Dir72]. Gravity [Gam56b, Gam61d, Gam62b, Gam62c, Gam67b, Gam67d, Gam67c, Gam02, Gam64b, Gam65b, Gam10, Las62, Hen63, Rog62]. Grawitacja [Gam65b]. Great [Gam88a, Sch12b, GT39a, GT39b, GT39d]. Green [Dys93]. Grenzfragen [GT56, GT58]. Griffin [Det55]. Group [Far01]. Growth [Gam34j, Ske54]. Guide [Sha53]. guided [HG07]. Gustaf [Gam40e].

H [Det55, Gam52d, Pra93, Uns60]. Haggerty [Rog62]. Hahn [Gam66d]. Half [Gam50a]. Hall [Smi61a]. Hans [Gam42i, BL09]. hardback [Per03]. Hardy [Pra93]. Harper [Wil71]. Harvey [Det55]. Hauptserie [PG27]. Heart [Gam61e, Gam62d]. Heat [Gam64c]. Heating [Lon72]. Hecht [GHJ47]. held [CCJ^{+}34, MR86]. Helical [Gam55f]. Helix [Wat02]. Henry [Ano55a, Gam50b]. Herman [Alp12]. Herrn [Rac35]. Heuer [Gam53g]. heutigen [Gam34a, Gam37b]. High [HS19]. Higham [Igg66]. Himself [GY67, Sus69]. histoire [Gam68f, Gam01b]. Historical [Kra96a]. historie [Gam68e]. History [Gam42a, Gam42i, Gam52a, Gam52d, Gam54b, Gam67e, UM86a, UM86b, WP85, AWCT09, Gor90, GA71, KLR13, Nye02]. Holland [Dys87]. home [Wei68]. Homogeneous [Gam68d]. Honest [Jud01]. honor [MF69]. honyaku [Gam42, GDWWxx]. Horizons [BPP^{+}11]. Hot [Hoy54, Ray04, Ray05, Kra96c]. hour [Gam50a]. Hoyle [AWCT09]. HQ [KLR13]. HQ-3 [KLR13]. Human [Gam46a, Gam47b, Gam11a]. Hydrogen [Gam51c, Sal52, Hoy46]. hypothesis [Pol72, Tri10].

i.e [HPA97b]. Ice [Gam48f]. Ideas [Gam34g, Gam65a]. Ideen [Gam65a]. IEF [Rig07]. II [Gam51b, Rig07]. Illus [Joh54a, Sit64b]. illustrations [Gam38c]. Illustrious [Fer68, Fer71]. immigrants [Fer68, Fer71]. Immobilized [Rig07]. Importance [Bad71, Gam41d]. Impossibility [Bre57]. Individualitat [Kuh67]. Infant [FB12b]. infini [Gam55d]. Infinity [Gam47e, Gam55d, GT56, GT58, Gam77, Gam88b, Inf48]. Influence [Bet97]. Influential [Rog10, Jud01]. Informal [AH71, Gam70, Ric71, Stu71]. Information [Bre57, Gam54d, Gam55c, Gam55e, GRY56b, GRY56a, Gam57a, YvdM72]. Initial [AFH53]. Innocence [Stu18]. Inquiry [Gam53g]. inspiration [Rub02]. inspire [CMT17]. institut [CCJ^{+}34]. Institute [CCJ^{+}34, Gor90]. Institution [HPA97b]. intellectual [Fer68, Fer71]. interest [Ped12]. Internal [GL33]. International [BKST^{+}07, CBKZ^{+}09, KLR13, MR86, CCJ^{+}34, Gam34e]. interplanètaires [Gam62e]. interplanetary [Gam62e]. Interpretation [Ano94, Stu94, Stu13]. Interrogating [OM18]. Interview [Wei68]. intraatómica [GS42]. Introduction [CO07, JLA^{+}04]. Intuition [Rub02]. invariance [HN72]. invented [Sch12b]. Invention [Hay98]. inventions [Sch12b]. Inventor [Wil71]. investigación [Gam63a]. ippan [GDWWxx]. irradiation [HS39]. Irving [Gam50c]. ISBN [Cas12a, Dys02, Per03]. ISBN-13 [Cas12a]. isoelectric [Rig06]. Isomeric [Gam34f]. Isotropic [Gam68d]. Israel [UM86a, UM86b]. iste [Gam80]. Italian [Gam32i, Gam01a, Gam10]. Italy [MR86]. IV [Far01, Gam93b]. Ivanenko [Oku02]. ix [Hob02].

J. [Rig95]. James [Jud01, Rog62, Wil71, Sch12b]. Jammer [Kuh67]. Jellinek [Gam51a, Gam51g]. Jeux [GS67]. Jim [Jud01]. John [Fea62, Fre61, GHJ47, Igg66, Nug54, Ske54, vN96]. Joseph [Det55]. journals [Bey49]. Journey [Dys02, FR13, Gam80]. Jr. [Rog62]. Jubilee [BCY95]. Judith [Dys02]. July [BL09, KLR13, MR86]. June [Gam37b, KLR13]. junge [Gam60]. Juni [Gam37b].

kagaku [GDWWxx]. Kaleidoscope [UM86a]. Kalinga [Ano56]. Kaliums [PG27]. kankei [Gam42]. Kapitza [Fre94b]. Kaplan [Det55]. Karl [Gam51g, Gam51a]. Kenneth [Gam53g]. Kernen [DG31]. Kernmassen [vW35]. Kerns [GH29]. Kernumwandlungen [Gam38c, Gam38a]. Key [Oku02]. Kinetics [Gam36a, Gam36b]. Kingdom [Hob02]. Klaus [Kuh67]. Know [Rub97]. knowledge [Gam65e, Kle00, Lal14]. Komplementaritat [Kuh67]. Korrespondenz [Kuh67]. Kosmos [Gam80, vdBS12]. Kurt [BPP^{+}11]. Kvanteteoriens [Gam68e].

L [Fre94b, Gam26, Kuh67, Oku02, Rog62]. Lancelot [Gam49f]. Lanczos [Rin09, Rin11]. Landau [Oku02]. Langmuir [Gam50c]. Large [Alp73]. last [Alp73]. later [AH96]. Law [Gam49f]. Layman [Gam42a, Gam42i, Sit64c, Sit64a]. Laymen [Gam42a, Smi61b]. Learns [Atw54, Bar53, Gam53e, Gam11b, Joh54a, Joh54b, R.53]. Lebensgeschichte [Gam54b]. legacy [Pus07, TTL07, Tur08]. Legend [OM18]. Lehrbuch [Gam51b, Gam51f]. Lemaître [Kra18]. Let [Ray04, Ray05]. Letter [GT37, GT38a, GT39c, Gam67f, Gam67g, Mar08]. letters [Gam67e]. Level [Ped12]. Levels [Gam33g, Gam75a, Gam32a, Gam33d]. Levi [Igg66]. L’évolution [Gam38g]. L’hypothèse [Pol72]. LI [Rut27]. Life [Cer05, Coc46, Gam46a, Gam47b, Gam53e, Gam54b, Gam63e, Gam63g, Gam11a, Gam11b, Nug54, Pra93, Ske54, Wei13, Atw54, Bar53, Joh54a, Joh54b, Nug54, R.53]. light [TTL07]. like [Gam42f]. Limiting [GIL28, GIL02b, Oku02]. limits [Gam34c, Rac35]. Line [AH71, Fre94a, Gam70, Gam93c, Ric71, Stu71, Fre94c, Wil71]. lines [PG27]. Linien [PG27]. Liniya [Gam93c]. Lippmann [Ano47]. Liquid [Ano94, Stu94, Stu97, CMT17]. Liquid-Drop [Ano94, Stu94, Stu97]. literature [HG07]. Little [Pra93]. Lives [Wil71]. Living [Gam55c, Gar07]. livre [Ano05]. Lorentz [Gam61f]. L’Origine [Gam33d]. Louis [Gam54h]. Lovett [Kuh67]. Lunar [Gam63c]. Lune [Gam55d].

M [Det55, Fea62, Fre61, Gam80]. M. [Gam26, SG12]. MA [Dys02]. Mach [Rin09, Rin11]. made [Haw11, Rub02, Wei13]. Magician [Gre90]. Mai [Gam34a]. Man [Ano99, Gam47a, Gam63d, Gam66c, Kuh67, Ske54]. människan [Gam47a]. Many [Gam42f]. March [BL09, Huf09]. Margenau [Gam50b]. Martin [Gam69b]. Martynas [Sus69]. Marvin [Pol58]. Mass [Gam30c, vW35]. Massive [Gam43a]. Masur [Igg66]. materials [Ano50b]. Materie [Gam51b, GI26, Gam51f]. math [GS58a, GS58b, GS67, Pol58]. mathematical [Gam54f, Nye02]. Mathematician [Pra93]. Mathematics [BPP^{+}11, BP64, BP81, Inf48, Ula72, URR86a, FF91, URR86b]. mathématiques [GS67]. Matter [Gam58b, Gam59d, Gam65c, Pra93, GI26, MF69, WH07, Dan65, Dwi66, Bin58, Fie59, PG66]. matters [Rei72a, Rei72b]. Max [Cas12a, Kuh67, Cas12b, Seg11]. May [Gam34a, TGF41]. Meaning [Nug54]. Mechanical [Mis08]. Mechanics [Gam40a, GC28, GC29, GH29, Kuh67]. Mechanism [Gam33e]. Medical [LT56]. Meet [GHJ47]. Meeting [CCJ^{+}34]. mehr [Gam63b]. Meitner [Stu13]. Memoirs [Gam60, Wil71, Dys02]. Memorial [BKST^{+}07, PD00, Rei72a, Rei72b]. Memories [AH72a, Tel97]. memory [BBC^{+}07]. mentor [AM23]. Meson [Gam47f]. Meson-Absorption [Gam47f]. Metal [BC05, Fre10, FN12]. Metal-Poor [BC05, Fre10, FN12]. metals [HS39]. Method [Ske54]. methodologies [Rig06, Rig07]. Meyer [Kuh67]. Meyer-Abich [Kuh67]. Michael [Kuh67, Rog62]. microcosmos [Gam80]. Microscope [BR85]. Microwave [Alp12, Nov07, Pen72, Wil79]. migration [Fer68, Fer71]. Mikrokosmos [Gam80]. Mind [Pra93]. Minutes [Wei77, Wei93]. Mirovay [Gam93c]. misadventures [Jud01]. Mistera [Gam94]. mittels [HS39]. Mixed [Gam48d]. Model [Ano94, CG39, Gam38b, GK45, Stu94, Stu97, Wat72]. Modern [BF86, BBBM90, CO07, Gam34g, Gam40c, Gam50b, Gam54c, Gam56b, Gam55a, GB68, Gam90, Igg66, Kra18, Nug54, Smi61b, Wei77, Wei93, Gam27, Gam38g, Nye02]. moderne [Gam38g]. modernen [GT56, GT58]. molecular [Woe67]. Molecules [Gam50c]. Mond [EG57a]. monde [SG12]. Moon [Sha53, EG57b, EG57a, Gam51d, Gam53a, Gam55d, Gam59a, GC71, Ano55a, Sha53, Ske54, Van53]. Moore [Gam66c, Kuh67, Sha53]. Most [Rog10, Haw11, Jud01]. motion [GLI26]. Move [GHJ47]. Moving [GR31, Wei72b, Wei85]. Moya [Gam93c]. Mr [Ano02, Gre00, Hob02, Per03, Gam11b, Gam12, Fre40, Joh54a, Joh54b, Ped12, Pom44]. Mr. [Bar53, Ber68, Gam39c, Gam42g, Gam44b, Gam46c, Gam53e, Gam65d, GY67, Gam80, Gam93a, Gam94, GO06, Rac35, Sta99, SG12, Boy93, Hoo93, Ano44, Atw54, M.40b, Mat66, McC40, Pra93, R.53, Sus69]. muerte [GS42]. München [Uns60]. Muscles [Gam67i]. Music [Ano47]. My [Gam70, AH71, Gam93c, Wil71, Ric71, Stu71, Wil71]. Mystery [FR13].

N [Nug54]. nach [Gam60]. Nachweis [HS39]. Nacimiento [GS42]. Naming [Gam68b, Kra14]. Nathan [Ske54]. Nature [Gam50b, Gam68c, Alp73]. Near [Gam4xa]. Nebulæ [GT39a, GT39b, GT39d]. Negative [BG61, Gam34h, Gam35b, YvdM72]. Nelson [Igg66]. Neuere [Hou30]. Neumann [vN96]. Neutrino [CR72, GS41, Gam41d, GS46, Gam49c]. Neutrinos [GS40, Gam1 , Gam42h, Gam48g, Gam49c]. Neutron [Alp48, GT38b, SCG08, HS39]. Neutron-Capture [Alp48, SCG08]. Neutronen [HS39]. Neutrons [Gam33f, Gam36c]. Newton [Det55, Gam62e]. Nicht [Gam63b]. Nickel [Hoy54]. Niels [Gam60, Kuh67, Gam60, Gam63d, Gam66c, Kuh67]. Niemeyer [Gam54h]. Nightmares [Pom44]. ninetieth [Che94a, Fre94c]. niveaux [Gam33d]. No [Gam63b]. Nobel [Kra17]. non [Ped12]. non-specialists [Ped12]. nonconservative [GLI26]. nonlocally [CMT17]. North [Dys87]. Note [Gam29b, Rac35]. nouveau [SG12]. Novæ [Gam38e]. noyau [Hei34]. noyaux [CCJ^{+}34, GR33, Gam35e]. nucleaires [Gam33d, Gam36b]. Nuclear [Ano94, BB36, Bet97, Gam28a, Gam30c, Gam32a, Gam32g, Gam33b, Gam33g, Gam34g, Gam34h, Gam34i, Gam35c, Gam36a, Gam37a, Gam38d, Gam38b, Gam39d, Gam39e, Gam47f, GC49, Gam75a, Gra64, Hoy54, Mla98, Ros72, Sal52, Sal96, Stu94, Stu18, Bey49, Gam32b, Gam32i, Gam33d, Gam35d, Gam36b, GA71, Hug93, RSJ07, Stu13, Tuc72, vW35, Gam38a, Gam38c]. nucleare [Gam32i]. nucléares [Gam36a]. Nuclei [BB36, DW48, Gam29c, Gam31a, Gam32c, Gam34f, Gam37a, Har32, LW46, RAC^{+}29, Wat46, CCJ^{+}34, DG31, GH32, GR33, Gam33i, Gam34c, Gam34e, Gam35a, Gam35e, GBK48, Rac35]. Nucleic [Bre57, Gam54d, Gam55e, GRY56b, GRY56a, Gam57a]. nucleocosmochronologies [Fow72]. Nucleoproteins [DGS^{+}56]. nucleosynthesis [AWCT09, Cla68, Cla83]. Nucleus [FR13, Gam30a, Gam32f, Gam34j, GC49, Gam61a, Stu97, Gam28b, GH29, Gam29a, Gam32d, Gam32h, Gam34d, Gam93b, Hei34, Hou30, vdBS12, Meg61, Smi61a, Smi61b]. Numbers [Alp73]. Numerology [GM54, Gam68c].

O [GHJ47]. Obituary [Ano68, Gam69b, Har07]. Object [GR31]. Objects [Gam67j]. observability [Gam27]. Observational [Gam68d]. Observations [Gam50c, Fow72]. Occasion [Fre94a]. Occurring [Hoy54]. October [CCJ^{+}34, Far01]. octobre [CCJ^{+}34]. Odessa [CBKZ^{+}09, Rya05]. old [CMT17, Fow72]. ondes [Gam26]. One [FB12a, Gam47e, Gam77, Gam88b, GLI26, Jud01, GT56, GT58, Gla52, Inf48, Nug54]. One-Shot [FB12a]. Ontstaan [Hun49]. Opinion [Ano47]. Oppenheimer [Rig95]. Ordinary [Cas12b, Seg11, Cas12a]. Origin [ABG48, Ano94, Gam35c, GT39a, GT39b, GT39d, Gam42e, GH45, Gam46b, Gam47c, Gam48e, Gam48f, Gam51e, Gam53d, Gam53b, Gam63e, Gam75a, Gam86, Stu94, Wei77, Wei93, AHG49b, Gam33d, Gam53f, Gam66b, Gam69c, Pen79, Rut27, Sin04]. originally [Bey49]. Origins [BK21, Cas12a, Cas12b, Igg66, Seg11, Tri10]. Orr [Det55]. ost [Gam32c]. Other [Gam61e, Gam62d, Rei72a, Rei72b]. Otto [Pra93, Smi61b, Gam66d]. Our [Bek86, Gam41a, Gam51h, Gam69a, Jud01, Sch12b]. Ours [Gam42f]. Outline [Gam33i, Gam93b]. Outlines [Gam32f]. Ovenden [Rog62]. Overlapping [Bre57]. Own [Bek86].

P [Ano55a, Per03, Fre94b]. Pacific [Ano98]. pages [Cas12a, Hob02, Wil71, Wil71]. Panel [GNF^{+}97]. paper [Hob02, Tur08, Gam51b]. Paperback [Gam65d, Gam93a, Gam12, Sta99, Per03, Hoo93, Ber68, Mat66, Ped12, Pra93, Boy93]. Papers [BF86, vN96, Ano50b, Gam55b, GG76, Haw11]. Part [Rig06, Rig07]. particle [Gam33c]. Particles [AG68, CG30, Gam33b, Lon72]. Pasadena [Tri10]. Past [Dix61, Gam41b, Gam48a, Gam59b, Gla49]. Patrick [Sha53]. Patterns [Ano47]. Paul [Det55]. peace [MW88]. people [Ber68, URR86b]. period [Coc46]. Perseus [Dys02]. Personal [Tel97, URR86a, Coc46]. personality [Pus07]. Petersburg [BCY95, PD00]. Petri [Uns60]. pH [Rig07]. phase [Gam26]. phase-wave [Gam26]. phases [Gam26]. Phenomena [Gam36c, Gam50c]. Phil [Gam67e]. Philosophical [Gam42j]. Philosophy [Gam50b, Les90, UM86a, UM86b]. Photograph [GR31]. Photonen [Gam51a, Gam51g]. Photosynthesis [BG61]. Phys. [Gam47c]. Physical [AFH53, Gam39f, Gam42j, Gam50b, Gam62f, Les90, M.40a, Pee71, Pee93, Nye02, Gor90]. physicist [BBC^{+}07, Wil71]. Physicists [Gam88a, Kuh67]. Physics [Anoxx, Azi67, BB36, Bet97, Dys93, Gam40c, Gam49f, GC60, Gam61b, Gam62a, Gam65a, Gam66e, Gam66f, GB68, GC69, Gam72, Gam75b, GC76, Gam85, Gam14, Kuh67, LT56, MR86, Mla98, Oku02, RW64, Smi61b, Stu18, TGF41, Bey49, Che94b, CR72, FF91, Gam27, Gam38g, Gam49a, Gam56c, Gam66g, Gam68f, Gam68e, Gam01b, Gam01a, Haw11, HN72, Hug93, KLR13, MW88, Meh75, Nad95, RSJ07, WP85, Wei72b, Wei85, WH07, CCJ^{+}34, Fea62, Fre61, GF42, TGF39, Gam50b, Gam54h, Her66, Kle66, Meg62, Van62]. Physik [Gam51b, Gam51f, Gam65a]. physique [CCJ^{+}34, Gam38g, Gam68f, Gam01b]. Pioneer [Kra18]. Planet [Gam63f, Gam69a, Gam69a]. Planetary [GH45, Gil12]. Planets [Gil12]. point [Gam38g]. polariz [Luk70]. Polish [Gam65b]. Politics [Dys02]. Polymath [Har01]. Polypeptide [GM54]. pomme [Gam62e]. Poor [BC05, Fre10, FN12]. Populations [Gam48d]. portraits [Far01]. positive [Gam33k, YvdM72]. Possibilities [Gam39f]. Possibility [Gam36c]. Possible [AG68, Fre40, GS40, Gam54f, Gam54g]. potassium [PG27]. Pp [Gam51b, Dys02, Joh54a, Sit64b, Smi61a]. Pre [Gor90, Ped12]. Pre-history [Gor90]. pre-university [Ped12]. Precious [FB12b]. precision [HS19, Tur08]. Predystoriya [Gor90]. preface [Cla83]. Prentice [Smi61a]. Prentice-Hall [Smi61a]. Present [Dix61, Gam41b, Gam48a, Gam59b, Gla49]. Press [Ano02, Boy93, Dys87, Hob02, Joh54a, Per03, Wil71]. Prevalence [Blo88]. Price [Dys02]. Priklyucheniy [Gam94]. Primordial [Gam54e, TTL07]. principal [PG27]. Principle [Gam49f, Gam27, Gam58c, Gam59c]. Principles [Cla68, Cla83, Gam40a, Gam64a, Gam81, Pee93, Gam68a, Wei72a]. Prism [UM86b]. Prize [Ano56, Kra17]. probabilities [DG31]. Probability [BG36, Gam47f]. Problem [CGT38, GL50, CMT17, Gam33j, Gam34d, GRY56b, GRY56a, Gam57a, GY58]. Problems [Cer05, Oku02, Gam35a, GBK48]. Proceedings [BCY95, MR86, BKST^{+}07, CBKZ^{+}09]. process [Gam32e, GN00]. Producing [Gam39a]. Production [Bet39, GT39c, Gam41d]. Prof [Ano56]. Professor [Ano68, Wei68]. Profile [Gre90]. Progress [Bai53, Det55]. Properties [Gam55f, Gam68d, MF69, CCJ^{+}34]. propriétés [CCJ^{+}34]. prospects [CR72]. protégé [AM23]. Protein [Gam54g, GY55, GY56, GY58]. Proteins [Bre57, Gam54d, Gam55e, Gam54f, GRY56b, GRY56a, Gam57a]. Protogalaxies [AGH67, Gam54e, Gam53b]. Proton [Gam35b, Gam33c, Kav72]. proton-proton [Kav72]. Protons [Gam34h]. Prout [Tri10]. Psychopathology [Nug54]. Public [Ano47]. Publications [HPA97a]. Puzzle [GS58a, GS58b, GS67, Pol58]. Puzzle-math [GS58a, GS58b, GS67, Pol58].

Quantenmechanik [GH29]. Quantentheorie [Gam28b, Gam29b, Gam29e, Hou30, Kuh67]. Quanti [Gam01a]. quantica [Gam32i]. quantique [Gam68f, Gam01b]. Quantities [AG68]. Quantum [Azi67, Gam28a, Gam32b, Gam35d, Gam66e, Gam66f, Gam72, Gam85, Gam07, GC29, Her66, Kuh67, Mis08, Opp28, Gam28b, Gam29b, Gam29e, GH29, Gam68f, Gam68e, Gam01b, Gam01a, Haw11, Hou30, KLR13, Gam32i].

Quasi [Gam67a]. Quasi-stellar [Gam67a]. Quasistellar [Gam67j]. Quelques [GS67]. Questioners [Kuh67]. qui [Gam68f, Gam01b].

R [Det55, GHJ47, Ske54, Smi61b, Gam80]. Radiation [AGH67, Gam33a, AH90, Che94b, Gam31b, Nov07, Wil79, Alp12]. Radiations [RCE30, RCE51]. radioactifs [GR33]. Radioactive [Gam32f, Gam32g, Gam34i, GC28, GC29, Gam29b, GH29, Gam31b, GR33, Gam34b, Rut27, RCE30, RCE51]. radioactivists [Hug93]. Radioactivity [Gam30a, Gam31a, Gam32c, Gam46a, Gam47b, Gam11a, GH32, Har32]. radioaktiven [Gam29b, GH29, Gam31b]. Radioaktivität [GH32, Pau32]. Radioaktivn [Gam32c]. Radioaktivnost [Gam30a]. Radioastronomy [Rya06]. Rakete [EG57a]. Ralph [Gam54h, Alp12, AM23, Har07]. Random [Web73]. Rapports [CCJ^{+}34]. Rate [GT38a]. rates [Kav72]. ratio [PG27]. Ray [BG36, BBC^{+}07]. rayons [Gam33d]. Rays [Gam30b, Gam75a, Gam33d, Gam33j, Rut27, SST72]. Re [Pra93]. Re-Reading [Pra93]. Reaction [Gam39a, Kav72]. Reactions [AHG48, AHG49a, Gam36a, GT38a, Gam39d, Gam39e, Hoy54, Sal52, Gam36b, Gam38a, Gam36a, Gam36b]. readership [Ped12]. Reading [Pra93]. readings [WH07]. Reality [Gam50b, Gam48g]. Received [Hoo93]. Reconstructing [Fre14]. recording [Luk70]. red [PG27, Gam39b, GT39c, Gam45b, GK45, GL50]. Red-Giant [Gam45b]. Reflections [AH72b, Pus07]. Reich [Uns60]. Reichenbach [Gam42i]. Reisen [Gam80]. Relation [AG68, Gam54g, Mis08, Gam54f]. Relative [Alp48, AH48, DW48, Gam41d]. Relativistic [Gam49e, Kra05]. Relativity [Gam42i, Gam42a, KE05, Nim11, Wei72a]. Remarkable [Rin11, Rin09]. Remarks [Gam61f]. Remember [Pra93]. Reminiscences [GNF^{+}97, Ric71, Ros72, Sal96, URR86a, Coc46]. Report [Gam38c]. reported [Bey49]. Reports [CCJ^{+}34]. research [CMT17]. Researchers [Gam65a]. reserves [Tuc72]. Responsibility [Rog62]. resulting [HS39]. resurfaces [CMT17]. Retrospect [BBBM90]. Returns [Ber68]. Rev [Gam47c]. Rev. [Gam69b]. Review [AH71, Ano44, Ano47, Ano50a, Ano98, Ano02, Atw54, Bar53, Bin58, Boy93, C.48, Cas12a, Dan65, Det55, Dix61, Dwi66, Dys87, Dys02, Fea62, Fie59, Fre40, Fre61, Gam40e, Gam42j, Gam42i, GHJ47, Gam49f, Gam50b, Gam50c, Gam51a, Gam51b, Gam51g, Gam51f, Gam52d, Gam53g, Gam54h, Gam66c, Gam66d, Gla49, Gla52, Gre00, Har32, Hen63, Her66, Hob02, Igg66, Inf48, Joh54a, Joh54b, Jud01, K.62, Kle66, Kuh67, Las62, M.40b, M.40a, Mat42, Mat66, McC40, Meg61, Meg62, Mul41, Nug54, Pau32, Per03, PG66, Pol58, Pom44, Pra93, R.53, Ric71, Rog62, Sha53, Sit64b, Sit64c, Ske54, Smi61a, Smi61b, Stu71, Sus69, Uns60, Van53, Van62, AM23]. Reviews [Ano40, Gam52a, Sit64a, Wil71]. Revisited [Mis08]. Revolution [Gam54h, Kle05, Ano05, Kle05]. Ribonucleic [GY55, GY56]. Riedman [Ske54]. Rio [GN00]. rise [Hug93]. Road [Kra96c]. Robert [Rig95, Alp12]. Roberts [Det55]. Rocket [EG57b, EG57a]. rockets [Gam62e]. Roger [Igg66]. Role [GS40, Gam52e]. Rose [Ske54]. Rotating [Gam46d]. roten [PG27]. round [DGS^{+}56]. round-table [DGS^{+}56]. Row [Wil71]. Rules [GT36]. Russell [Gre00, Hob02, Per03]. Russia [BCY95]. Russian [GLI26, Gam27, GIL28, Gam30a, Gam32d, Gam32e, Gam32f, Gam32c, Gam33a, Gam33h, Gam33f, Gam33i, Gam33j, Gam33c, Gam34b, Gam34j, Gam57a, Gam93c, Gam94, GIL02a, Gor90, ZN73]. Ruth [Ano47, Gam66c, Kuh67]. Rutherford [Bad71, Coc46, Wil83]. rystede [Gam68e].

S [Det55, Ske54, Uns60]. said [Ray04, Ray05]. Sarah [Ske54]. Satellites [Rog62]. say [Wei13]. Scanning [BR85]. Scatchard [Det55]. School [CBKZ^{+}09, MR86, Rya06]. Schrödinger [Pra93]. Schuman [Ano55a]. Science [Bai53, Dys02, Gam42a, GT56, GT58, Gam66c, Gam77, Gam88b, Hay98, Kuh67, Nug54, Rog62, Sit64c, Sit64a, Ske54, URR86b, UM86a, UM86b, Web73, Nye02, Sab96, Sha07, Gla52, Det55, Nug54]. Sciences [Gam62f, Kra17, Nye02, Gor90]. Scientific [Gam53g, Gam66d, Har01, Alp73, Bey49, HG07, Haw11, Pus07, Ano55b]. Scientists [Rog10, Jud01, Sch12b]. sconvolsero [Gam01a]. Scope [Ped12]. Second [Stu18]. Segrè [Cas12a, Wil71]. Selection [GT36]. Selective [Gam36c, GT38a, Gam38b]. Selig [GHJ47]. seltsame [Gam80]. Seminar [BCY95]. Sengo [Gam42, GDWWxx]. Sep [BCY95]. Separation [Gam48e, Gam86]. sept [Ano05, Kle05]. septième [CCJ^{+}34]. sequel [Jud01]. Series [Ske54, PG27]. Serious [GO06]. session [PD00]. seven [Kle05]. seven-times [Kle05]. Seventh [TGF41, CCJ^{+}34, Far01]. Several [Wil71]. shapes [HS19]. Shattuck [Igg66]. Shell [CG39, GK45]. Shell-Source [CG39]. Shook [Azi67, Gam66e, Gam66f, Gam72, Gam75b, Gam85, Her66, Kle66, Kuh67, Gam66g, Gam68f, Gam68e, Gam01b, Gam01a, Haw11]. Shoolery [Dys02]. Shot [FB12a]. shuppan [Gam42, GDWWxx]. Side [Gam61e, Gam62d]. Sigmund [Nug54]. Signatures [FB12a]. simple [Wil83]. Simpson [Nug54]. since [Meh75]. Single [GR31]. Sixty [FR13]. skakade [Gam66g]. Sketch [Gam34j]. Sketches [Pra93]. Sky [Gam58b, Gam65c, Gam59d, Bin58, Dan65, Dwi66, Fie59, PG66]. Social [Wil71]. Society [Ano98]. Sociology [UM86a, UM86b]. Soft [Nug54]. sol [GS42]. Solar [Gil12]. Solarized [Luk70]. Solvay [CCJ^{+}34, Far01, Meh75, CCJ^{+}34]. som [Gam66g]. Some [GT37, Tel97, Tuc72]. Somerville [Ske54]. Sonne [Gam47d, Gam67h]. Soul [Gam40e]. sound [Luk70]. Source [CG39, Gam38b, GK45, Gam38a, SST72]. Sources [Gam38d, Gam38c, GC49, Kuh67]. sous [CCJ^{+}34]. Space [Gam52f, Gam66a, Kra02]. Spacecraft [Rog62]. Spanish [Gam42b, GS42, Gam63a, Gam14, Lal14]. Special [Mis08, Nim11]. specialists [Ped12]. Spectrum [HS19, MP31]. Speculations [GT56, GT58, Gam77, Gam88b, Gla52, Tuc72]. Spektrum [MP31]. Spherical [BGK51]. Spider [Nug54]. Spin [Gam32g, Gam34i, Gam4xc]. sponsored [HPA97b]. Springer [Gam51b]. Springer-Verlag [Gam51b]. St [BCY95]. St. [PD00]. Stabilitätsgrenzen [Gam34c, Rac35]. stability [Gam34c, Gam35a, GBK48, Rac35]. stability-problems [Gam35a, GBK48]. Stage [Gam45b]. Stages [AFH53]. Stairway [Gam55d]. Stand [Gam34a, Gam37b]. Stannard [Gre00, Hob02, Per03]. Star [ABN02, Gam38b, Gam64d, Gam67h, Sit64b, Sit64c, Sit64a, Rub02]. Stars [BC05, Bet39, BBFH57, Fre10, FN12, GL33, GT38b, Gam40d, Gam41c, Gam43a, Gam43b, Gam44a, GK45, Gam51c, Gil12, Hoy54, Sal52, Wat48, BBC^{+}07, Gam33h, Gam38a, Gam38g, Gam38c, Nad95]. Start [Gam5x]. State [Gam33b, Gam54a, Hoy90]. States [BB36]. Stationary [BB36]. Statistical [GY55, GY56, YvdM72]. Statistics [Gam67a]. status [Gam34a, Gam37b]. steady [Gam54a, Hoy90]. steady-state [Gam54a, Hoy90]. Stellar [BGK51, CGT38, CG39, Fre10, Gam38d, Gam38f, Gam39d, Gam39e, Gam39f, Gam40b, GS40, GS41, GS42, Gam45a, Gam45b, Gam47d, Gam48d, Gam52b, Gam05, Gil12, Mul41, Cla68, Cla83, GS46, Gam67a]. Stern [Gam67h, Pol58]. Sternbildung [Gam47d]. Sterne [Gam38c, Gam38a]. Sternen [Gam67h]. storia [Gam01a]. Stories [Gam39c]. Story [Azi67, Gam66e, Gam66f, Gam72, Gam85, Kuh67, Gam68f, Gam68e, Gam01b, Gam01a, Gam42b, GHJ47]. Strahlen [Gam31b]. strange [Gam80]. Strangest [Fre40]. Street [Ano55a]. Stroenie [Gam32c, ZN73]. Stromberg [Gam40e]. Structure [Bek86, CCJ^{+}34, Gam29c, Gam30b, Gam34h, Gam34j, Gam37a, RAC^{+}29, Gam29a, GH32, Gam32b, Gam32h, Gam32i, Gam33i, Gam34e, Gam35d, Gam93b, Hei34, MP31, CCJ^{+}34, Rut27, ZN73]. Struktur [Gam29a, Gam51b, Gam51f]. struttura [Gam32i]. Studie [Kuh67]. Studies [Gam32f, UM86a, UM86b, Bey49]. Study [Gam34j, Gam63a]. stuff [Haw11]. style [Pus96]. Sub [Gam52b]. Sub-atomic [Gam52b]. subatomare [Gam47d]. Subatomic [Gam40b, GS42, Gam45a, Gam47d, Gam05, Mul41]. substances [RCE30, RCE51]. Successive [Gam29d]. Summary [Gam38c]. Summer [CBKZ^{+}09, Gin94, Rya06]. Sun [Gam39a, Gam40b, Gam41a, GS42, Gam45a, Gam47d, Gam49b, Gam52b, Gam05, Mul41, Gam4xb, Gam51h, Gam64d, Gam67h, Hun49, M.40a, Ano40, Ano47, Sit64b, Sit64c, Sit64a]. Supernovae [Gam49g, Gam1 , Gam42h]. Support [Gam42j]. sure [Hei34]. Surface [Gam67i]. survey [Gam65e]. Swedish [Gam46c, Gam47a, Gam66g]. Symposium [HPA97b, Ano98]. Synthesis [BBFH57, Hoy54, GY58, Hoy46]. System [GH45, Kra17]. Systems [Gam55f, Gil12, GLI26].

table [DGS^{+}56]. Technicolor [Gar07]. Technology [Gam62f]. Teil [Gam51a, Gam51g]. Teller [Dys02, Blo88, MF69, MW88]. Temperature [GL33, Mis08, Che94b]. Tension [Gam67i]. Tentative [Gam38e]. tenu [CCJ^{+}34]. Teoria [Gam32i, Gam01a]. ThC [MP31, MP31]. their [AM23, SST72]. Theoretical [Anoxx, GF42, Kra18, TGF39, TGF41, Che94b, Hei34]. Theoretischen [Gam51b, Gam51f]. Theorie [Gam31b, Gam34a, Gam37b, MP31, vW35, Gam26]. Theories [Gam42j, Gam52a, Gam52d, GB68, Kra96a, Gam66b]. théoriques [Hei34]. Theory [Alp48, Gam28a, Gam32f, GT37, Gam38e, GS41, Gam42a, Gam42i, GH45, GC49, Gam63c, Gam66e, Gam66f, Gam72, Gam85, Her66, Lin00, Lip86, Opp28, Stu86, Wei72a, Gam26, GI26, Gam28b, Gam29b, Gam29e, Gam31b, Gam32d, Gam32b, Gam32i, Gam33i, Gam33k, Gam34d, Gam34a, Gam35d, Gam37b, GS46, Gam54a, Gam68f, Gam68e, Gam93b, Gam01b, Gam01a, Hou30, Hoy90, MP31, vW35, BK21, Azi67, Kuh67]. there [Ray04, Ray05]. Thermal [AGH67]. Thermo [Gam38b]. Thermo-Nuclear [Gam38b]. thermodynamics [YvdM72]. Thermonuclear [AHG48, AHG49a, GT38a, Cri72]. Things [Gam5x]. Third [KLR13]. thirteen [Bey49]. Thirty [Gam66e, Gam66f, Gam72, Gam75b, Gam85, Azi67, Gam66g, Gam68f, Gam68e, Gam01b, Gam01a, Her66, Kle66, Kuh67]. Thoughts [Gam50c]. Three [Gam47e, GT56, GT58, Gam77, Gam88b, Wei77, Wei93, Inf48]. ThreeInfinity [Gla52, Nug54]. tierra [Gam42b]. Time [Bek86, Gam67b, Kra02, Mis08, Rog10, Alp73]. times [Kle05]. Tod [Gam47d]. today [CMT17, Smi61b]. tomorrow [Gam49a]. Tompkins [Ano44, Atw54, Bar53, Ber68, Boy93, Fre40, Gam39c, Gam42g, Gam80, Hoo93, Joh54a, Joh54b, M.40b, Mat66, McC40, Ped12, Pom44, Pra93, R.53, Sus69, Gam42g, Gam44b, Gam46c, Gam53e, Gam65d, GY67, Gam80, Gam93a, Gam94, GO06, Gam11b, Gam12, Hob02, Sta99, SG12, Ano02, Gre00, Per03]. Tompkinsa [Gam94]. Topological [Gam55f]. tosho [Gam42, GDWWxx]. tot [vdBS12]. tour [HG07]. Touring [Sha53]. Townes [Det55]. Tracks [Gam38f]. Traditions [KLR13]. Transfer [Bre57, Gam54d, Gam55c, Gam55e, GRY56b, GRY56a, Gam57a]. Transformation [GT37, Gam32h, Gam33f]. Transformations [Gam29d, Gam35c, Gam37a, Gam38c, KLR13]. Transition [Blo88, GIL28, GIL02b, Oku02, DG31]. translator [Gam54h]. treasury [FF91]. Tredive [Gam68e]. tree [URR86b]. Trent’anni [Gam01a]. Trente [Gam68f, Gam01b]. Trettio [Gam66g]. tribe [Gam91]. Triplet [Bre57]. truly [Alp73]. Truth [BPP^{+}11]. Tunneling [BR85, Nim11]. Turbulence [Gam52e, Gam52f, Gam66a]. Turbulent [Gam54e]. Turn [Kuh67, Rub02]. Twentieth [Dys02]. Twentieth-Century [Dys02]. Two [Gam47e, GT56, GT58, Gam77, Gam88b, Gla52, Inf48, Kra96a, Nug54, Gam67e]. Types [Gam48d].

U [Boy93]. Übergangswahrscheinlichkeiten [DG31]. Übersetz [Uns60]. Ukraine [CBKZ^{+}09, RSJ07]. uncertainty [Gam07, Gam58c]. undergraduate [Ped12]. understand [Gam32e]. Unendlichkeit [GT56, GT58]. unique [Pus96, Pus07]. Unitary [Gam49f]. United [Hob02]. Units [Gam68b]. Univ [Joh54a]. Universal [Gam4xc, GIL02a, Gam65e]. Universe [ABN02, AHG48, AHG49a, AFH53, DW48, Fre10, FN12, FB12b, GT39a, GT39b, Gam46b, Gam46d, Gam47a, Gam47c, Gam48b, Gam52c, Gam52e, Gam53d, Gam56a, Gam58a, Gam61c, Gam68d, Gam04a, Kra96a, LW46, M.40a, Mis08, Ray04, Uns60, Wat46, ZN73, Gam53f, Gam56c, KE05, Ray05, Sin04, Wat72, BBC^{+}07, Gam40e, Gam51e, Gam54a, Gam67e, Gam10, Rub97, Wei77, Wei93, Gam40e]. University [Ano02, HPA97b, Hob02, Per03, Wil71, Ped12]. universo [Gam10]. universum [Gam47a]. Unravelling [FR13]. unser [Gam69a]. unusual [MF69]. uranium [HS39]. Urans [HS39]. URCA [GN00]. using [Gam63b]. utforskar [Gam46c].

Vanished [Gam67a]. Varenna [MR86]. Variability [Gam67j, Dir72]. Variables [GL50]. vary [Alp73]. Velocity [GR31]. Vergaan [Hun49]. Verhalten [HS39]. Verhältnis [PG27]. Verlag [Gam51b, Uns60]. vers [Gam55d]. verständlich [GT56, GT58]. Verstandliche [Gam51a, Gam51g]. Very [BC05, Hoy54]. View [Wei77, Wei93, Gam38g]. Views [Gam56b]. Viking [Cas12a, Sit64b, Wil71]. violet [PG27]. violetten [PG27]. Visual [Ano50b]. Void [GHJ47]. Vol [Gam51b]. Volume [LT56, Rei72a, Rei72b, UM86a, UM86b]. vs [Gam1 , Gam42h]. Vselennoi [ZN73]. vue [Gam38g].

W [GHJ47, Gam54h, Rog62, Uns60]. Waerden [Kuh67]. Walk [Web73]. Walter [Ano47, Det55, Gam51f, Gam51b]. war [MW88, Rig95]. Wars [Stu18]. was [Wei13]. Washington [GF42, HPA97b, TGF39, TGF41, Anoxx]. Washton [Ske54]. Watershed [Gin94]. Watson [Det55, Jud01]. Watt [Sch12b]. Wave [GC28, Gam26, GI26]. waves [Lon72]. Way [Ske54]. WC [Gam43b]. Weber [Igg66]. Weiss [Igg66]. Weizel [Gam51b, Gam51f]. Weizsäcker [GH45]. Wellenmechanik [Gam51a, Gam51g]. Wellentheorie [GI26]. were [Bey49]. West [Gam51b]. Whitehead [Nug54]. Who [Gam63d]. Whyte [Gam49f]. Will [Gam51h]. William [Igg66]. Wissenschaft [GT56, GT58]. without [Sal52]. WN [Gam43b]. Wonderland [Gam39c, Gam42g, Fre40, M.40b, McC40]. words [Che94b]. work [AH90, Coc46, Hou30, Rac35]. World [AH71, Cer05, Fre94a, Fre94c, GIL28, Gam53g, Gam66c, Gam70, Gam93c, GIL02b, Gre00, Hob02, Kuh67, Oku02, Per03, Ric71, Ske54, Sta99, Stu71, Stu18, Tuc72, Wil71, Ano02, Del72, FF91, Haw11, SG12, Nug54]. World-Line [Fre94a]. Worlds [Fre40, Gam42f]. writer [Sab96]. Writing [Rog62]. wrong [Tur08]. Wyler [Ske54].

X [Dys02]. xiv [Sit64b]. xxi [Cas12a]. xxx [Gam57b].

Yadra [Gam32c]. Yadro [Gam30a]. Ycas [Sus69]. Year [FR13, Rya05, Coc46]. Years [Azi67, Gam46a, Gam47b, Gam66e, Gam66f, Gam72, Gam75b, Gam85, Gam11a, Her66, Kle66, Kuh67, Mla98, AH96, Gam66g, Gam68f, Gam68e, Gam01b, Gam01a]. Ylem [Ano54]. York [Ano55a, Cas12a, Joh54a, Sit64b, Smi61a, Wil71]. young [Ber68, Gam60].

Zerfalls [Gam31b, Gam29b, Gam34a, Gam37b]. Zertrümmerung [Gam31b]. Zon [Hun49]. Zum [Gam60]. zur [Gam29b, Kuh67, Rac35, GI26, Gam28b, Gam29e, GH29, vW35]. Zusammenfassender [Gam38c]. zwei [GT56, GT58].

References

Alpher:1948:OCE

[ABG48] Ralph A. Alpher, Hans Bethe, and George Gamow. The origin of chemical elements. Physical Review, 73(7):803–804, April 1, 1948. CODEN PHRVAO. ISSN 0031-899X (print), 1536-

6065 (electronic). URL http://prola.aps.org/abstract/PR/

v73/i7/p803_1. See historical retrospective [Tur08] and comments [AWCT09].

Abel:2002:FFS

[ABN02] Tom Abel, Greg L. Bryan, and Michael L. Norman. The formation of the first star in the universe. Science, 295(5552):93–98, 2002. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (elec-

tronic). URL http://www.sciencemag.org/content/295/5552/
93.abstract.

Alpher:1953:PCI

[AFH53] Ralph A. Alpher, James W. Follin, Jr., and Robert C. Herman. Physical conditions in the initial stages of the expanding universe. Physical Review, 92(6):1347–1361, December 1953. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL

http://link.aps.org/doi/10.1103/PhysRev.92.1347.

Alpher:1968:PRB

[AG68] Ralph A. Alpher and George Gamow. A possible relation between cosmological quantities and the characteristics of elementary particles. Proceedings of the National Academy of Sciences of the United States of America, 61(2):363–366, October 15, 1968. CO-DEN PNASA6. ISSN 0027-8424 (print), 1091-6490 (electronic).

URL http://adsabs.harvard.edu/abs/1968PNAS...61..363A;
 http://www.jstor.org/stable/pdfplus/59081.pdf.

Alpher:1967:TCR

[AGH67] Ralph A. Alpher, George Gamow, and Robert Herman. Thermal cosmic radiation and the formation of protogalaxies. Proceedings of the National Academy of Sciences of the United States of America, 58(6):2179–2186, December 15, 1967. CO-DEN PNASA6. ISSN 0027-8424 (print), 1091-6490 (electronic).

URL http://adsabs.harvard.edu/abs/1967PNAS...58.2179A;
 http://www.jstor.org/stable/pdfplus/58699.pdf.

Alpher:1948:RAE

[AH48] Ralph A. Alpher and Robert C. Herman. On the relative abundance of the elements. Physical Review, 74(12):1737–1742, December 1948. CODEN PHRVAO. ISSN 0031-899X (print), 1536-

6065 (electronic). URL http://link.aps.org/doi/10.1103/
PhysRev.74.1737.

Alpher:1971:BRG

[AH71] Ralph A. Alpher and Robert Herman. Book review: George Gamow, My World Line: An Informal Autobiography. Physics Today, 24(3):51–52, March 1971. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://www.physicstoday.

org/resource/1/phtoad/v24/i3/p51_s1.

Alpher:1972:MG

[AH72a] Ralph A. Alpher and Robert Herman. Memories of Gamow. In Reines [Rei72a], pages 304–313. ISBN 0-87081-025-1. LCCN

QC780 .C65. URL http://adsabs.harvard.edu/abs/1972cht..
conf..304A.

Alpher:1972:RBB

[AH72b] Ralph A. Alpher and Robert Herman. Reflections on “big bang” cosmology. In Reines [Rei72a], pages 1–14. ISBN 0-87081-025-

1. LCCN QC780 .C65. URL http://adsabs.harvard.edu/abs/
1972cht..conf....1A.

Alpher:1973:ABG

[AH73] R. A. Alpher and R. Herman. Alpher, Bethe and Gamow. In Weber [Web73], page ?? ISBN 0-85498-027-X. LCCN Q167.W42. Compiled by R. L. Weber, edited by E. Mendoza, with a foreword by William Cooper.

Alpher:1990:EWB

[AH90] Ralph A. Alpher and Robert Herman. Early work on ‘big-bang’ cosmology and the cosmic blackbody radiation. In Bertotti et al. [BBBM90], pages 129–158. ISBN 0-521-37213-5. LCCN QB981

.M774 1990. URL http://www.loc.gov/catdir/description/
cam024/90041803.html;   http://www.loc.gov/catdir/toc/
cam028/90041803.html.

Alpher:1996:CGB

[AH96] Ralph A. Alpher and Robert Herman. Celebration of Gamow’s birth: 90 years later. Astronomical and Astrophysical Transactions, 10:1–2, 1996. CODEN AATREG. ISSN 1055-6796 (print),

1476-3540 (electronic). URL http://adsabs.harvard.edu/abs/
1996A%26AT...10....1A.

Alpher:2001:GBB

[AH01] Ralph Alpher and Robert Herman. Genesis of the Big Bang. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2001. ISBN 0-19-511182-6. 256 pp. LCCN ????

Alpher:1948:TRE

[AHG48] Ralph A. Alpher, Robert Herman, and George A. Gamow. Thermonuclear reactions in the expanding universe. Physical Review, 74(9):1198–1199, November 1, 1948. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

URL http://adsabs.harvard.edu/abs/1948PhRv...74.1198A;
 http://prola.aps.org/abstract/PR/v74/i9/p1198_2. See er-

ratum [AHG49a].

Alpher:1949:ETR

[AHG49a] Ralph A. Alpher, Robert Herman, and George A. Gamow. Erratum: Thermonuclear reactions in the expanding universe. Physical Review, 75(4):701, February 15, 1949. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http:/

/adsabs.harvard.edu/abs/1949PhRv...75..701A;   http://
prola.aps.org/abstract/PR/v75/i4/p701_1. See [AHG48].

Alpher:1949:OE

[AHG49b] Ralph A. Alpher, Robert C. Herman, and George A. Gamow. On the origin of the elements. Physical Review, 75(2):332, January 15, 1949. CODEN PHRVAO. ISSN 0031-899X (print), 1536-

6065 (electronic). URL http://prola.aps.org/abstract/PR/
v75/i2/p325_1.

Alpher:1948:NCT

[Alp48] Ralph A. Alpher. A neutron-capture theory of the formation and relative abundance of the elements. Physical Review, 74(11): 1577–1589, December 1948. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://link.aps.org/doi/

10.1103/PhysRev.74.1577.

Alpher:1973:LNC

[Alp73] Ralph A. Alpher. Large numbers, cosmology and Gamow: Are the fundamental constants of nature truly constant, or do they vary with time? An account of George Gamow’s last scientific enquiry. American Scientist, 61(1):52–58, January/February 1973. CO-DEN AMSCAC. ISSN 0003-0996 (print), 1545-2786 (electronic).

URL http://adsabs.harvard.edu/abs/1973AmSci..61...52A;
 http://www.jstor.org/stable/27843563.

Alpher:2012:RAR

[Alp12] Victor S. Alpher. Ralph A. Alpher, Robert C. Herman, and the Cosmic Microwave Background Radiation. Physics in Perspective (PIP), 14(3):300–334, September 2012. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://

link.springer.com/article/10.1007/s00016-012-0088-7.

Alpher:2023:GGR

[AM23] Victor S. Alpher and Simon A. Mitton. George Gamow and Ralph Alpher: a review of their cosmological collaboration as mentor and protégé 1942–1955. European Physical Journal H, 48(1):??, December 2023. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-

6467 (electronic). URL https://link.springer.com/article/
10.1140/epjh/s13129-023-00057-0.

Anonymous:1940:BRB

[Ano40] Anonymous. Book reviews: The Birth and Death of the Sun, by G. Gamow. Popular Astronomy, 48:397–??, December 1940. CO-

DEN ???? ISSN ???? URL http://adsabs.harvard.edu/abs/
1940PA.....48..397G.

Anonymous:1944:BRB

[Ano44] Anonymous. Book review: Mr. Tompkins Explores the Atom, by G. Gamow. Popular Astronomy, 52:362–??, 1944. CODEN ????

ISSN ???? URL http://adsabs.harvard.edu/abs/1944PA....
.52..362G.

Anonymous:1947:RBO

[Ano47] Anonymous. Review: Public Opinion by Walter Lippmann; Patterns of Culture by Ruth Benedict; The Birth and Death of the Sun by George Gamow; You and Music by Christian Darnton. Journal of Educational Sociology (New York), 20(5):314–315, January 1947. CODEN ???? ISSN 0885-3525 (print), 2326-4055 (electronic).

URL http://www.jstor.org/stable/pdfplus/2263833.pdf.

Anonymous:1950:BRB

[Ano50a] Anonymous. Book Review: Biography of the Earth, by George Gamow. Popular Astronomy, 58:367–??, October 1950. CO-

DEN ???? ISSN ???? URL http://adsabs.harvard.edu/abs/
1950PA.....58..367G.

Anonymous:1950:VMG

[Ano50b] Anonymous. Visual materials from the George Gamow and Barbara Gamow papers, 1950.

Anonymous:1954:AYW

[Ano54] Anonymous. After Ylem, what? Physics Today, 7(3):9, March 1954. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699

(electronic). URL http://www.physicstoday.org/resource/1/

PHTOAD/v7/i3. See also [AFH53, Gam52c].

Anonymous:1955:GGB

[Ano55a] Anonymous. Gamow, George. The moon. New York (20 E. 70th Street) : Henry Schuman, Inc. 1954. 118 P. $2.50. Science Education, 39(3):252, 1955. CODEN SEDUAV. ISSN 0036-8326 (print),

1098-237X (electronic). URL http://adsabs.harvard.edu/abs/
1955SciEd..39Q.252..

Anonymous:1955:NAS

[Ano55b] Anonymous. The new astronomy : a Scientific American book. Simon and Schuster, New York, NY, USA, 1955. x + 243 pp. LCCN ????

Anonymous:1956:KPP

[Ano56] Anonymous. The Kalinga Prize: Prof. George Gamow. Nature, 178 (4538):836, October 20, 1956. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v178/n4538/pdf/178836a0.pdf.

Anonymous:1968:OPG

[Ano68] Anonymous. Obituary: Professor George Gamow. Nature, 220 (5168):723, November 16, 1968. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.

com/nature/journal/v220/n5168/pdf/220723a0.pdf.

Anonymous:1969:GGD

[Ano69] Anonymous. G. Gamow, died 1968 Aug. 19. Ríse Hvezd, 50:18–??,

1969. ISSN 0035-5550. URL http://adsabs.harvard.edu/abs/
1969Rise...50...18..

Anonymous:1994:EOL

[Ano94] Anonymous. Errata: The Origin of the Liquid-Drop Model and the Interpretation of Nuclear Fission. Perspectives on Science, 2 (2):254, Summer 1994. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (electronic). Correction of printers error in bottom three equations from page 90.

Anonymous:1995:GGA

[Ano95] Anonymous. George Gamov and astrophysics. In A. M. Bykov, R. A. Chevalier, and D. G. Yakovlev, editors, Proceedings of the Jubilee Gamov Seminar at St. Petersburg, September 12–14, 1994, to commemorate the 90th anniversary of the birth of George Gamow. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1995. URL http://adsabs.harvard.

edu/abs/1995gga..conf.....B.

Anonymous:1998:BRB

[Ano98] Anonymous. Book review: The George Gamow Symposium / Astronomical Society of the Pacific, 1997. The Observatory, 118 (1146):311–??, October 1998. CODEN OBSEAR. ISSN 0029-7704.

URL http://adsabs.harvard.edu/abs/1998Obs...118..311H;
 http://www.ulo.ucl.ac.uk/obsmag/.

Anonymous:1999:CM

[Ano99] Anonymous. The common man. Physics in Perspective (PIP), 1 (3):336, October 1999. CODEN PHPEF2. ISSN 1422-6944 (print),

1422-6960 (electronic).  URL http://link.springer.com/
article/10.1007/s000160050024;   http://www.springerlink.
com/content/y1h02gr3rv7wurrl/.

Anonymous:2000:GG

[Ano00] Anonymous. Gamow, George (1904–68). In P. Murdin, editor, Encyclopedia of Astronomy and Astrophysics. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, Novem-

ber 2000. URL http://adsabs.harvard.edu/abs/2000eaa..
bookE3639..

Anonymous:2002:BRB

[Ano02] Anonymous. Book Review: The new world of Mr. Tompkins / Cambridge University Press, 2001. Monthly Notes of the Astronomical Society of South Africa, 61:78–??, June 2002. CODEN ???? ISSN 0024-8266 (print), 2224-8366 (electronic). URL http:

//adsabs.harvard.edu/abs/2002MNSSA..61...78G.

Anonymous:2005:CLB

[Ano05] Anonymous. Critique de livre: Il était sept Fois La révolution: Albert Einstein et les autres. La Recherche, ??(386):11–??, May 1, 2005. CODEN RCCHBV. ISSN 0029-5671 (print), 1625-9955 (elec-

tronic). URL http://www.larecherche.fr/idees/livres/il-
etait-sept-fois-revolution-albert-einstein-autres-01-05-
2005-73924.

Anonymous:20xx:WCT

[Anoxx] Anonymous. Washington conferences on theoretical physics. Web

document, 20xx. URL http://home.gwu.edu/~kargaltsev/HEA/
washington-conferences.html. Undated. The page includes a

photograph of a plaque with the preface “The most famous event at this 5th Washington Conference on Theoretical Physics came from the announcement by Niels Bohr at the 1939 conference, in the Hall of Government, Room 209, that the nucleus of uranium had been split by bombardment with neutrons, with significant energy released. This was the dawn of the atomic age.” and the engraving: “In this room, January 26, 1939, Niels Bohr made the first pubic announcement of the successful disintegration of uranium into barium with the attendant release of approximately two hundred million electron volts of energy per disintegration. This announcement was heard by the physicists listed below who where attending the fifth of the conferences on theoretical physics which are sponsored jointly by the Carnegie Institution of Washington and The George Washington University.” The participant listed on the plaque are: L. H. Adams; Donald Hatch Andrews; Ferdinand G. Brickwedde; Gerhard Heinrich Dieke; George A. Gamow; Maria Goeppert-Mayer; M. H. Hebb; Karl Ferdinand Herzfeld; J. H. Hibben; J. H. Hoge; D. R. Inglis; F. G. Keyes; F. C. Kracek; R. Myers; H. M. O’Bryan; E. Posnjak; A. E. Ruark; R. B. Scott; Francis B. Silsbee; C. Starr; Otto Stern; Edward Teller; Harold C. Urey; and B. D. van Evera.

Atwood:1954:RBT

[Atw54] K. C. Atwood. Review: Mr. Tompkins Learns the Facts of Life by George Gamow. Quarterly Review of Biology, 29(4):343–344, December 1954. CODEN QRBIAK. ISSN 0033-5770 (print),

1539-7718 (electronic). URL http://www.jstor.org/stable/
pdfplus/2815365.pdf.

Amett:2009:HHN

[AWCT09] David Amett, George Wallerstein, Ken Croswell, and Michael S. Turner. αβγ, Hoyle, and the history of nucleosynthesis. Physics Today, 62(5):10–11, May 2009. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://www.

physicstoday.org/resource/1/phtoad/v62/i5/p10_s1. See

[ABG48, Tur08].

Azimov:1967:BGG

[Azi67] Isaac Azimov. Books: George Gamow, Thirty Years That Shook Physics: The Story of Quantum Theory. The Physics Teacher, 5 (7):350–351, 1967. CODEN PHTEAH. ISSN 0031-921X (print),

1943-4928 (electronic). URL http://link.aip.org/link/?PTE/
5/350/1.

Badash:1971:IBE

[Bad71] Lawrence Badash. The importance of being Ernest Rutherford. Science, 173(4000):873, September 3, 1971. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http:/

/adsabs.harvard.edu/abs/1971Sci...173..873B;   http://
www.jstor.org/stable/1731789.

Baitsell:1953:SP

[Bai53] George A. Baitsell, editor. Science in Progress. Yale University Press, New Haven, CT, USA, 1953. xiv + 285 pp. LCCN ????

Barr:1953:BRG

[Bar53] E. Scott Barr. Book review: George Gamow, Mr. Tompkins Learns the Facts of Life. American Journal of Physics, 21(8):649– 650, November 1953. CODEN AJPIAS. ISSN 0002-9505 (print),

1943-2909 (electronic). URL http://ajp.aapt.org/resource/1/
ajpias/v21/i8/p649_s1.

Bethe:1936:NPS

[BB36] H. A. Bethe and R. F. Bacher. Nuclear physics A. Stationary states of nuclei. Reviews of Modern Physics, 8(2):82–229, April 1936. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.8.82;   http://rmp.aps.org/abstract/RMP/v8/
i2/p82_1.

Bertotti:1990:MCR

[BBBM90] B. Bertotti, R. Balbinot, S. Bergia, and A. Messina, editors. Modern Cosmology in Retrospect. Cambridge University Press, Cambridge, UK, 1990. ISBN 0-521-37213-5. xx + 426 pp. LCCN QB981

.M774 1990. URL http://www.loc.gov/catdir/description/
cam024/90041803.html;   http://www.loc.gov/catdir/toc/
cam028/90041803.html.

Bernardini:2007:CGR

[BBC^{+}07] M. G. Bernardini, C. L. Bianco, P. Chardonnet, F. Fraschetti, R. Ruffini, and S.-S. Xue. Cosmic gamma ray bursts: the brightest stars of the Universe. In memory of George Gamow: a bright physicist. In Bisnovatyĭ-Kogan et al. [BKST^{+}07], pages 153–?? ISBN 1-904868-38-X. LCCN QB460 .G356. URL http://adsabs.

harvard.edu/abs/2007acag.conf..153B.

Burbidge:1957:SES

[BBFH57] E. Margaret Burbidge, G. R. Burbidge, William A. Fowler, and F. Hoyle. Synthesis of the elements in stars. Reviews of Modern Physics, 29(4):547–650, October 1957. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.29.547;
 http://rmp.aps.org/abstract/RMP/v29/i4/p547_1.

Beers:2005:DAV

[BC05] Timothy C. Beers and Norbert Christlieb. The discovery and analysis of very metal-poor stars in the galaxy. Annual Review of Astronomy and Astrophysics, 43(??):531–580, September 2005. CODEN ARAAAJ. ISSN 0066-4146 (print), 1545-

4282 (electronic). URL http://www.annualreviews.org/doi/
abs/10.1146/annurev.astro.42.053102.134057.

Bykov:1995:GGA

[BCY95] A. M. Bykov, R. A. Chevalier, and D. G. Yakovlev. George Gamov and astrophysics. Proceedings. Jubilee Gamow Seminar, St. Petersburg (Russia), 12–14 Sep 1994. Space Science Reviews, 74: ??, November 1995. CODEN SPSRA4. ISSN 0038-6308 (print),

1572-9672 (electronic). URL http://adsabs.harvard.edu/abs/
1995SSRv...74.....B.

Bekenstein:1986:FSC

[Bek86] Jacob D. Bekenstein. The fine-structure constant: From Eddington’s time to our own. In Ullmann-Margalit [UM86b], pages 209– 224. ISBN 90-277-2160-2, 90-277-2161-0 (paperback), 94-009-4566-3 (e-book). ISSN 0068-0346. LCCN Q174 .B67 vol. 95 Q175. URL

http://www.springerlink.com/content/978-94-009-4566-1;
 https://link.springer.com/book/10.1007/978-94-009-4566-
1.

Berger:1968:BYP

[Ber68] Carl F. Berger. Books for young people: George Gamow, Mr. Tompkins Returns and Mr. Tompkins in Paperback. The Physics Teacher, 6(9):473–474, 1968. CODEN PHTEAH. ISSN 0031-921X (print), 1943-4928 (electronic). URL http://link.aip.org/

link/?PTE/6/473/1.

Bethe:1939:EPSa

[Bet39] Hans A. Bethe. Energy production in stars. Physical Review, 55 (5):434–456, March 1, 1939. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://link.aps.org/doi/

10.1103/PhysRev.55.434;   http://prola.aps.org/abstract/

PR/v55/i5/p434_1. See further work [Hoy46, Hoy54]. According to [BL09, page 41], this paper won Bethe the A. Cressy Morrison Prize of the New York Academy of Sciences, and almost three decades later, the 1967 Nobel Prize in Physics.

Bethe:1997:IGE

[Bet97] Hans Bethe. Influence of Gamow on early astrophysics and on early accelerators in nuclear physics. In Harper et al. [HPA97b], pages 44–48. ISBN 1-886733-49-X. LCCN QC16.G37 G46 1996.

URL http://adsabs.harvard.edu/abs/1997ASPC..129...44B.

Beyer:1949:FNP

[Bey49] Robert T. (Robert Thomas) Beyer, editor. Foundations of nuclear physics: facsimiles of thirteen fundamental studies as they were originally reported in the scientific journals. Dover, New York, NY, USA, 1949. 272 pp. LCCN QC173 .B485. See original paper [HS39] and later translation [Gra64].

Bernstein:1986:CCP

[BF86] Jeremy Bernstein and Gerald Feinberg, editors. Cosmological Constants: Papers in Modern Cosmology. Columbia University Press, New York, NY, USA, 1986. ISBN 0-231-06376-8. ix + 328 pp. LCCN QB985 .C66 1986.

Bloch:1936:PRE

[BG36] Felix Bloch and George Gamow. On the probability of γ-ray emission. Physical Review, 50(3):260, August 1, 1936. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL

http://prola.aps.org/abstract/PR/v50/i3/p260_1.

Brittin:1961:NEP

[BG61] Wesley Brittin and George Gamow. Negative entropy and photosynthesis. Proceedings of the National Academy of Sciences of the United States of America, 47(5):724–727, May 15, 1961. CODEN PNASA6. ISSN 0027-8424 (print), 1091-6490 (electronic). URL

http://www.jstor.org/stable/pdfplus/70982.pdf.

Belzer:1950:DEG

[BGK50] Jack Belzer, George Gamow, and Geoffrey Keller. Dynamics of elliptical galaxies. In Proceedings, Scientific Computation Forum, pages 67–69. International Business Machines Corporation, New York, NY, USA, 1950.

Belzer:1951:SDS

[BGK51] J. Belzer, G. Gamow, and G. Keller. On the stellar dynamics of spherical galaxies. Astrophysical Journal, 113:166–180, January 1951. CODEN ASJOAB. ISSN 0004-637X (print), 1538-4357

(electronic). URL http://adsabs.harvard.edu/abs/1951ApJ..
.113..166B.

Binnie:1958:RBE

[Bin58] W. P. Binnie. Review: Matter, Earth, and Sky by George Gamow. Science, 128(3324):587–588, September 12, 1958. CODEN SCI-EAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http:

//www.jstor.org/stable/pdfplus/1754727.pdf.

Bagdonas:2021:FOB

[BK21] Alexandre Bagdonas and Alexei Kojevnikov. Funny origins of the Big Bang Theory. Historical Studies in the Natural Sciences, 51(1):87–??, February 2021. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic). URL https://online.

ucpress.edu/hsns/article/51/1/87/116291/Funny-Origins-
of-the-Big-Bang-Theory.

Bisnovatyi-Kogan:2007:ACA

[BKST^{+}07] G. S. (Gennadiĭ Semenovich) Bisnovatyĭ-Kogan, S. Silich, E. Terlevich, R. Terlevich, and A. Zhuk, editors. Astrophysics and cosmology after Gamow: proceedings of the Gamow Memorial International Conference. Cambridge Scientific Publishers, Cottenham, Cambridge, 2007. ISBN 1-904868-38-X. LCCN QB460 .G356. URL

http://adsabs.harvard.edu/abs/2007acag.conf.....B.

Brown:2009:HAB

[BL09] Gerald E. Brown and Sabine Lee. Hans Albrecht Bethe: July 2, 1906–March 6, 2005. Biographical memoirs — National Academy of Sciences of the United States of America, 91:31–56, 2009. CO-DEN BMNSAC. ISBN 0-309-14560-0. ISSN 0077-2933. URL http:

//www.nap.edu/catalog.php?record_id=12776.

Bloom:1988:PGT

[Blo88] Stewart D. Bloom. On the prevalence of the Gamow–Teller transition. In Mark and Wood [MW88], pages 15–37. ISBN 90-277-2775-9. LCCN QC71 .E62 1988.

Boyle:1993:BRB

[Boy93] B. Boyle. Book Review: Mr. Tompkins in paperback / Cambridge U Press, 1993. The Observatory, 113(1117):318–??, December 1993. CODEN OBSEAR. ISSN 0029-7704. URL http://adsabs.

harvard.edu/abs/1993Obs...113Q.318B;   http://www.ulo.
ucl.ac.uk/obsmag/.

Beckenbach:1964:ACM

[BP64] Edwin F. Beckenbach and George Pólya, editors. Applied Combinatorial Mathematics. Wiley, New York, NY, USA, 1964. xxi + 608 pp. LCCN QA164 .B4 1964.

Beckenbach:1981:ACM

[BP81] Edwin F. Beckenbach and George Pólya, editors. Applied Combinatorial Mathematics. R. E. Krieger Pub. Co., Malabar, FL, USA, 1981. ISBN 0-89874-172-6. xxi + 608 pp. LCCN QA164 .B4 1981.

Baaz:2011:KGF

[BPP^{+}11] Matthias Baaz, Christos H. Papadimitriou, Hilary W. Putnam, Dana S. Scott, and Charles L. Harper, Jr., editors. Kurt Gödel and the Foundations of Mathematics: Horizons of Truth. Cambridge University Press, Cambridge, UK, 2011. ISBN 0-521-76144-1 (print), 0-511-97423-X (e-book). xxiv + 515 pp. LCCN QA9.65 .K87 2011.

Binnig:1985:STM

[BR85] Gerd Binnig and Heinrich Rohrer. The scanning tunneling microscope. Scientific American, 253(2):50–56, August 1985. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v253/
n2/pdf/scientificamerican0885-50.pdf;   http://www.nobelprize.
org/nobel_prizes/physics/laureates/1986/.

Brenner:1957:IAO

[Bre57] S. Brenner. The impossibility of all overlapping triplet codes in information transfer from nucleic acid to proteins. Proceedings of the National Academy of Sciences of the United States of America, 43 (8):687–694, August 15, 1957. CODEN PNASA6. ISSN 0027-8424 (print), 1091-6490 (electronic). URL http://www.jstor.

org/stable/89586.

C:1948:RBE

[C.48] M. S. C. Review: Biography of the Earth by George Gamow. The Journal of Geology, 56(6):593–594, November 1948. CODEN JGEOAZ. ISSN 0022-1376 (print), 1537-5269 (electronic). URL

http://www.jstor.org/stable/pdfplus/30080854.pdf.

Cassidy:2012:BRG

[Cas12a] David C. Cassidy. Book review: Gino Segrè, Ordinary Geniuses: Max Delbrück, George Gamow, and the Origins of Genomics and Big Bang Cosmology, Viking Adult, New York, 2011, xxi + 330 pages, ISBN-13 978-0-670-02276-2. Physics in Perspective (PIP), 14(2):248–250, June 2012. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Cassidy:2012:OGM

[Cas12b] David C. Cassidy. Ordinary geniuses: Max Delbrück, George Gamow, and the origins of genomics and Big Bang cosmology. Physics in Perspective (PIP), 14(2):248–250, June 2012. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Chakrabarti:2009:ACA

[CBKZ^{+}09] Sandip K. Chakrabarti, G. S. (Gennadiĭ Semenovich) Bisnovatyĭ-Kogan, Alexander I. Zhuk, et al., editors. Astrophysics and cosmology after Gamow: proceedings of the 4th Gamow International Conference on Astrophysics and Cosmology after Gamow and The 9th Gamow Summer School, Odessa, Ukraine, 17–23 August 2009, volume 1206 of AIP conference proceedings. American Institute of Physics, Woodbury, NY, USA, 2009. ISBN 0-7354-0743-6 (paperback). ISSN 0094-243X (print), 1551-7616 (electronic), 1935-0465. LCCN QB460 .G37 2009.

URL http://adsabs.harvard.edu/abs/2010AIPC.1206.....
C;   http://scitation.aip.org/dbt/dbt.jsp?KEY=APCPCS\%
26Volume=1206\%26Issue=1.

Cockcroft:1934:SPN

[CCJ^{+}34] J. Cockcroft, J. Chadwick, F. Joliot, J. Joliot, N. Bohr, G. Gamov, P. A. M. Dirac, and W. Heisenberg, editors. Structure et propriétés des noyaux atomiques. Rapports et discussions du septième conseil de physique tenu à Bruxelles du 22 au 29 octobre 1933 sous les auspices de l’institut international de physique Solvay. (French) [Structure and properties of atomic nuclei. Reports and discussions of the Seventh Meeting on Physics held in Brussels from 22 to 29 October 1933 under the auspices of the Solvay International Institute of Physics]. Gauthier-Villars et cie, Paris, France, 1934. LCCN ???? Publiés par la commission administrative de l’institut.

Cernobai:2005:GGC

[Cer05] V. Cernobai. G. Gamow and calculability problems of the world constants and the life code. Odessa Astronomical Publications = Izvestija Odesskoj Astronomičeskoj Observatorii, 17:21–24, 2005. CODEN ???? ISSN 1810-4215. URL http://adsabs.harvard.

edu/abs/2005OAP....17...21C.

Chadwick:1930:ADP

[CG30] J. Chadwick and George Gamow. Artificial disintegration by αparticles. Nature, 126(3167):54–55, July 12, 1930. CODEN NAT-UAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Critchfield:1939:SSS

[CG39] C. L. Critchfield and G. Gamow. The shell-source stellar model. Astrophysical Journal, 89:244–254, March 1939. CODEN ASJOAB. ISSN 0004-637X (print), 1538-4357 (electronic). URL

http://adsabs.harvard.edu/abs/1939ApJ....89..244C.

Chandrasekhar:1938:PSE

[CGT38] Subrahmanyan Chandrasekhar, George Gamow, and Merle A. Tuve. The problem of stellar energy. Nature, 141(3578):982, May 28, 1938. CODEN NATUAS. ISSN 0028-0836 (print),

1476-4687 (electronic). URL http://adsabs.harvard.edu/abs/
1938Natur.141..982C.

Chernin:1994:GAN

[Che94a] Artur D. Chernin. Gamow in America: 1934–1968 (on the ninetieth anniversary of G A Gamov’s birth). Physics-Uspekhi, 37 (8):791–801, 1994. CODEN PHUSEY. ISSN 1063-7869 (print),

1468-4780 (electronic). URL http://adsabs.harvard.edu/
abs/1994PhyU...37..791C;   http://stacks.iop.org/1063-
7869/37/i=8/a=A04.

Chernin:1994:HGC

[Che94b] Artur D. Chernin. How Gamow calculated the temperature of the background radiation or a few words about the fine art of theoretical physics. Physics-Uspekhi, 37(8):813–820, 1994. CO-DEN PHUSEY. ISSN 1063-7869 (print), 1468-4780 (electronic).

URL http://adsabs.harvard.edu/abs/1994PhyU...37..813C;
 http://stacks.iop.org/1063-7869/37/i=8/a=A06.

Chernin:1995:GGB

[Che95] A. D. Chernin. George Gamow and the Big Bang. Space Science Reviews, 74(3–4):447–454, November 1995. CODEN SPSRA4. ISSN 0038-6308 (print), 1572-9672 (electronic). URL http://

adsabs.harvard.edu/abs/1995SSRv...74..447C.

Clayton:1968:PSE

[Cla68] Donald D. Clayton. Principles of stellar evolution and nucleosynthesis. McGraw-Hill, New York, NY, USA, 1968. xi + 612 pp. LCCN QB464 .C55.

Clayton:1983:PSE

[Cla83] Donald D. Clayton. Principles of stellar evolution and nucleosynthesis: with a new preface. University of Chicago Press, Chicago, IL, USA, 1983. ISBN 0-226-10953-4 (paperback), 0-226-10952-6 (hardcover). xxii + 612 pp. LCCN QB806

.C55 1983. URL http://www.loc.gov/catdir/enhancements/
fy0608/83005106-d.html;   http://www.loc.gov/catdir/enhancements/
fy0608/83005106-t.html;   http://www.loc.gov/catdir/enhancements/
fy0609/83005106-b.html.

Choksi:2017:OPR

[CMT17] Rustum Choksi, Cyrill B. Muratov, and Ihsan Topaloglu. An old problem resurfaces nonlocally: Gamow’s liquid drops inspire today’s research and applications. Notices of the American Mathematical Society, 64(11):1275–1283, December 2017. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Carroll:2007:IMA

[CO07] Bradley W. Carroll and Dale A. Ostlie. An Introduction to Modern Astrophysics. Pearson Addison-Wesley, San Francisco, CA, USA, second edition, 2007. ISBN 0-8053-0402-9, 0-321-44284-9 (paperback). ???? pp. LCCN QB461 .C35 2007. URL http:/

/catdir.loc.gov/catdir/toc/ecip0613/2006015391.html;
http://www.gbv.de/dms/ilmenau/toc/512485305.PDF.

Cockcroft:1946:RLW

[Coc46] J. D. Cockcroft. Rutherford: Life and work after the year 1919, with personal reminiscences of the Cambridge period. Proceedings of the Physical Society, London, 58(6):625–633, November 1946. CODEN PPSOAU. ISSN 0959-5309 (print), 2051-2171

(electronic). URL http://stacks.iop.org/0959-5309/58/i=6/
a=301.

Cowan:1972:NPP

[CR72] Clyde L. Cowan and Frederick Reines. Neutrino physics — prospects. In Reines [Rei72a], pages 150–168. ISBN 0-87081-025-

1. LCCN QC780 .C65. URL http://adsabs.harvard.edu/abs/
1972cht..conf..150C.

Critchfield:1972:AFT

[Cri72] Charles L. Critchfield. Analytic forms of the thermonuclear function. In Reines [Rei72a], pages 186–191. ISBN 0-87081-025-1.

LCCN QC780 .C65. URL http://adsabs.harvard.edu/abs/
1972cht..conf..186C.

Danos:1965:BRG

[Dan65] Michael Danos. Book review: George Gamow, Matter, Earth, and Sky. Physics Today, 18(11):63–65, November 1965. CO-DEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

URL http://www.physicstoday.org/resource/1/phtoad/v18/
i11/p63_s3.

Delbruck:1972:W

[Del72] Max Delbrück. Out of this world. In Reines [Rei72a], pages 280– 288. ISBN 0-87081-025-1. LCCN QC780 .C65.

Demiannski:2007:GGG

[Dem07] M. Demiański. George Gamow and the genetic code. In Bisnovatyĭ-Kogan et al. [BKST^{+}07], pages 65–?? ISBN 1-904868-38-X.

LCCN QB460 .G356. URL http://adsabs.harvard.edu/abs/
2007acag.conf...65D.

Dethier:1955:RBP

[Det55] V. G. Dethier. Review: Science in Progress, by George Gamow, Walter Orr Roberts, Joseph Kaplan, F. M. Carpenter, D. M. S. Watson, Donald R. Griffin, Paul R. Burkholder, E. Newton Harvey, Charles H. Townes, George Scatchard, George A. Baitsell. Quarterly Review of Biology, 30(4):429–??, December 1955. CO-DEN QRBIAK. ISSN 0033-5770 (print), 1539-7718 (electronic).

URL http://www.jstor.org/stable/pdfplus/2814309.pdf.

Delbruck:1931:UAK

[DG31] Max Delbrück and George Gamow. Übergangswahrscheinlichkeiten von angeregten Kernen. (German) [Transition probabilities of excited nuclei]. Zeitschrift für Physik, 72(7–8):492–499, July 1931. CODEN ZEPYAA. ISSN 0044-3328. URL http://www.

springerlink.com/content/jlt1564970867882/.

Dounce:1956:NRT

[DGS^{+}56] A. L. Dounce, G. Gamow, S. Spiegelman, P. Newmark, D. Harker, and M. Soodak. Nucleoproteins round-table discussion. Journal of Cellular and Comparative Physiology, 47(S1):103–112, May 1956. CODEN JCLLAX. ISSN 0021-9541 (print), 1097-4652 (electronic).

Dirac:1972:VGC

[Dir72] Paul A. M. Dirac. The variability of the gravitational constant. In Reines [Rei72a], pages 56–59. ISBN 0-87081-025-1. LCCN QC780

.C65. URL http://adsabs.harvard.edu/abs/1972cht..conf..
.56D.

Dixey:1961:RBE

[Dix61] F. Dixey. Review: Biography of the Earth: Its Past, Present and Future by George Gamow. The Geographical Journal, 127 (1):112–113, March 1961. CODEN GGJOAR. ISSN 0016-7398 (print), 1475-4959 (electronic). URL http://www.jstor.org/

stable/pdfplus/1793234.pdf.

DeToledo:1948:RAN

[DW48] Paulo Saraiva De Toledo and Gleb Wataghin. On the relative abundances of nuclei in the universe. Physical Review, 73(1):79–80, January 1948. CODEN PHRVAO. ISSN 0031-899X (print), 1536-

6065 (electronic). URL http://link.aps.org/doi/10.1103/

PhysRev.73.79.2. See remarks in [Dys93] about the relation of this work to [ABG48], and the subsequent incorrect neglect of Wataghin’s work. See also related papers [LW46, Wat46, Wat48].

Dwight:1966:BRG

[Dwi66] C. Harrison Dwight. Book review: George Gamow, Matter, Earth, and Sky. American Journal of Physics, 34(5):447–448, May 1966. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic).

URL http://ajp.aapt.org/resource/1/ajpias/v34/i5/p447_
s2.

Dyson:1987:BRB

[Dys87] J. Dyson. Book review: Gamow cosmology / North Holland Press, 1986. Astrophysics and Space Science, 138(1):227, November 1987. CODEN APSSBE. ISSN 0004-640X (print), 1572-946X (elec-

tronic). URL http://adsabs.harvard.edu/abs/1987Ap%26SS.
138..227D.

Dyson:1993:GGP

[Dys93] Freeman Dyson. George Green and physics. Physics World, 6 (8):33–38, August 1993. CODEN PHWOEW. ISSN 0953-8585 (print), 2058-7058 (electronic). URL iopscience.iop.org/pwa/

full/pwa-pdf/6/8/phwv6i8a28.pdf.

Dyson:2002:BRB

[Dys02] Freeman J. Dyson. Book review: Memoirs: A Twentieth-Century Journey in Science and Politics. Edward Teller with Judith Shoolery. 602 pp. Perseus, Cambridge, MA, 2001. Price: $35.00 ISBN 0-7382-0532-X. American Journal of Physics, 70(4):462–463, April 2002. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (elec-

tronic). URL http://ajp.aapt.org/resource/1/ajpias/v70/
i4/p462_s1.

Ehricke:1957:RMR

[EG57a] Krafft Ehricke and George Gamow. Mit einer Rakete um den Mond. [A rocket around the Moon]. Vasiona, 5(??):68–71, 1957. CODEN ???? ISSN 0506-4295.

Ehricke:1957:RAM

[EG57b] Krafft Ehricke and George Gamow. A rocket around the Moon. Scientific American, 196(6):47–53, June 1957. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (elec-

tronic). URL http://chelsea-sulis.blogspot.com/2011/08/
cow-jumped-over-moon-1957.html.

Fara:2001:GPI

[Far01] Patricia Fara. Group portraits IV — the Seventh Solvay Conference [Bruxelles, 22–29 October 1933]. Endeavour, 25(4):137–138, December 2001. CODEN ENDEAS. ISSN 0160-9327 (print), 1873-1929 (electronic).

Frebel:2012:CSF

[FB12a] Anna Frebel and Volker Bromm. Chemical signatures of the first galaxies: Criteria for one-shot enrichment. Astrophysical Journal, 759(2):115:1–115:8, November 10, 2012. CODEN ASJOAB. ISSN 0004-637X (print), 1538-4357 (electronic). URL http://stacks.

iop.org/0004-637X/759/i=2/a=115.

Frebel:2012:PFI

[FB12b] Anna Frebel and Volker Bromm. Precious fossils of the infant universe. Physics Today, 65(4):49–54, April 2012. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http:

//scitation.aip.org/content/aip/magazine/physicstoday/
article/65/4/10.1063/PT.3.1519.

Feather:1962:BRG

[Fea62] Norman Feather. Book review: George Gamow and John M. Cleveland, Physics: Foundations and Frontiers. Physics Today, 15(7):58, July 1962. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699

(electronic). URL http://www.physicstoday.org/resource/1/
phtoad/v15/i7/p58_s1.

Fermi:1968:III

[Fer68] Laura Fermi. Illustrious immigrants; the intellectual migration from Europe, 1930/41. University of Chicago Press, Chicago, IL, USA, 1968. xi + 440 pp. LCCN E184.A1 F47.

Fermi:1971:III

[Fer71] Laura Fermi. Illustrious immigrants; the intellectual migration from Europe, 1930/41. University of Chicago Press, Chicago, IL, USA, second edition, 1971. ISBN 0-226-24376-1, 0-226-24378-8 (paperback). xi + 431 pp. LCCN E184.A1 F47 1971.

Ferris:1991:WTP

[FF91] Timothy Ferris and Clifton Fadiman, editors. The world treasury of physics, astronomy, and mathematics. Little, Brown and Co., Boston, MA, USA, 1991. ISBN 0-316-28129-8. xv + 859 pp. LCCN QC71 .W67 1991. Foreword by Clifton Fadiman.

Field:1959:RBE

[Fie59] George B. Field. Review: Matter, Earth and Sky by George Gamow. American Scientist, 47(3):294A, ???? 1959. CODEN AMSCAC. ISSN 0003-0996 (print), 1545-2786 (electronic). URL

http://www.jstor.org/stable/27827408.

Frebel:2012:MPS

[FN12] Anna Frebel and John E. Norris. Metal-poor stars and the chemical enrichment of the universe. In Gilmore [Gil12], chapter 3. ISBN 94-007-5612-7. LCCN QB460-466; QB857 .P53 2013.

Fowler:1972:NOO

[Fow72] William A. Fowler. New observations and old nucleocosmochronologies. In Reines [Rei72a], pages 67–123. ISBN 0-87081-

025-1. LCCN QC780 .C65. URL http://adsabs.harvard.edu/
abs/1972cht..conf...67F.

Fernandez:2013:UMA

[FR13] Bernard Fernandez and Georges Ripka. Unravelling the Mystery of the Atomic Nucleus — a Sixty Year Journey 1896–1956. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2013. ISBN 1-4614-4180-3 (hardcover), 1-4614-4181-1 (ebook). xviii + 522 pp. LCCN QC773 .F47 2013.

Freeman:1940:RSA

[Fre40] Ira M. Freeman. Review: Strangest of all possible worlds: Mr Tompkins in Wonderland by George Gamow. The Scientific Monthly, 51(5):471–472, November 1940. CODEN SCMOAA. ISSN 0096-3771 (print), 2327-7513 (electronic). URL http://www.

jstor.org/stable/pdfplus/17405.pdf.

Freeman:1961:BRG

[Fre61] Ira M. Freeman. Book review: George Gamow and John M. Cleveland, Physics: Foundations and Frontiers. American Journal of Physics, 29(1):60, January 1961. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). URL http://ajp.aapt.org/

resource/1/ajpias/v29/i1/p60_s1.

Frenkel:1994:GGWa

[Fre94a] V. Y. Frenkel. George Gamow — world-line 1904–1933 — (on the occasion of the 90th anniversary of his birth). Uspekhi Fizicheskikh Nauk, 164(8):845–866, August 1994. CODEN UFNAAG. ISSN 0042-1294 (print), 1996-6652 (electronic).

Frenkel:1994:CBG

[Fre94b] Viktor Ya. Frenkel’. Correspondence between G A Gamow and P L Kapitza. Physics-Uspekhi, 37(8):803–811, 1994. CODEN PHUSEY. ISSN 1063-7869 (print), 1468-4780 (electronic). URL

http://stacks.iop.org/1063-7869/37/i=8/a=A05.

Frenkel:1994:GGWb

[Fre94c] Viktor Ya. Frenkel’. George Gamow: World line 1904–1933 (on the ninetieth anniversary of G A Gamov’s birth). Physics-Uspekhi, 37(8):767–789, 1994. CODEN PHUSEY. ISSN 1063-7869 (print), 1468-4780 (electronic). URL http://stacks.iop.org/

1063-7869/37/i=8/a=A03.

Frebel:2010:SAE

[Fre10] Anna Frebel. Stellar archaeology: Exploring the universe with metal-poor stars. Astronomische Nachrichten, 331(5):474–488, May 20, 2010. CODEN ASNAAN. ISSN 0004-6337 (print), 1521-

3994 (electronic). URL http://onlinelibrary.wiley.com/doi/
10.1002/asna.201011362/abstract.

Frebel:2014:RCE

[Fre14] Anna Frebel. Reconstructing the cosmic evolution of the chemical elements. Dædalus, 143(4):71–80, Fall 2014. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://www.

mitpressjournals.org/doi/pdfplus/10.1162/DAED_a_00307.

Graetzer:1971:DNF

[GA71] Hans G. Graetzer and David L. Anderson. The discovery of nuclear fission: a documentary history, volume 20 of Van Nostrand Reinhold momentum books. Van Nostrand Reinhold, New York, NY, USA, 1971. viii + 120 pp. LCCN QC790 .G68.

Gamow:1949:SSH

[Gam42] George Gamow. Sengo shuppan honyaku genshiryoku, butsurigaku kankei tosho. Hakuyosha, Tokyo, Japan, 1949–1967 (Showa 24–42). ???? (volumes 24) pp. LCCN CLC QD461 nos. 1-24.

Gamow:1926:TOP

[Gam26] G. Gamow. Sur la théorie des ondes de phases de M. L. de Broglie. (French) [On the phase-wave theory of M. L. de Broglie]. Comptes rendus de l’Académie des sciences, Paris, 183:875–876, November 13, 1926. CODEN ???? ISSN ????

Gamow:1927:PFO

[Gam27] G. A. Gamow. The principle of fundamental observability in modern physics. (Russian). Uspekhi Fizicheskikh Nauk, 5(??):386–??, ???? 1927. CODEN UFNAAG. ISSN 0042-1294 (print), 1996-6652 (electronic).

Gamow:1928:QTN

[Gam28a] George Gamow. The quantum theory of nuclear disintegration. Nature, 122(3082):805–806, November 24, 1928. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Gamow:1928:QAG

[Gam28b] George Gamow. Zur Quantentheorie des Atomkernes. (German) [On the quantum theory of the atomic nucleus]. Zeitschrift für Physik, 51(3–4):204–212, March 1928. CODEN ZEPYAA.

ISSN 0044-3328. URL http://www.springerlink.com/content/

mw52h8867mr4x185/. Reprinted in [Bey49, pages 77–85].

Gamow:1929:SAG

[Gam29a] G. Gamow. Über die Struktur des Atomkernes. (German) [On the structure of the atomic nucleus]. Physikalische Zeitschrift, 30: 717–720, 1929. CODEN PHZTAO. ISSN 0369-982X.

Gamow:1929:BQR

[Gam29b] George Gamow. Bemerkung zur Quantentheorie des radioaktiven Zerfalls. (German) [Note on the quantum theory of radioactive decay]. Zeitschrift für Physik, 53(7–8):601–604, July 1929. CODEN

ZEPYAA. ISSN 0044-3328. URL http://www.springerlink.
com/content/w211101j12115804/.

Gamow:1929:DSA

[Gam29c] George Gamow. Discussion on the structure of atomic nuclei. Proceedings of the Royal Society of London. Series A, Containing Papers of a Mathematical and Physical Character, 123(792):386–387, April 6, 1929. ISSN 0950-1207 (print), 2053-9150 (electronic). URL

http://www.jstor.org/stable/pdfplus/95202.pdf.

Gamow:1929:ST

[Gam29d] George Gamow. Successive α-transformations. Nature, 123(3103): 606, April 20, 1929. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Gamow:1929:QAG

[Gam29e] George Gamow. Zur Quantentheorie der Atomzertrümmerung. (German) [On the quantum theory of atomic fission]. Zeitschrift für Physik, 52(7–8):510–515, July 1929. CODEN ZEPYAA.

ISSN 0044-3328. URL http://www.springerlink.com/content/
t240444152t66876/.

Gamow:1930:AYR

[Gam30a] G. A. Gamow. Atomnoe Yadro i Radioaktivnost’ (Russian) [The Atomic Nucleus and Radioactivity]. Modern Trends in Scientific Thought. GIZ, Moscow and Leningrad, USSR, 1930. 80 pp. LCCN ????

Gamow:1930:FSR

[Gam30b] George Gamow. Fine structure of α-rays. Nature, 126(3176):397, September 13, 1930. CODEN NATUAS. ISSN 0028-0836 (print),

1476-4687 (electronic). URL http://www.nature.com/nature/
journal/v126/n3176/full/126397a0.html.

Gamow:1930:MDC

[Gam30c] George Gamow. Mass defect curve and nuclear constitution. Proceedings of the Royal Society of London. Series A, Containing Papers of a Mathematical and Physical Character, 126(803):632–644, March 3, 1930. ISSN 0950-1207 (print), 2053-9150 (electronic).

URL http://www.jstor.org/stable/pdfplus/95297.pdf.

Gamow:1931:CAN

[Gam31a] George Gamow. Constitution of Atomic Nuclei and Radioactivity. Clarendon Press, Oxford, UK, 1931. 4 + 114 + 1 pp. LCCN QC173 .G3.

Gamow:1931:TRZ

[Gam31b] George Gamow. Über die Theorie des radioaktiven Zerfalls, der Zertrümmerung und die Anregung durch Strahlen. (German) [On the theory of radioactive decay, the destruction and the excitation by radiation]. Physikalische Zeitschrift, 32(??):651–655, September 1, 1931. CODEN PHZTAO. ISSN 0369-982X.

Gamow:1932:NL

[Gam32a] G. Gamow. Nuclear α-and p-levels. Physikalische Zeitschrift der Sowjetunion, 1(3):433–435, 1932. CODEN PHZSAL. ISSN 0369-9811.

Gamow:1932:QTN

[Gam32b] G. Gamow. Quantum theory of nuclear structure. Reale Accademia d’Italia. Fondazione Alessandro Volta. Atti dei Convegni I: Convegno di Fisica nucleare, Ottobre 1931. Roma, 1:65–81, October 1932. CODEN ???? ISSN ???? May be the same as [Gam32i].

Gamow:1932:SAY

[Gam32c] G. A. Gamow. Stroenie Atomnogo Yadra i Radioaktivn ost’. (Russian) [The Constitution of the Atomic Nuclei and Radioactivity]. GTTI, Moscow and Leningrad, USSR, 1932. 146 pp. LCCN ????

Gamow:1932:EDT

[Gam32d] George Gamow. Essay on the development of the theory of atomic nucleus. (Russian). Uspekhi Fizicheskikh Nauk, 12(??):31–??, 1932. CODEN UFNAAG. ISSN 0042-1294 (print), 1996-6652 (electronic).

Gamow:1932:NAU

[Gam32e] George Gamow. A new attempt to understand the process of decay. (Russian). Sorena, ??(??):16–38, ???? 1932. CODEN ???? ISSN ????

Gamow:1932:ODS

[Gam32f] George Gamow. Outlines of the development of the studies on the atomic nucleus: Theory of radioactive decay (Russian). Physics-Uspekhi, 10(??):531–544, ???? 1932. CODEN PHUSEY. ISSN 1063-7869 (print), 1468-4780 (electronic).

Gamow:1932:RDN

[Gam32g] George Gamow. Radioactive disintegration and nuclear spin. Nature, 129(3256):470, March 26, 1932. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Gamow:1932:SAN

[Gam32h] George Gamow. The structure of the atomic nucleus and the transformation of the elements. Sorena, ??(??):16–38, ???? 1932. CO-DEN ???? ISSN ????

Gamow:1932:TQD

[Gam32i] George Gamow. Teoria quantica delta struttura nucleare. (Italian) [Quantum theory of nuclear structure]. Nuovo Cimento, 9(??): xxxii–xxxv, 1932. CODEN ???? ISSN ????

Gamow:1933:CRR

[Gam33a] George Gamow. Cosmic radiation. (Russian). Sorena, 1(??):36–56,

???? 1933. CODEN ???? ISSN ????

Gamow:1933:FSN

[Gam33b] George Gamow. Fundamental state of nuclear α-particles. Nature, 131(3313):618–619, April 29, 1933. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Gamow:1933:PEP

[Gam33c] George Gamow. Is the proton an elementary particle?. (Russian). Sorena, 9(??):105–??, ???? 1933. CODEN ???? ISSN ????

Gamow:1933:LRN

[Gam33d] George Gamow. L’origine des rayons et les niveaux d’énergie nucleaires. (French). [The origin of rays and nuclear energy levels]. Technical report, 98000 Institut Solvay (Physique), Brussels, Belgium, October 1933.

Gamow:1933:MED

[Gam33e] George Gamow. Mechanism of γ-excitation by β-disintegration. Nature, 131(3298):57–58, January 14, 1933. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Gamow:1933:NAT

[Gam33f] George Gamow. Neutrons and artificial transformation of elements. (Russian). Priroda (Moscow, Russian Federation) [Nature], 1(??): 16–21, ???? 1933. CODEN PRIRA3. ISSN 0032-874X.

Gamow:1933:NEL

[Gam33g] George Gamow. Nuclear energy levels. Nature, 131(3308):433, March 25, 1933. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Gamow:1933:FES

[Gam33h] George Gamow. On the formation of the elements in stars. (Russian). Uspekhi Astronomicheskich Nauk, 2(??):72–83, ???? 1933. CODEN ???? ISSN ????

Gamow:1933:ODT

[Gam33i] George Gamow. Outline of development of the theory of the structure of atomic nuclei. (Russian). Uspekhi Fizicheskikh Nauk, 13 (??):46–57, ???? 1933. CODEN UFNAAG. ISSN 0042-1294 (print), 1996-6652 (electronic).

Gamow:1933:PCR

[Gam33j] George Gamow. The problem of cosmic rays. (Russian). Priroda (Moscow, Russian Federation) [Nature], 3–4(??):36–39, ???? 1933. CODEN PRIRA3. ISSN 0032-874X.

Gamow:1933:TDE

[Gam33k] George Gamow. The theory of Dirac electrons and positive. Sorena, 8(??):25–30, ???? 1933. CODEN ???? ISSN ????

Gamow:1934:HSM

[Gam34a] G. Gamow. Über den heutigen Stand (20. Mai 1934) der Theorie des β-Zerfalls. (German) [On the current status (20 May 1934) of the theory of β decay]. Physikalische Zeitschrift, 35:533–542, 1934. CODEN PHZTAO. ISSN 0369-982X.

Gamow:1934:ARE

[Gam34b] George Gamow. Artificial radioactive elements. (Russian). Sorena, 6(??):3–7, ???? 1934. CODEN ???? ISSN ????

Gamow:1934:ESA

[Gam34c] George Gamow. Empirische Stabilitätsgrenzen von Atomkernen. (German) [Empirical stability limits of atomic nuclei]. Zeitschrift für Physik, 89(9–10):592–596, September 1934. CODEN ZEPYAA.

ISSN 0044-3328. URL http://www.springerlink.com/content/
x6v71670rm3t1576/. See note [Rac35].

Gamow:1934:EDT

[Gam34d] George Gamow. Essay on the development of the theory of atomic nucleus. Decay problem. Uspekhi Fizicheskikh Nauk, 14(??):389– 406, ???? 1934. CODEN UFNAAG. ISSN 0042-1294 (print), 1996-6652 (electronic).

Gamow:1934:ICS

[Gam34e] George Gamow. International Congress on the structure of atomic nuclei. Sorena, ??(??):16–21, ???? 1934. CODEN ???? ISSN ????

Gamow:1934:IN

[Gam34f] George Gamow. Isomeric nuclei? Nature, 133(3370):833, June 2, 1934. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Gamow:1934:MIN

[Gam34g] George Gamow. Modern ideas on nuclear constitution. Nature, 133 (3368):744–747, May 19, 1934. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Gamow:1934:NPN

[Gam34h] George Gamow. Negative protons and nuclear structure. Physical Review, 45(10):728–729, May 15, 1934. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://prola.

aps.org/abstract/PR/v45/i10/p728_1.

Gamow:1934:NSR

[Gam34i] George Gamow. Nuclear spin of radioactive elements. Proceedings of the Royal Society of London. Series A, Containing Papers of a Mathematical and Physical Character, 146(856):217–222, August 1, 1934. ISSN 0950-1207 (print), 2053-9150 (electronic). URL

http://www.jstor.org/stable/pdfplus/2935489.pdf.

Gamow:1934:SGS

[Gam34j] George Gamow. A sketch of the growth of the study of the structure of the atomic nucleus. (Russian). Physics-Uspekhi, 14(??):??, ???? 1934. CODEN PHUSEY. ISSN 1063-7869 (print), 1468-4780 (electronic).

Gamow:1935:GSP

[Gam35a] G. Gamow. General stability-problems of atomic nuclei. In Papers and discussion of the International Conference on Physics. London, 1934, V. A., volume 1, pages 60–71. ????, ????, 1935.

Gamow:1935:NP

[Gam35b] George Gamow. The negative proton. Nature, 135(3421):858–861, May 24, 1935. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Gamow:1935:NTO

[Gam35c] George Gamow. Nuclear transformations and the origin of the chemical elements. Ohio Journal of Science, 35(5):406– 413, September 1935. CODEN OJSCA9. ISSN 0030-0950.

URL http://hdl.handle.net/1811/2763;   https://kb.osu.
edu/dspace/bitstream/handle/1811/2763/V35N05_406.pdf.

Gamow:1935:QTN

[Gam35d] George Gamow. Quantum theory of nuclear structure. In International conference in physics. London, 1934. Papers and discussions: V. A. London, volume ??, pages 60–71. ????, ????, 1935.

Gamow:1935:NAF

[Gam35e] Georges Gamow. Les noyaux atomiques. (French) [Atomic nuclei]. Annales de l’Institut Henri Poincaré, 5(2):89–114, 1935. CODEN AIHPA2. ISSN 0365-320x (print), 2400-4855 (electronic). URL

http://www.numdam.org/item?id=AIHP_1935__5_2_89_0.

Gamow:1936:CRN

[Gam36a] George Gamow. Cinétique des réactions nucléares. (French) [Kinetics of Nuclear Reactions]. Hermann, Paris, France, 1936. ???? pp. LCCN ????

Gamow:1936:LCR

[Gam36b] George Gamow. … Cinétique des réactions nucléaires. (French) [Kinetics of nuclear reactions]. Exposés de physique théorique, publié sous la direction de m. Louis de Broglie… xvi. Hermann, Paris, France, 1936. 21 + 1 pp. LCCN Q111 .A3 no. 369.

Gamow:1936:PSP

[Gam36c] George Gamow. Possibility of selective phenomena for fast neutrons. Physical Review, 49(12):946, June 15, 1936. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL

http://prola.aps.org/abstract/PR/v49/i12/p946_1.

Gamow:1937:SAN

[Gam37a] George Gamow. Structure of Atomic Nuclei and Nuclear Transformations, volume 235 of The international series of monographs on physics. Clarendon Press, Oxford, UK, second edition, 1937. xi + 1 + 270 + 1 pp. LCCN QC173 .G3 1937.

Gamow:1937:HSJ

[Gam37b] George Gamow. Über den heutigen Stand (1. Juni 1937) der Theorie des β-Zerfalls [On the current status (1 June 1933) of the theory of β decay]. Physikalische Zeitschrift, 38(??):800–814, ???? 1937. CODEN PHZTAO. ISSN 0369-982X.

Gamow:1938:KES

[Gam38a] G. Gamow. Kernumwandlungen als Energiequelle der Sterne. (German) [Nuclear reactions as an energy source of stars]. Zeitschrift f”ur Astrophysik, 16:113–160, 1938. CODEN ZEASAJ. ISSN 0372-8331.

Gamow:1938:SMS

[Gam38b] G. Gamow. A star model with selective thermo-nuclear source. Astrophysical Journal, 87:206–208, March 1938. CODEN ASJOAB. ISSN 0004-637X (print), 1538-4357 (electronic). URL http://

adsabs.harvard.edu/abs/1938ApJ....87..206G.

Gamow:1938:ZBK

[Gam38c] G. Gamow. Zusammenfassender Bericht. Kernumwandlungen als Energiequelle der Sterne. Mit 10 Abbildungen. (German) [Summary report. Nuclear transformations as energy sources of stars. With 10 illustrations]. Zeitschrift f”ur Astrophysik, 16(3):113–160, 1938. CODEN ZEASAJ. ISSN 0372-8331. URL http://adsabs.

harvard.edu/abs/1938ZA.....16..113G.

Gamow:1938:NES

[Gam38d] George Gamow. Nuclear energy sources and stellar evolution. Physical Review, 53(7):595–604, April 1938. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

URL http://adsabs.harvard.edu/abs/1938PhRv...53..595G;
 http://prola.aps.org/abstract/PR/v53/i7/p595_1.

Gamow:1938:TTN

[Gam38e] George Gamow. Tentative theory of novæ. Physical Review, 54(6): 480, September 1938. CODEN PHRVAO. ISSN 0031-899X (print),

1536-6065 (electronic). URL http://adsabs.harvard.edu/abs/
1938PhRv...54..480G;   http://prola.aps.org/abstract/PR/
v54/i6/p480_1.

Gamow:1938:TSE

[Gam38f] George Gamow. Tracks of stellar evolution. Physical Review, 53 (11):907–908, June 1938. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://adsabs.harvard.

edu/abs/1938PhRv...53..907G;   http://prola.aps.org/abstract/
PR/v53/i11/p907_1.

Gamow:1938:LED

[Gam38g] Georges Gamow. L’évolution des étoiles du point de vue de la physique moderne. (French) [The evolution of stars in the view of modern physics]. Annales de l’Institut Henri Poincaré, 8(5): 193–211, 1938. CODEN AIHPA2. ISSN 0365-320x (print), 2400-

4855 (electronic). URL http://www.numdam.org/item?id=AIHP_
1938__8_5_193_0.

Gamow:1939:EPR

[Gam39a] G. Gamow. The energy-producing reaction in the sun. Astrophysical Journal, 89:130–133, January 1939. CODEN ASJOAB. ISSN 0004-637X (print), 1538-4357 (electronic). URL http://adsabs.

harvard.edu/abs/1939ApJ....89..130G.

Gamow:1939:ERG

[Gam39b] George Gamow. The evolution of Red Giants. Physical Review, 55(8):796–797, April 1939. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://prola.aps.org/

abstract/PR/v55/i8/p796_2.

Gamow:1939:MTW

[Gam39c] George Gamow. Mr. Tompkins in Wonderland: or, Stories of c, G, and h. Cambridge University Press, Cambridge, UK, 1939. x + 91 pp. LCCN ???? Illustrated by John Hookham.

Gamow:1939:NRSa

[Gam39d] George Gamow. Nuclear reactions in stellar evolution. Nature, 144(3648):575–577, September 30, 1939. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://adsabs.

harvard.edu/abs/1939Natur.144..575G.

Gamow:1939:NRSb

[Gam39e] George Gamow. Nuclear reactions in stellar evolution*. Nature, 144(3649):620–622, October 7, 1939. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://adsabs.

harvard.edu/abs/1939Natur.144..620G.

Gamow:1939:PPS

[Gam39f] George Gamow. Physical possibilities of stellar evolution. Physical Review, 55(8):718–725, April 1939. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http:/

/adsabs.harvard.edu/abs/1939PhRv...55..718G;   http://
link.aps.org/doi/10.1103/PhysRev.55.718.

Gamow:1940:BPN

[Gam40a] George Gamow. Basic principles of the new mechanics. The Scientific Monthly, 51(4):358–364, October 1940. CODEN SCMOAA. ISSN 0096-3771 (print), 2327-7513 (electronic). URL http://www.

jstor.org/stable/pdfplus/17355.pdf.

Gamow:1940:BDS

[Gam40b] George Gamow. The Birth and Death of the Sun: Stellar Evolution and Subatomic Energy. Viking Press, New York, NY, USA, 1940. xiv + 1 + 238 pp. LCCN QB44 .G26. URL http://adsabs.

harvard.edu/abs/1940QB44.G26........

Gamow:1940:DMP

[Gam40c] George Gamow. Dualism in the modern physics. The Scientific Monthly, 50(1):78–79, January 1940. CODEN SCMOAA. ISSN 0096-3771 (print), 2327-7513 (electronic). URL http://www.

jstor.org/stable/pdfplus/17022.pdf.

Gamow:1940:ES

[Gam40d] George Gamow. The evolution of the stars. Nature, 146(3690): 97, July 20, 1940. CODEN NATUAS. ISSN 0028-0836 (print),

1476-4687 (electronic). URL http://adsabs.harvard.edu/abs/
1940Natur.146...97G.

Gamow:1940:RUS

[Gam40e] George Gamow. Review: Has the Universe a soul? : The Soul of the Universe by Gustaf Stromberg. The Scientific Monthly, 51(6):564–565, December 1940. CODEN SCMOAA. ISSN 0096-3771 (print), 2327-7513 (electronic). URL http:/

/adsabs.harvard.edu/abs/1940SciMo..51..564S;   http://
www.jstor.org/stable/pdfplus/17151.pdf.

Gamow:1941:OSB

[Gam41a] G. Gamow. Our sun is bound to explode. Popular Astronomy, 49: 360–??, August 1941. CODEN ???? ISSN ???? URL http://

adsabs.harvard.edu/abs/1941PA.....49..360G.

Gamow:1941:BEP

[Gam41b] George Gamow. Biography of the Earth: Its Past, Present, and Future. Viking Press, New York, NY, USA, 1941. xiii, [1], 242 pp. LCCN QE501 .G35.

Gamow:1941:HSB

[Gam41c] George Gamow. How stars are born. American Weekly, ??(??):??, June 22, 1941. CODEN ???? ISSN ????

Gamow:1941:RID

[Gam41d] George Gamow. Relative importance of different elements for neutrino production. Physical Review, 59(7):617–618, April 1, 1941. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065

(electronic). URL http://prola.aps.org/abstract/PR/v59/
i7/p617_1.

Gamow:1942:LHT

[Gam42a] G. Gamow. A layman’s history of the theory of Relativity. (books on science for laymen: From Copernicus to Einstein). The Scientific Monthly, 55:569–??, December 1942. CODEN SCMOAA. ISSN 0096-3771 (print), 2327-7513 (electronic). URL http://

adsabs.harvard.edu/abs/1942SciMo..55..569R.

Gamow:1942:BTS

[Gam42b] George Gamow. Biografía de la tierra. (Spanish) [Story of the Earth]. Espasa-Calpe argentina, s. a., Buenos Aires, Argentina y México, Méico, 1942. 3 + 9–264 pp. LCCN QE501 .G353. Translation to Spanish by Manuel Balanzat.

Gamow:1942:BE

[Gam42c] George Gamow. Biography of the Earth. Science, 96(2482):84–85, July 24, 1942. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

Gamow:1942:BEI

[Gam42d] George Gamow. Biography of the Earth. Science, 96(2482):84– 85, July 24, 1942. CODEN SCIEAS. ISSN 0036-8075 (print),

1095-9203 (electronic). URL http://www.jstor.org/stable/
pdfplus/1670096.pdf.

Gamow:1942:COC

[Gam42e] George Gamow. Concerning the origin of chemical elements. Journal of the Washington Academy of Sciences, 32(12):353–355, December 15, 1942. CODEN JWASA3. ISSN 0043-0439. URL https:

//archive.org/stream/journalofwashin321942wash#page/353/

.

Gamow:1942:MMW

[Gam42f] George Gamow. Many more worlds like ours. American Weekly, ??(??):??, January 4, 1942. CODEN ???? ISSN ????

Gamow:1942:MTD

[Gam42g] George Gamow. Mr. Tompkins i Drømmeland. (Danish) [Mr. Tompkins in Wonderland]. Gyldendalske Boghandel Nordisk Forlag, København, Danmark, 1942. 95 pp. Forord af Niels Bohr.

Gamow:1942:NVS

[Gam42h] George Gamow. Neutrinos vs. supernovæ. The Scientific Monthly, 54(1):65–70, January 1942. CODEN SCMOAA. ISSN 0096-3771 (print), 2327-7513 (electronic). URL http:/

/adsabs.harvard.edu/abs/1942SciMo..54...65G;   http://
www.jstor.org/stable/pdfplus/17474.pdf.

Gamow:1942:RLH

[Gam42i] George Gamow. Review: a layman’s history of the theory of relativity: From Copernicus to Einstein by Hans Reichenbach. The Scientific Monthly, 55(6):569–??, December 1942. CODEN SC-MOAA. ISSN 0096-3771 (print), 2327-7513 (electronic). URL

http://www.jstor.org/stable/pdfplus/17860.pdf.

Gamow:1942:RDP

[Gam42j] George Gamow. Review: Do physical theories support philosophical doctrines? The Scientific Monthly, 54(5):473–474, May 1942. CODEN SCMOAA. ISSN 0096-3771 (print), 2327-7513 (elec-

tronic). URL http://www.jstor.org/stable/pdfplus/17571.
pdf.

Gamow:1943:CEM

[Gam43a] G. Gamow. Contractive evolution of massive stars. Astrophysical Journal, 98:498–499, November 1943. CODEN ASJOAB. ISSN 0004-637X (print), 1538-4357 (electronic). URL http://adsabs.

harvard.edu/abs/1943ApJ....98..498G.

Gamow:1943:WWS

[Gam43b] G. Gamow. On WC and WN stars. Astrophysical Journal, 98: 500, November 1943. CODEN ASJOAB. ISSN 0004-637X (print),

1538-4357 (electronic). URL http://adsabs.harvard.edu/abs/
1943ApJ....98..500G.

Gamow:1944:ECS

[Gam44a] George Gamow. The evolution of contracting stars. Physical Review, 65(1–2):20–32, January 1, 1944. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http:/

/adsabs.harvard.edu/abs/1944PhRv...65...20G;   http://
prola.aps.org/abstract/PR/v65/i1-2/p20_1.

Gamow:1944:MTE

[Gam44b] George Gamow. Mr. Tompkins Explores the Atom. Macmillan Publishing Company, New York, NY, USA, 1944. x + 97 pp. LCCN QC173 .G32.

Gamow:1945:BDS

[Gam45a] George Gamow. The Birth and Death of the Sun: Stellar Evolution and Subatomic Energy. Pelican books. Penguin, New York, NY, USA, 1945. xv + 2 + 219 pp. LCCN QB44 .G26 1945. URL

http://adsabs.harvard.edu/abs/1945QB44.G26........

Gamow:1945:RGS

[Gam45b] George Gamow. The Red-Giant stage of stellar evolution. Physical Review, 67(3–4):120–121, February 1, 1945. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

URL http://adsabs.harvard.edu/abs/1945PhRv...67..120G;
 http://prola.aps.org/abstract/PR/v67/i3-4/p120_2.

Gamow:1946:AEC

[Gam46a] George Gamow. Atomic Energy in Cosmic and Human Life: Fifty Years of Radioactivity. Cambridge University Press, Cambridge, UK, 1946. x + 161 pp. LCCN QC173 .G28.

Gamow:1946:EUO

[Gam46b] George Gamow. Expanding universe and the origin of elements. Physical Review, 70(7–8):572–573, October 1, 1946. CO-DEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

URL http://adsabs.harvard.edu/abs/1946PhRv...70..572G;
 http://prola.aps.org/abstract/PR/v70/i7-8/p572_2. See

erratum [Gam47c].

Gamow:1946:MTU

[Gam46c] George Gamow. Mr. Tompkins utforskar atomen. (Swedish) [Mr. Tompkins explores the atom]. Natur och kultur, Stockholm, Sweden, 1946. 102 pp. LCCN QC778 .G37.

Gamow:1946:RU

[Gam46d] George Gamow. Rotating universe? Nature, 158(4016):549, October 19, 1946. CODEN NATUAS. ISSN 0028-0836 (print),

1476-4687 (electronic). URL http://adsabs.harvard.edu/abs/

1946Natur.158..549G. This short letter is reprinted in its entirety in [Rin09, Figure 2, page 500].

Gamow:1947:AMO

[Gam47a] George Gamow. Atomen, människan och universum. (Swedish) [The Atom, Man, and the Universe]. Bonnier, Stockholm, Sweden, 1947. ???? pp. LCCN QC778 .G358.

Gamow:1947:AEC

[Gam47b] George Gamow. Atomic Energy in Cosmic and Human Life: Fifty Years of Radioactivity. Cambridge University Press, Cambridge, UK, 1947. ???? pp. LCCN QC173 .G28 1947.

Gamow:1947:EEU

[Gam47c] George Gamow. Erratum: Expanding universe and the origin of elements Phys. Rev. 70, 572–573 (1946). Physical Review, 71(4):273, February 15, 1947. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http:/

/adsabs.harvard.edu/abs/1947PhRv...71..273G;   http://
prola.aps.org/abstract/PR/v71/i4/p273_1. See [Gam46b].

Gamow:1947:GTS

[Gam47d] George Gamow. Geburt und Tod der Sonne: Sternbildung und subatomare Energie. (German) [The Birth and Death of the Sun: Stellar Evolution and Subatomic Energy], volume 3 of Wissenschaft und Kultur. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1947. xviii + 284 pp. LCCN QB44 .G263. Translated from English to German by Baron E. (Emanuel) von der Pahlen.

Gamow:1947:OTT

[Gam47e] George Gamow. One, Two, Three, … Infinity. Viking Press, New York, NY, USA, 1947. ???? pp. LCCN ????

Gamow:1947:PNM

[Gam47f] George Gamow. Probability of nuclear meson-absorption. Physical Review, 71(8):550–551, April 15, 1947. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://prola.

aps.org/abstract/PR/v71/i8/p550_2.

Gamow:1948:BEP

[Gam48a] George Gamow. Biography of the Earth: Its Past, Present and Future. N.A.L. Mentor books; A Pelican Mentor book. New American Library, New York, NY, USA, 1948. 194 pp. LCCN QE501 .G35 1948.

Gamow:1948:EU

[Gam48b] George Gamow. The evolution of the universe. Nature, 162(4122): 680–682, October 30, 1948. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://adsabs.harvard.

edu/abs/1948Natur.162..680G.

Gamow:1948:GF

[Gam48c] George Gamow. Galaxies in flight. Scientific American, 179(1): 20–24, July 1948. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Gamow:1948:MTS

[Gam48d] George Gamow. Mixed types of stellar populations. Nature, 162(4125):816, November 20, 1948. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://adsabs.

harvard.edu/abs/1948Natur.162..816G.

Gamow:1948:OES

[Gam48e] George Gamow. The origin of elements and the separation of galaxies. Physical Review, 74(4):505–506, August 15, 1948. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065

(electronic). URL http://adsabs.harvard.edu/abs/1948PhRv.
..74..505G;   http://prola.aps.org/abstract/PR/v74/i4/

p505_2. Reprinted in [BF86].

Gamow:1948:OI

[Gam48f] George Gamow. Origin of the ice. Scientific American, 179(4): ??, October 1948. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Gamow:1948:RN

[Gam48g] George Gamow. The reality of neutrinos. Physics Today, 1(3):4–7, 30–31, July 1948. CODEN PHTOAD. ISSN 0031-9228 (print),

1945-0699 (electronic). URL http://link.aip.org/link/?PTO/
1/4/1.

Gamow:1949:PT

[Gam49a] George Gamow. Any physics tomorrow? Physics Today, 2(1):16– 21, January 1949. CODEN PHTOAD. ISSN 0031-9228 (print),

1945-0699 (electronic). URL http://link.aip.org/link/?PTO/

2/16/1. See minor erratum [Gam49d].

Gamow:1949:BDS

[Gam49b] George Gamow. The Birth and Death of the Sun. Viking Press, New York, NY, USA, 1949. xvi + 245 pp. LCCN QB44 .G26

1949. URL http://adsabs.harvard.edu/abs/1949QB44.G26...

..... Original edition 1945.

Gamow:1949:EN

[Gam49c] George Gamow. Die Existenz der Neutrinos. (German) [Existence of the neutrino]. Physikalische Blätter, 5(3):108–114, March 1949. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (elec-

tronic). URL http://onlinelibrary.wiley.com/doi/10.1002/
phbl.19490050302/abstract.

Gamow:1949:E

[Gam49d] George Gamow. Errata. Physics Today, 2(4):35, April 1949. CO-DEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

URL http://www.physicstoday.org/resource/1/phtoad/v2/

i4/p35_s1. See [Gam49a].

Gamow:1949:RC

[Gam49e] George Gamow. On relativistic cosmogony. Reviews of Modern Physics, 21(3):367–373, July 1949. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//adsabs.harvard.edu/abs/1949RvMP...21..367G;   http:/
/link.aps.org/doi/10.1103/RevModPhys.21.367;   http://
rmp.aps.org/abstract/RMP/v21/i3/p367_1.

Gamow:1949:RBP

[Gam49f] George Gamow. Review: Unitary Principle in Physics and Biology by Lancelot Law Whyte. Science, 110(2862):483–??, November 4, 1949. CODEN SCIEAS. ISSN 0036-8075 (print),

1095-9203 (electronic). URL http://www.jstor.org/stable/
pdfplus/1676041.pdf.

Gamow:1949:S

[Gam49g] George Gamow. Supernovae. Scientific American, 181(6):??, December 1949. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Gamow:194x:NE

[Gam4xa] George Gamow. Near the end? Time, 53(??):??, ???? ??, 194x. CODEN TYMEA9. ISSN 0040-781x (print), 2169-1665 (electronic).

Gamow:194x:SAF

[Gam4xb] George Gamow. Sun’s atomic fuel. Science Illustrated, 2(??):??, 194x. CODEN ???? ISSN ????

Gamow:194x:US

[Gam4xc] George Gamow. Universal spin. Newsweek, 28(??):??, 194x. CO-DEN ???? ISSN 0028-9604.

Gamow:1950:HHC

[Gam50a] George Gamow. Half an hour of creation. . .. Physics Today, 3(8): 16–21, August 1950. CODEN PHTOAD. ISSN 0031-9228 (print),

1945-0699 (electronic). URL http://link.aip.org/link/?PTO/
3/16/1.

Gamow:1950:RBN

[Gam50b] George Gamow. Review: The Nature of Physical Reality: A Philosophy of Modern Physics by Henry Margenau. Science, 112 (2901):155–??, August 4, 1950. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.jstor.

org/stable/pdfplus/1679660.pdf.

Gamow:1950:RTO

[Gam50c] George Gamow. Review: Thoughts, observations, and conclusions: Phenomena, Atoms and Molecules by Irving Langmuir. The Scientific Monthly, 70(5):340–341, May 1950. CODEN SCMOAA. ISSN 0096-3771 (print), 2327-7513 (electronic). URL http://www.

jstor.org/stable/pdfplus/19953.pdf.

Gamow:1951:BRK

[Gam51a] George Gamow. Book review: Karl Jellinek, Verstandliche Elemente der Wellenmechanik: Photonen, Freie Elektronen, Einelektronige Atome, Teil I. Science, 113(2930):218, February 23, 1951. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

Gamow:1951:BRW

[Gam51b] George Gamow. Book review: Walter Weizel, Lehrbuch der Theoretischen Physik: Struktur der Materie, Vol. II, West Berlin, Germany: Springer-Verlag, 1950. Pp. 772–1,540. DM 69.90 bound, DM 66 paper. Science, 114(2953):137, August 3, 1951. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

Gamow:1951:HEE

[Gam51c] George Gamow. Hydrogen exhaustion and explosions of stars. Nature, 168(4263):72–73, July 14, 1951. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://adsabs.

harvard.edu/abs/1951Natur.168...72G.

Gamow:1951:M

[Gam51d] George Gamow. The Moon. Abelard-Schuman, London, UK, revised edition, 1951. 127 pp. LCCN ???? Illustrations by Bunji Tagawa.

Gamow:1951:OEU

[Gam51e] George Gamow. The origin and evolution of the Universe. American Scientist, 39(3):392–406, ???? 1951. CODEN AMSCAC. ISSN 0003-0996 (print), 1545-2786 (electronic). URL http://www.

jstor.org/stable/27826381.

Gamow:1951:RBT

[Gam51f] George Gamow. Review: Lehrbuch der Theoretischen Physik: Struktur der Materie by Walter Weizel. Science, 114(2953):137– ??, August 3, 1951. CODEN SCIEAS. ISSN 0036-8075 (print),

1095-9203 (electronic). URL http://www.jstor.org/stable/
pdfplus/1678852.pdf.

Gamow:1951:RBE

[Gam51g] George Gamow. Review: Verstandliche Elemente der Wellenmechanik: Photonen, Freie Elektronen, Einelektronige Atome by Teil I. Karl Jellinek. Science, 113(2930):218–??, February 23, 1951. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

URL http://www.jstor.org/stable/pdfplus/1679431.pdf.

Gamow:1951:WOS

[Gam51h] George Gamow. Will our Sun ever explode? Proceedings of the American Academy of Arts and Sciences, 79(4):291–294, July 1951. CODEN PAAAAV. ISSN 0065-6836. URL http://www.jstor.

org/stable/pdfplus/20023615.pdf.

Gamow:1952:HCT

[Gam52a] G. Gamow. History of cosmological theories. (book reviews: Cosmology). The Scientific Monthly, 75:65–??, July 1952. CODEN SCMOAA. ISSN 0096-3771 (print), 2327-7513 (electronic). URL

http://adsabs.harvard.edu/abs/1952SciMo..75...65G.

Gamow:1952:BDS

[Gam52b] George Gamow. The Birth and Death of the Sun: Stellar Evolution and Sub-atomic Energy, volume M77 of A Mentor book. New American Library, New York, NY, USA, 1952. 219 pp. LCCN QB44

.G26 1952. URL http://adsabs.harvard.edu/abs/1952QB44.
G26........

Gamow:1952:CU

[Gam52c] George Gamow. The Creation of the Universe. Viking Press, New York, NY, USA, 1952. xii + 147 pp. LCCN QB981 .G3. URL

http://adsabs.harvard.edu/abs/1952QB981.G3........

Gamow:1952:RHC

[Gam52d] George Gamow. Review: History of cosmological theories: Cosmology by H. Bondi. The Scientific Monthly, 75(1):65–??, July 1952. CODEN SCMOAA. ISSN 0096-3771 (print), 2327-7513 (elec-

tronic). URL http://www.jstor.org/stable/pdfplus/20618.
pdf.

Gamow:1952:RTE

[Gam52e] George Gamow. The role of turbulence in the evolution of the universe. Physical Review, 86(2):251, April 15, 1952. CO-DEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

URL http://adsabs.harvard.edu/abs/1952PhRv...86..251G;
 http://prola.aps.org/abstract/PR/v86/i2/p251_1.

Gamow:1952:TS

[Gam52f] George Gamow. Turbulence in space. Scientific American, 186 (6):??, June 1952. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Gamow:1953:M

[Gam53a] G. Gamow. The Moon. H. Schuman, New York, NY, USA, 1953.

URL http://adsabs.harvard.edu/abs/1953QB581.G3........

Gamow:1953:OP

[Gam53b] G. Gamow. Origin of protogalaxies. Astronomical Journal, 58: 39, March 1953. CODEN ANJOAA. ISSN 0004-6256 (print),

1538-3881 (electronic). URL http://adsabs.harvard.edu/abs/
1953AJ.....58Q..39G.

Gamow:1953:A

[Gam53c] George Gamow. The alphabet. Physics Today, 6(2):20, February 1953. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699

(electronic). URL http://www.physicstoday.org/resource/1/
phtoad/v6/i2/p20_s1.

Gamow:1953:EUO

[Gam53d] George Gamow. Expanding universe and the origin of galaxies. Det Kongelige Danske Videnskabernes Selskab, Matematisk-Fysiske Meddelelser, 27(10):1–15 (??), ???? 1953. CODEN ???? ISSN ????

Gamow:1953:MTL

[Gam53e] George Gamow. Mr. Tompkins Learns the Facts of Life. Cambridge University Press, Cambridge, UK, 1953. 87 pp. LCCN QH309 .G3.

Gamow:1953:OEU

[Gam53f] George Gamow. The origin and evolution of the universe. In Baitsell [Bai53], page ?? LCCN ????

Gamow:1953:RBE

[Gam53g] George Gamow. Review: The End of the World: A Scientific Inquiry by Kenneth Heuer. The Scientific Monthly, 77(6):319– 320, December 1953. CODEN SCMOAA. ISSN 0096-3771 (print),

2327-7513 (electronic). URL http://www.jstor.org/stable/
pdfplus/21026.pdf.

Gamow:1954:SST

[Gam54a] G. Gamow. On the steady-state theory of the Universe. Astronomical Journal, 59(1217):200, June 1954. CODEN ANJOAA. ISSN 0004-6256 (print), 1538-3881 (electronic). URL http://adsabs.

harvard.edu/abs/1954AJ.....59..200G.

Gamow:1954:LEG

[Gam54b] George Gamow. Die Lebensgeschichte der Erde. (German) [The Life History of Earth]. Bruckmann, München, Germany, 1954. 183 pp. LCCN ????

Gamow:1954:MC

[Gam54c] George Gamow. Modern cosmology. Scientific American, 190(3): 61–??, March 1954. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Gamow:1954:ITN

[Gam54d] George Gamow. On information transfer from nucleic acids to proteins. Det Kongelige Danske Videnskabernes Selskab, Biologiske Meddelelser, 22(3):??, ???? 1954. CODEN ???? ISSN ????

Gamow:1954:FPT

[Gam54e] George Gamow. On the formation of protogalaxies in the turbulent primordial gas. Proceedings of the National Academy of Sciences of the United States of America, 40(6):480–484, June 15, 1954. CODEN PNASA6. ISSN 0027-8424 (print), 1091-6490 (electronic).

URL http://adsabs.harvard.edu/abs/1954PNAS...40..480G;
 http://www.jstor.org/stable/pdfplus/88967.pdf.

Gamow:1954:PMR

[Gam54f] George Gamow. Possible mathematical relation between deoxyribonucleic acid and proteins. Det Kongelige Danske Videnskabernes Selskab, Biologiske Meddelelser, 22(3):1–13, ???? 1954. CODEN ???? ISSN ????

Gamow:1954:PRB

[Gam54g] George Gamow. Possible relation between deoxyribonucleic acid and protein. Nature, 173(4398):318, February 13, 1954. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Gamow:1954:RRP

[Gam54h] George Gamow. Review: The Revolution in Physics by Louis de Broglie; Ralph W. Niemeyer (translator). The Scientific Monthly, 78(1):48, January 1954. CODEN SCMOAA. ISSN 0096-3771 (print), 2327-7513 (electronic). URL http://www.jstor.org/

stable/21151.

Gamow:1966:MC

[Gam55a] G. Gamow. Modern cosmology. In The new astronomy : a Scientific American book [Ano55b], pages 3–?? LCCN ???? URL

http://adsabs.harvard.edu/abs/1966neas.book....3G.

Gamow:1955:GGP

[Gam55b] George Gamow. George Gamow papers, [ca. 1935–ca. 1955]. George Washington University library archives, 1955.

Gamow:1955:ITL

[Gam55c] George Gamow. Information transfer in the living cell. Scientific American, 193(4):70–78, October 1955. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Gamow:1955:LEV

[Gam55d] George Gamow. Lune, escale vers l’infini. (French) [The Moon: Stairway to Infinity]. Plon, Paris, France, 1955. viii + 103 pp. LCCN ???? Translated to French by Denise Meunier.

Gamow:1955:ITN

[Gam55e] George Gamow. On information transfer from nucleic acids to proteins. Det Kongelige Danske Videnskabernes Selskab, Biologiske Meddelelser, 22(8):1–7, ???? 1955. CODEN ???? ISSN ????

Gamow:1955:TPC

[Gam55f] George Gamow. Topological properties of coiled helical systems. Proceedings of the National Academy of Sciences of the United States of America, 41(1):7–9, January 15, 1955. CODEN PNASA6. ISSN 0027-8424 (print), 1091-6490 (electronic). URL http://www.

jstor.org/stable/pdfplus/88999.pdf.

Gamow:1956:EU

[Gam56a] George Gamow. Evolutionary universe. Scientific American, 195 (3):136–154, September 1956. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Gamow:1956:GCM

[Gam56b] George Gamow. Gravity: Classic and Modern Views. University of Colorado, Boulder, CO, USA, 1956. 157 pp. LCCN ????

Gamow:1956:PEU

[Gam56c] George Gamow. The physics of the expanding universe. In Arthur Beer, editor, Vistas in Astronomy, volume 2 (of 41), pages 1726– 1732. Pergamon, New York, NY, USA, 1956. ISSN 0083-6656 (print), 1872-9207 (electronic). URL http://adsabs.harvard.

edu/abs/1956VA......2.1726G.

Gamow:1957:PIT

[Gam57a] George Gamow. The problem of information transfer from nucleic acids to proteins. (Russian). In D. M. Frank, editor, Voprosi Biofiziki [Problems of Biophysics], pages 205–263. ????, ????, 1957. LCCN ????

Gamow:1957:X

[Gam57b] George Gamow. xxx. In Voprosy Biofiziki. (Russian) [Problems in Biophysics], pages 203–?? Inostrannaya Literatura, Moscow, USSR, 1957. LCCN ????

Gamow:1958:CUI

[Gam58a] George Gamow. The creation of the universe. The Sewanee Review, 66(3):413–422, Summer 1958. CODEN ???? ISSN 0037-3052 (print), 1934-421X (electronic). URL http://www.jstor.

org/stable/pdfplus/27538746.pdf.

Gamow:1958:MES

[Gam58b] George Gamow. Matter, Earth, and Sky. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1958. 593 pp. LCCN QC171 .G3.

Gamow:1958:PU

[Gam58c] George Gamow. The Principle of Uncertainty. Scientific American, 198(1):51–57, January 1958. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Gamow:1959:M

[Gam59a] G. Gamow. The Moon. Abelard-Schuman, London, UK, re-

vised edition, 1959. URL http://adsabs.harvard.edu/abs/
1959QB581.G3........

Gamow:1959:BEP

[Gam59b] George Gamow. Biography of the Earth: Its Past, Present, and Future, volume C53 of Compass books. Viking Press, New York, NY, USA, revised edition, 1959. xiii + 242 pp. LCCN QE501 .G35 1959a.

Gamow:1959:EP

[Gam59c] George Gamow. The Exclusion Principle. Scientific American, 201 (1):74–86, July 1959. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Gamow:1959:MES

[Gam59d] George Gamow. Matter, Earth, and sky. Macmillan Publishing Company, New York, NY, USA, 1959.

Gamow:195x:ST

[Gam5x] George Gamow. Start of things. Newsweek, 39(??):??, ???? ??, 195x. CODEN ???? ISSN 0028-9604.

Gamow:1960:JNB

[Gam60] George Gamow. Der junge Niels Bohr Zum 75. Geburtstag, nach Erinnerungen. (German) [The young Niels Bohr: 75th birthday memoirs]. Physikalische Blätter, 16(10):525–527, October 1960. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (elec-

tronic). URL http://onlinelibrary.wiley.com/doi/10.1002/
phbl.19600161008/abstract.

Gamow:1961:AN

[Gam61a] George Gamow. The Atom and its Nucleus. A Spectrum book, S-St-1. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1961. 153 pp. LCCN QC171 .G29.

Gamow:1961:BPa

[Gam61b] George Gamow. Biography of Physics, volume TB567 of Harper torchbooks. The Cloister library. Harper, New York, NY, USA, 1961. 338 pp. LCCN QC7 .G263 1964.

Gamow:1961:CU

[Gam61c] George Gamow. The Creation of the Universe. Viking Press, New York, NY, USA, revised edition, 1961. 147 pp. LCCN QB981 .G3

1961. URL http://adsabs.harvard.edu/abs/1961QB981.G3...
.....

Gamow:1961:G

[Gam61d] George Gamow. Gravity. Scientific American, 204(3):94–106, March 1961. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-

7087 (electronic). URL http://www.scientificamerican.com/
article.cfm?id=gamow-gravity;   http://www.scientificamerican.
com/mar2011/gamow.

Gamow:1961:HOS

[Gam61e] George Gamow. The heart on the other side. University of Colorado Literary Magazine, 72(??):??, 1961. CODEN ???? ISSN ????

Gamow:1961:RLC

[Gam61f] George Gamow. Remarks on Lorentz contraction. Proceedings of the National Academy of Sciences of the United States of America, 47(5):728–729, May 15, 1961. CODEN PNASA6. ISSN 0027-8424 (print), 1091-6490 (electronic). URL http://www.jstor.

org/stable/pdfplus/70983.pdf.

Gamow:1962:BPb

[Gam62a] George Gamow. Biography of Physics. Harper torchbooks. Science library. Hutchinson Science Library, London, UK, 1962. x + 338 pp. LCCN ????

Gamow:1962:Ga

[Gam62b] George Gamow. Gravity. Doubleday, Garden City, NY, USA, 1962. 157 pp. LCCN ????

Gamow:1962:Gb

[Gam62c] George Gamow. Gravity. Heinemann, London, UK, 1962. 157 pp. LCCN ????

Gamow:1962:HOS

[Gam62d] George Gamow. The heart on the other side. In The expert dreamers [Poh62], pages 51–61. LCCN PZ1 .P745 Ex. Reprinted in 1968 by Avon Books, New York.

Gamow:1962:GPN

[Gam62e] George Gamow. La Gravitation: de la pomme de Newton aux fusèes interplanètaires. (French) [Gravitation: from Newton’s apple to interplanetary rockets]. Payot, Paris, France, 1962. 152 pp. LCCN ???? Translated to French by Dr J. Mètadier.

Gamow:1962:PST

[Gam62f] George Gamow. The physical sciences and technology. In The Great Ideas Today, page ?? Encyclopaedia Britannica, Inc., Chicago, IL, USA, 1962. LCCN ????

Gamow:1963:IDA

[Gam63a] G. Gamow. La investigación del atomo. (Spanish) [The study of the atom]. Fondo de Cultura Economica, México, México, second edi-

tion, 1963. URL http://adsabs.harvard.edu/abs/1963lida.
book.....G.

Gamow:1963:MPA

[Gam63b] G. Gamow. Nicht mehr per Sie mit dem Atom. (German) [No more by using the atom]. Physik-Verlag, Mosbach, West Germany, 1963.

URL http://adsabs.harvard.edu/abs/1963nmps.book.....G.

Gamow:1963:ELT

[Gam63c] George Gamow. Epilogue. on lunar theory. In R. C. Sheriff, editor, The Hopkins Manuscript, page ?? Macmillan Publishing Company, New York, NY, USA, 1963. LCCN ????

Gamow:1963:NBM

[Gam63d] George Gamow. Niels Bohr, the man who explained the atom. Science Digest, ??(??):??, May 1963. CODEN ???? ISSN ????

Gamow:1963:OL

[Gam63e] George Gamow. The origin of life. Transactions of the Bose Research Institute (Calcutta), 24(??):??, ???? 1963. CODEN TBI-CAQ. ISSN 0006-7903.

Gamow:1963:PCE

[Gam63f] George Gamow. A Planet Called Earth. Viking Press, New York, NY, USA, 1963. 257 pp. LCCN QE501 .G36.

Gamow:1963:WL

[Gam63g] George Gamow. What is life? Transactions of the Bose Research Institute (Calcutta), 24(??):??, ???? 1963. CODEN TBICAQ. ISSN 0006-7903.

Gamow:1964:CPG

[Gam64a] George Gamow. Combinatorial principles in genetics. In Beckenbach and Pólya [BP64], chapter 17, pages 515–?? LCCN QA164 .B4 1964.

Gamow:1964:GFG

[Gam64b] George Gamow. Gravitaatio. (Finnish) [Gravity]. WSOY, Porvoo, Finland, 1964. 129 pp. LCCN ????

Gamow:1964:HE

[Gam64c] George Gamow. Heat as energy. In Rapport and Wright [RW64], pages 47–66.

Gamow:1964:SCS

[Gam64d] George Gamow. A Star Called the Sun. Viking Press, New York, NY, USA, 1964. 208 pp. LCCN ???? URL http://adsabs.

harvard.edu/abs/1964QB521.G26.......

Gamow:1965:BPF

[Gam65a] George Gamow. Der Biographie der Physik : Forscher — Ideen — Experimente. (German) [The Biography of Physics: Researchers — Ideas — Experiments]. Econ-Verlag, Düsseldorf, West Germany, 1965. 392 pp. LCCN ????

Gamow:1965:GPG

[Gam65b] George Gamow. Grawitacja. (Polish) [Gravity]. Nowošci Nauki i Techniki. Wiedza Powszechna, Warszawa, Poland, 1965. 107 + 1 pp. LCCN ???? Translated to Polish by Maria Hurwicowa.

Gamow:1965:MES

[Gam65c] George Gamow. Matter, Earth, and Sky. Prentice-Hall, Upper Saddle River, NJ 07458, USA, second edition, 1965. xiv + 624 pp. LCCN QC171 .G3 1965.

Gamow:1965:MTP

[Gam65d] George Gamow. Mr. Tompkins in Paperback. Cambridge University Press, Cambridge, UK, 1965. xii + 186 pp. LCCN QC71 G35

Gamow:1965:NSU

[Gam65e] George Gamow. A new survey of universal knowledge. In Encyclopaedia Britannica, volume 6, pages 578–582. Encyclopaedia Britannica, Chicago, IL, USA; London, UK; Toronto, ON, Canada, 1965. LCCN ????

Gamow:1966:TS

[Gam66a] G. Gamow. Turbulence in space. In The new astronomy : a Scientific American book [Ano55b], pages 66–?? LCCN ???? URL

http://adsabs.harvard.edu/abs/1966neas.book...66G.

Gamow:1966:CTO

[Gam66b] George Gamow. Cosmological theories of the origin of chemical elements. In Robert E. Marshak and J. Warren Blaker, editors, Perspectives in modern physics: essays in honor of Hans A. Bethe on the occasion of his 60th birthday, July 1966, page ?? Wiley, New York, NY, USA, 1966. LCCN ????

Gamow:1966:RBB

[Gam66c] George Gamow. Review of Niels Bohr: The Man. His Science, and the World They Changed, by Ruth Moore. New York Times Book Review, October 23, 1966. CODEN ???? ISSN 0028-7806.

Gamow:1966:RBH

[Gam66d] George Gamow. Review of Otto Hahn: A Scientific Autobiography. New York World-Journal-Tribune, December 1966. CODEN ???? ISSN ????

Gamow:1966:TYSa

[Gam66e] George Gamow. Thirty Years That Shook Physics: the Story of Quantum Theory. Doubleday, Garden City, NY, USA, 1966. xvi + 224 pp. LCCN QC174.1 .G3 1966b.

Gamow:1966:TYSb

[Gam66f] George Gamow. Thirty Years That Shook Physics: the Story of Quantum Theory, volume S45 of Science study series. Anchor Books, New York, NY, USA, 1966. xvi + 224 pp. LCCN QC174.1 .G3.

Gamow:1966:TAS

[Gam66g] George Gamow. Trettio år som skakade fysiken. (Swedish) [Thirty years that shook physics]. Prisma, Stockholm, Sweden, 1966. 155 pp. LCCN ????

Gamow:1967:CVC

[Gam67a] George Gamow. Case of the vanished correlation in statistics of quasi-stellar. Nature, 216(5114):461–462, November 4, 1967. CO-DEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

URL http://adsabs.harvard.edu/abs/1967Natur.216..461G.

Gamow:1967:DGC

[Gam67b] George Gamow. Does gravity change with time? Proceedings of the National Academy of Sciences of the United States of America, 57(2):187–193, February 15, 1967. CO-DEN PNASA6. ISSN 0027-8424 (print), 1091-6490 (electronic).

URL http://adsabs.harvard.edu/abs/1967PNAS...57..187G;
 http://www.jstor.org/stable/pdfplus/57929.pdf.

Gamow:1967:EGC

[Gam67c] George Gamow. Electricity, gravity, and cosmology. Physical Review Letters, 19(13):759–761, September 25, 1967. CODEN PRL-TAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145.

URL http://adsabs.harvard.edu/abs/1967PhRvL..19..759G;
 http://prl.aps.org/abstract/PRL/v19/i13/p759_1. See er-

ratum [Gam67d].

Gamow:1967:EEG

[Gam67d] George Gamow. Erratum: “Electricity, Gravity, and Cosmology”. Physical Review Letters, 19(17):1000, October 23, 1967. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-

0145. URL http://adsabs.harvard.edu/abs/1967PhRvL..19.

1000G. See [Gam67c].

Gamow:1967:HU

[Gam67e] George Gamow. History of the Universe and two letters to Phil. Science, 158(3802):766–769, November 10, 1967. CO-DEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

URL http://adsabs.harvard.edu/abs/1967Sci...158..766G;
 http://www.jstor.org/stable/pdfplus/1723077.pdf.

Gamow:1967:LEa

[Gam67f] George Gamow. Letter to the editor. Science, 158(3802):767–768, November 10, 1967. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

Gamow:1967:LEb

[Gam67g] George Gamow. Letter to the editor. Science, 158(3802):768, November 10, 1967. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

Gamow:1967:SSS

[Gam67h] George Gamow. Sonne — Stern unter Sternen. (German) [A Star Called the Sun]. Ehrenwirth, München, Germany, 1967. 222 pp. LCCN ????

Gamow:1967:STC

[Gam67i] George Gamow. Surface tension and the contraction of muscles. Proceedings of the National Academy of Sciences of the United States of America, 57(3):696–697, March 15, 1967. CO-DEN PNASA6. ISSN 0027-8424 (print), 1091-6490 (electronic).

URL http://www.jstor.org/stable/pdfplus/57576.pdf.

Gamow:1967:VEC

[Gam67j] George Gamow. Variability of elementary charge and quasistellar objects. Physical Review Letters, 19(16):913–914, October 16, 1967. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (elec-

tronic), 1092-0145. URL http://prl.aps.org/abstract/PRL/
v19/i16/p913_1.

Gamow:1968:CPG

[Gam68a] George Gamow. Combinatorial principles in genetics. In M. E. Deza, editor, Applied Combinatorial Mathematics. (Russian), pages 289–308. Mir, Moscow, USSR, 1968. LCCN ????

Gamow:1968:NU

[Gam68b] George Gamow. Naming the units. Nature, 219(5115):765, August 17, 1968. CODEN NATUAS. ISSN 0028-0836 (print),

1476-4687 (electronic). URL http://adsabs.harvard.edu/abs/
1968Natur.219..765G.

Gamow:1968:NCN

[Gam68c] George Gamow. Numerology of the constants of nature. Proceedings of the National Academy of Sciences of the United States of America, 59(2):313–318, February 15, 1968. CODEN PNASA6. ISSN 0027-8424 (print), 1091-6490 (electronic). URL http://www.

jstor.org/stable/pdfplus/58638.pdf.

Gamow:1968:OPH

[Gam68d] George Gamow. Observational properties of the homogeneous and isotropic expanding universe. Physical Review Letters, 20 (23):1310–1312, June 3, 1968. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145. URL http:/

/adsabs.harvard.edu/abs/1968PhRvL..20.1310G;   http://
link.aps.org/doi/10.1103/PhysRevLett.20.1310.

Gamow:1968:TAR

[Gam68e] George Gamow. Tredive år der rystede fysikken, Kvanteteoriens historie. (Danish) [Thirty years that shook physics; the story of quantum theory], volume 204 of Gyldendals uglebøger. Gyldendal, Oslo, Norway, 1968. ISBN ???? 199 pp. LCCN ????

Gamow:1968:TAQ

[Gam68f] George Gamow. Trente années qui ébranlèrent la physique: histoire de la physique quantique. (French) [Thirty years that shook physics; the story of quantum theory]. Dunod, Paris, France, 1968. x + 194 pp. LCCN ???? Translated to French by Geneviève Guéron.

Gamow:1969:EUP

[Gam69a] George Gamow. Erde, unser Planet (German) [Earth, our planet]. Franz Ehrenwirth Verlag, München, West Germany, 1969. 264

(est.) pp. LCCN ???? URL http://adsabs.harvard.edu/abs/
1969eup..book.....G.

Gamow:1969:ORM

[Gam69b] George Gamow. Obituary: The Rev. Martin Davidson. Irish Astronomical Journal, 9:170, December 1969. CODEN IRAJAW. ISSN

0021-1052. URL http://adsabs.harvard.edu/abs/1969IrAJ..
..9..170..

Gamow:1969:OG

[Gam69c] George Gamow. On the origin of galaxies. In Mark and Fernbach [MF69], pages 11–22. ISBN 0-470-56990-5. LCCN Q171 .P98. URL

http://adsabs.harvard.edu/abs/1969pmuc.book...11G.

Gamow:196x:ACE

[Gam6x] George Gamow. Astronomy on Christmas Eve. Boy’s Life, ??(??): ??, 196x. CODEN ???? ISSN ????

Gamow:1970:MWL

[Gam70] George Gamow. My World Line: An Informal Autobiography. Viking Press, New York, NY, USA, 1970. 178 pp. LCCN ???? URL

http://adsabs.harvard.edu/abs/1970mwla.book.....G. Fore-

word by Stanisław M. Ulam.

Gamow:1972:TYS

[Gam72] George Gamow. Thirty Years That Shook Physics: the Story of Quantum Theory, volume 38 of Science study series. Heinemann Educational, London, UK, 1972. ISBN 0-435-55071-3. viii + 224 + 12 pp. LCCN QC173.98 .G35 1972.

Gamow:1975:ORN

[Gam75a] George Gamow. 6. the origin of γ-rays and the nuclear energy levels. In Mehra [Meh75], chapter 8, pages 220–222. ISBN 90-277-0635-2. LCCN QC1.S792 M43.

Gamow:1975:TYS

[Gam75b] George Gamow. Thirty Years That Shook Physics. Chonpa Kwahaksa, Seoul, Korea, 1975. ISBN ???? 271 pp. LCCN ???? Translation to Korean by Chong-hum Kim.

Gamow:1977:OTT

[Gam77] George Gamow. One, Two, Three, …, Infinity: Facts and Speculations of Science. Penguin, New York, NY, USA, 1977. ISBN 0-14-004666-6. ???? pp. LCCN Q162 .G23 1977.

Gamow:1980:MIR

[Gam80] George Gamow. M(iste)r Tompkins seltsame Reisen durch Kosmos und Mikrokosmos. (German) [Mr. Tompkins’ strange journey through the cosmos and microcosmos]. Friedrich Vieweg und Sohn, Braunschweig, Germany, 1980. ISBN 3-528-08419-7. xii + 182 pp. LCCN ????

Gamow:1981:CPG

[Gam81] George Gamow. Combinatorial principles in genetics. In Beckenbach and Pólya [BP81], chapter 17, pages 515–?? ISBN 0-89874-172-6. LCCN QA164 .B4 1981.

Gamow:1985:TYS

[Gam85] George Gamow. Thirty Years That Shook Physics: the Story of Quantum Theory. Dover, New York, NY, USA, 1985. ISBN 0-486-24895-X (paperback). xiv + 224 + 9 pp. LCCN QC174.12

.G35 1985. US$4.95. URL http://www.loc.gov/catdir/
description/dover032/85006797.html.

Gamow:1986:OES

[Gam86] George Gamow. The origin of elements and the separation of galaxies. In Bernstein and Feinberg [BF86], pages 114–116. ISBN 0-231-06376-8. LCCN QB985 .C66 1986. URL http://adsabs.

harvard.edu/abs/1986coco.conf..114G.

Gamow:1988:GPG

[Gam88a] George Gamow. The Great Physicists from Galileo to Einstein. Dover, New York, NY, USA, 1988. ISBN 0-486-25767-3. vi + 338 pp. LCCN QC7 .G27 1988; QC7 .G14b 1988. URL http://www.

loc.gov/catdir/description/dover032/88017677.html.

Gamow:1988:OTT

[Gam88b] George Gamow. One, Two, Three, …, Infinity: Facts and Speculations of Science. Dover, New York, NY, USA, 1988. ISBN 0-486-25664-2 (paperback). xii + 340 pp. LCCN Q162

.G23 1988. US$6.95. URL http://www.loc.gov/catdir/
description/dover032/88018955.html.

Gamow:1990:MC

[Gam90] G. Gamow. Modern cosmology. In Leslie [Les90], pages 51–?? ISBN 0-02-370021-1. LCCN QB981 .P563 1990. URL http://

adsabs.harvard.edu/abs/1990pcp..book...51G.

Gamow:1991:GTE

[Gam91] George Gamow. Gay tribe of electrons. In Ferris and Fadiman [FF91], pages 38–49. ISBN 0-316-28129-8. LCCN QC71 .W67 1991. Foreword by Clifton Fadiman.

Gamow:1993:MTP

[Gam93a] George Gamow. Mr. Tompkins in Paperback. Cambridge University Press, Cambridge, UK, canto edition, 1993. ISBN 0-521-44771-2 (paperback). xvi + 185 pp. LCCN QC71 .G25 1993. URL http:/

/adsabs.harvard.edu/abs/1993mtp..book.....G;   http://
www.loc.gov/catdir/description/cam026/93199164.html;
http://www.loc.gov/catdir/toc/cam028/93199164.html.

Gamow:1993:ODT

[Gam93b] George Gamow. An outline of the development of the theory of the structure of the atomic nucleus (I, IV). Physics-Uspekhi, 36(4):

267–278, April 1993. CODEN PHUSEY. ISSN 1063-7869 (print),

1468-4780 (electronic). URL http://ufn.ru/en/articles/1993/
4/e/.

Gamow:1993:MML

[Gam93c] Georgii Gamow. Moya mirovay a liniya. (Russian) [My world line]. Kodry, 8(??):139–??, ???? 1993. CODEN ???? ISSN ????

Gamow:1994:PMT

[Gam94] George Gamow. Priklyucheniy a Mistera Tompkinsa. (Russian) [The Adventures of Mr. Tompkins]. Byuro Kvantum, Moscow, Russia, 1994. ISSN ???? ???? pp. LCCN ????

Gamow:2001:TCS

[Gam01a] George Gamow. Trent’anni che sconvolsero la fisica: la storia della Teoria dei Quanti. (Italian) [Thirty years that shook physics; the story of quantum theory]. Zanichelli, Bologna, Italy, 2001. ISBN ???? 205 pp. LCCN ????

Gamow:2001:TAQ

[Gam01b] George Gamow. Trente années qui ébranlèrent la physique: histoire de la physique quantique. (French) [Thirty years that shook physics; the story of quantum theory]. Dunod, Paris, France, 2001. ISBN ???? x + 194 pp. LCCN ???? Translated to French by J. Gabay.

Gamow:2002:G

[Gam02] George Gamow. Gravity. Dover, New York, NY, USA, dover edition, 2002. ISBN 0-486-42563-0 (paperback). 157 pp. LCCN QC178

.G3 2002. URL http://www.loc.gov/catdir/description/
dover033/2002034827.html.

Gamow:2004:CU

[Gam04a] George Gamow. The Creation of the Universe. Dover science books. Dover, New York, NY, USA, 2004. ISBN 0-486-43868-6. xii + 146 pp. LCCN QB981 .G3 2004.

Gamow:2004:GF

[Gam04b] George Gamow. Galaxies in flight. Resonance, 9(7):93–100, July 2004. CODEN RESOFE. ISSN 0971-8044 (print), 0973-712X (elec-

tronic). URL http://www.ias.ac.in/resonance/July2004/
pdf/July2004Classics.pdf;   http://www.springerlink.com/
content/0971-8044/9/7/.

Gamow:2005:BDS

[Gam05] George Gamow. The Birth and Death of the Sun: Stellar Evolution and Subatomic Energy. Dover, New York, NY, USA, 2005. ISBN 0-486-44231-4 (paperback). 219 pp. LCCN QB45.2

.G36 2005. URL http://www.loc.gov/catdir/enhancements/
fy0624/2005043294-d.html.

Gamow:2007:QU

[Gam07] George Gamow. Quantum uncertainty. In Whitfield and Hicks [WH07], page ?? ISBN 1-880323-91-5. LCCN QC7.5 .W53 2007.

URL http://store.greatbooks.org/adu-wtm.html. Foreword

by Alan Lightman.

Gamow:2010:GFC

[Gam10] George Gamow. Gravità: la forza che governa l’universo. (Italian) [Gravity: the force that governs the Universe]. Dedalo, Bari, Italy, 2010. ISBN 88-220-0248-2. 152 pp. LCCN ???? Translated to Italian by Andrea Migliori. Introduction by Elena Ioli. Foreword by Gino Segrè. Postscript by Silvio Bergia.

Gamow:2011:AEC

[Gam11a] George Gamow. Atomic Energy in Cosmic and Human Life: Fifty Years of Radioactivity. Cambridge University Press, Cambridge, UK, 2011. ISBN 1-107-40208-5. 161 pp. LCCN ????

Gamow:2011:MTL

[Gam11b] George Gamow. Mr Tompkins Learns the Facts of Life. Cambridge University Press, Cambridge, UK, 2011. ISBN 1-107-40207-7 (paperback). xi + 87 + 6 pp. LCCN ????

Gamow:2012:MTP

[Gam12] George Gamow. Mr Tompkins in Paperback. Cambridge University Press, Cambridge, UK, 2012. ISBN 1-107-60468-0. Reprint of the 1993 edition.

Gamow:2014:BF

[Gam14] George Gamow. Biografía de la física. (Spanish) [Biography of Physics], volume C34 of El Libro de bolsillo. Ciencias. Alianza Editorial, Madrid, Spain, 2014. ISBN 84-206-8970-X. 486 + 12 pp. LCCN ???? Spanish translation by Fernando Vela.

Gamow:1941:NVS

[Gam1 ] George Gamow. Neutrinos vs. supernovæ. The Scientific Monthly, 54(??):??, ???? 1941 (??). CODEN SCMOAA. ISSN 0096-3771 (print), 2327-7513 (electronic).

Garrity:2007:GGL

[Gar07] Peter Garrity. The galloping Gamows: In living technicolor. Documentary film (77 minutes), 2007. URL http://www.imdb.com/

title/tt1578747/;   http://wwwbooksurge.com/.

Gamow:1968:BTM

[GB68] George Gamow and Richard Blade. Basic theories in modern physics. Listed in [HPA97a]. The book was still unfinished when Gamow died., 1968.

Gamow:1948:GSP

[GBK48] George Gamow, J. Belzer, and J. Keller. General stabilityproblems of atomic nuclei. In ????, editor, Proceedings of the Scientific Computation Forum, pages 67–69. IBM Corporation, New York, NY, USA, 1948. LCCN ????

Gurney:1928:WMR

[GC28] Ronald W. Gurney and Edward U. Condon. Wave mechanics and radioactive disintegration. Nature, 122(3073):439, September 22, 1928. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687

(electronic). URL http://www.nature.com/nature/journal/
v122/n3073/pdf/122439a0.pdf. See also [Gam28b, GH29, GC29,

Alp73].

Gurney:1929:QMR

[GC29] R. W. Gurney and E. U. Condon. Quantum mechanics and radioactive disintegration. Physical Review (2), 33(2):127–140, February 1929. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (elec-

tronic). URL http://link.aps.org/doi/10.1103/PhysRev.33.
127.

Gamow:1949:TAN

[GC49] G. Gamow and C. L. Critchfield. Theory of Atomic Nucleus and Nuclear Energy-sources. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, third edition, 1949.

Gamow:1960:PFF

[GC60] George Gamow and John M. Cleveland. Physics: Foundations and Frontiers. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1960. 551 pp. LCCN QC23 .G195.

Gamow:1969:PFF

[GC69] George Gamow and John M. Cleveland. Physics: Foundations and Frontiers. Prentice-Hall, Upper Saddle River, NJ 07458, USA, second edition, 1969. ISBN 0-13-672451-5. xvi + 591 pp. LCCN QC23 .G195 1969.

Gamow:1971:M

[GC71] George Gamow and Hal Clement. The Moon. Abelard-Schuman, London, UK, revised edition, 1971. ISBN 0-200-71761-8. 125 pp.

LCCN QB581 .G3 1971. URL http://adsabs.harvard.edu/abs/
1971QB581.G3........

Gamow:1976:PFF

[GC76] George Gamow and John M. Cleveland. Physics: Foundations and Frontiers. Prentice-Hall, Upper Saddle River, NJ 07458, USA, third edition, 1976. ISBN 0-13-672535-X. xvi + 605 pp. LCCN QC23 .G195 1976. US$12.95.

Gamow:19xx:SSH

[GDWWxx] George Gamow, Charles Darwin, H. G. (Herbert George) Wells, and G. P. (George Philip) Wells. Sengo shuppan honyaku ippan kagaku tosho. Kindai Kagakusha, [etc.], Tōkyō, 19xx. various (37 volumes) pp. LCCN CLC Q158 nos. 1-37.

Gamow:1942:EAW

[GF42] George Gamow and J. A. Fleming. The Eighth Annual Washington Conference of Theoretical Physics. Science, 95(2475):579– 581, June 5, 1942. CODEN SCIEAS. ISSN 0036-8075 (print),

1095-9203 (electronic). URL http://www.jstor.org/stable/
pdfplus/1667728.pdf.

Gamow:1976:GGB

[GG76] George Gamow and Barbara Gamow. George Gamow and Barbara Gamow papers, 1915–1975, 1976. URL http://hdl.loc.

gov/loc.mss/eadmss.ms010191;   http://hdl.loc.gov/loc.
mss/eadmss.ms010191.3.

Gamow:1929:QRK

[GH29] George Gamow and Fritz G. Houtermans. Zur Quantenmechanik des radioaktiven Kerns. (German) [On the quantum mechanics of the radioactive nucleus]. Zeitschrift für Physik, 52(7–8):496–509, July 1929. CODEN ZEPYAA. ISSN 0044-3328. URL http://

www.springerlink.com/content/w44101m41820l7p8/. See also

[Gam28b, GC28, Alp73].

Gamow:1932:BAR

[GH32] George Gamow and Fritz Houtermans. Der Bau des Atomkerns und die Radioaktivität. (German) [The structure of atomic nuclei and radioactivity], volume 1 of Neue Probleme der Physik un Chemie. S. Hirzel, Leipzig, Germany, 1932. 5 + 147 + 1 pp. LCCN QC173 .G314.

Gamow:1945:NTC

[GH45] G. Gamow and J. A. Hynek. A new theory by C. F. von Weizsäcker of the origin of the planetary system. Astrophysical Journal, 101: 249–254, March 1945. CODEN ASJOAB. ISSN 0004-637X (print),

1538-4357 (electronic). URL http://adsabs.harvard.edu/abs/
1945ApJ...101..249G.

Gamow:1947:RTM

[GHJ47] George Gamow, Paul R. Heyl, and James J. Jelinek. Review: They move in the void:: Meet the Atoms by O. R. Frisch, The Atomic Story by John W. Campbell, Explaining the Atom by Selig Hecht. The Scientific Monthly, 65(2):165–167, August 1947. CO-DEN SCMOAA. ISSN 0096-3771 (print), 2327-7513 (electronic).

URL http://www.jstor.org/stable/pdfplus/19016.pdf.

Gamow:1926:WMG

[GI26] George Gamow and Dmitri Iwanenko. Zur Wellentheorie der Materie. (German) [On the wave theory of matter]. Zeitschrift für Physik, 39(10–11):865–868, October 1926. CODEN ZEPYAA.

ISSN 0044-3328. URL http://www.springerlink.com/content/
h8170266rv350j66/.

Gamow:1928:WCL

[GIL28] Georgii Antonovich Gamow, Dmitrii Dmitrievich Ivanenko, and Lev Davidovich Landau. World constants and limiting transition (Russian). Zhurnal Russkogo Fiziko-Khimicheskogo Obshchestva, Chast Fizicheskaya, 60(1A):13–17, ???? 1928. CODEN ZRFFAS.

ISSN 0372-9834. Reprinted in [GIL02a, GIL02b]. See comments about this little-known paper in [Oku02]. The journal table of contents gives the German translation of the title as Über die Weltkonstanten und den Grenzübergang.

Gamow:2002:UCB

[GIL02a] George Gamow, D. Ivanenko, and L. Landau. Universal constants and boundary crossings (Russian). Yadernaya Fizika, 65(7):1406– 1408, ???? 2002. CODEN IDFZA7. ISSN 0044-0027. Reprint of [GIL28].

Gamow:2002:WCL

[GIL02b] George Gamow, D. Ivanenko, and L. Landau. World constants and limiting transition. Physics of Atomic Nuclei, 65(7): 1373–1375, 2002. CODEN PANUEO. ISSN 1063-7788 (print),

1562-692X (electronic). URL http://www.springerlink.com/

content/0881529622k1t658/. English translation and reprint of Russian original [GIL28, GIL02a]. See comments [Oku02].

Gilmore:2012:PSS

[Gil12] Gerard Gilmore, editor. Planets, Stars and Stellar Systems: Solar and stellar planetary systems, volume 5. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2012. ISBN 94-007-5612-7. LCCN QB460-466; QB857 .P53 2013.

Gingerich:1994:SWA

[Gin94] Owen Gingerich. The summer of 1953: a watershed for astrophysics. Physics Today, 47(12):34–40, December 1994. CO-DEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

URL http://www.physicstoday.org/resource/1/PHTOAD/v47/

i12. Contains several mentions and anecdotes of George Gamow.

Gamow:1945:SSM

[GK45] George Gamow and G. Keller. A shell source model for Red Giant stars. Reviews of Modern Physics, 17(2–3):125–137, April 1945. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://adsabs.harvard.edu/
abs/1945RvMP...17..125G;   http://link.aps.org/doi/10.
1103/RevModPhys.17.125;   http://rmp.aps.org/abstract/
RMP/v17/i2-3/p125_1.

Gamow:1933:ITS

[GL33] George Gamow and Lev Landau. Internal temperature of stars. Nature, 132(3336):567, October 7, 1933. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://adsabs.

harvard.edu/abs/1933Natur.132..567G.

Gamow:1950:PRG

[GL50] George Gamow and C. Longmire. The problem of Red Giants and Cepheid variables. Physical Review, 79(4):743–744, August 15, 1950. CODEN PHRVAO. ISSN 0031-899X (print), 1536-

6065 (electronic). URL http://prola.aps.org/abstract/PR/
v79/i4/p743_1.

Glass:1949:RBE

[Gla49] Bentley Glass. Review:: Biography of the Earth. Its Past, Present and Future by George Gamow. Quarterly Review of Biology, 24 (4):354–??, December 1949. CODEN QRBIAK. ISSN 0033-5770 (print), 1539-7718 (electronic). URL http://www.jstor.org/

stable/pdfplus/2809601.pdf.

Glass:1952:RBT

[Gla52] Bentley Glass. Review: One Two Three… Infinity. Facts and Speculations of Science by George Gamow. Quarterly Review of Biology, 27(1):135–??, March 1952. CODEN QRBIAK. ISSN 0033-5770 (print), 1539-7718 (electronic). URL http://www.jstor.

org/stable/pdfplus/2812841.pdf.

Gamow:1926:MNS

[GLI26] George Gamow, Lev D. Landau, and Dimitri D. Ivanenko. On the motion of nonconservative systems with one degree of freedom. (Russian). Zhurnal Russkogo Fiziko-Khimicheskogo Obshchestva, Chast Fizicheskaya, 58(3A):477–482, ???? 1926. CODEN ZRF-FAS. ISSN 0372-9834.

Gamow:1954:NPC

[GM54] George Gamow and N. Metropolis. Numerology of polypeptide chains. Science (New Series), 120(3124):779–780, November 12, 1954. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http:

//www.jstor.org/stable/1681634.

Grib:2000:GRD

[GN00] A. Grib and M. Novello. Gamow in Rio and the discovery of the URCA process. Astronomical and Astrophysical Transactions, 19:669–673, 2000. CODEN AATREG. ISSN 1055-6796 (print),

1476-3540 (electronic). URL http://adsabs.harvard.edu/abs/
2000A%26AT...19..669G.

Gamow:1997:PDR

[GNF^{+}97] I. Gamow, M. Nirenberg, J. Follin, V. Rubin, A. Rich, R. Alpher, R. Herman, and P. Abelson. Panel discussion and reminiscences of George Gamow. In Harper et al. [HPA97b], pages 127–?? ISBN 1-886733-49-X. LCCN QC16.G37 G46 1996. URL http://adsabs.

harvard.edu/abs/1997ASPC..129..127G.

Gamow:2006:MTG

[GO06] George Gamow and Robert Oerter, editors. Mr. Tompkins Gets Serious: the Essential George Gamow. Pi Press, New York, NY, USA, the masterpiece science edition, 2006. ISBN 0-13-187291-5. xvi + 399 pp. LCCN QC71 .G26 2006. URL http://www.

loc.gov/catdir/toc/ecip0518/2005025070.html. Foreword by

R. Igor Gamow.

Gorelik:1990:PFG

[Gor90] G. E. Gorelik. Predystoriya FIANa. G. A. Gamow. (Russian) [The pre-history of the Physical Institute of the Academy of Sciences. G. A. Gamow]. HE Preprint 41, ????, Moscow, Russia, 1990.

Gamow:1931:DVO

[GR31] George Gamow and Léon Rosenfeld. On the determination of the velocity of an object moving in a fluid on the basis of a single photograph. Originally written in German, and reproduced in English translation in [Del72, pages 285–287]. Submitted to the journal Physica, but rejected by editor Paul Ehrenfest., June 7, 1931.

Gamow:1933:DEN

[GR33] George Gamow and S. Rosenblum. Les diamètres effectifs des noyaux radioactifs. (French) [The effective diameters of radioactive nuclei]. Comptes Rendus des Séances de L’Académie des Sciences, 197(??):1620–1622, December 18, 1933. CODEN ???? ISSN ????

Graetzer:1964:DNF

[Gra64] Hans G. Graetzer. Discovery of nuclear fission. American Journal of Physics, 32(9):9–15, January 1964. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). See [HS39, Bey49].

Greenstein:1990:MPG

[Gre90] G. Greenstein. The magician + a profile of George Gamow. American Scholar, 59(1):118–125, Winter 1990. ISSN 0003-0937 (print), 2162-2892 (electronic).

Greenberger:2000:BRG

[Gre00] Daniel M. Greenberger. Book review: George Gamow and Russell Stannard, The New World of Mr. Tompkins. Physics Today, 53(6):57, June 2000. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://www.physicstoday.

org/resource/1/phtoad/v53/i6/p57_s1.

Gamow:1956:PITb

[GRY56a] G. Gamow, A. Rich, and M. Ycas. The problem of information transfer from the nucleic acids to proteins. Advances in biological and medical physics, 4(??):23–68, ???? 1956. ISSN 0065-2245.

Gamow:1956:PITa

[GRY56b] George Gamow, Alexander Rich, and Martynas Yčas. The problem of information transfer from nucleic acids to proteins. In Lawrence and Tobias [LT56], pages 23–68.

Gamow:1940:PRN

[GS40] George Gamow and M. Schoenberg. The possible role of neutrinos in stellar evolution. Physical Review, 58(12):1117, December 15, 1940. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (elec-

tronic). URL http://adsabs.harvard.edu/abs/1940PhRv..
.58.1117G;   http://prola.aps.org/abstract/PR/v58/i12/
p1117_1.

Gamow:1941:NTS

[GS41] George Gamow and M. Schoenberg. Neutrino theory of stellar collapse. Physical Review, 59(7):539–547, April 1, 1941. CO-DEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

URL http://adsabs.harvard.edu/abs/1941PhRv...59..539G;
 http://prola.aps.org/abstract/PR/v59/i7/p539_1.

Gamow:1942:NMD

[GS42] George Gamow and Ernesto R. Sábato. Nacimiento y muerte del sol: evolución estelar y energía intraatómica. (Spanish) [The Birth and Death of the Sun: Stellar Evolution and Subatomic Energy]. Esapsa-Calpe Argentina, S.A., Buenos Aires y México, 1942. 257

+ 1 pp. LCCN QB44 .G265. URL http://adsabs.harvard.edu/
abs/1942QB44.G265.......

Gamow:1946:NTS

[GS46] G. Gamow and M. Schoenberg. Neutrino theory of stellar collapse. Publications of the American Astronomical Society, 10:126–130, 1946. CODEN ???? ISSN ???? URL http://adsabs.harvard.

edu/abs/1946PAAS...10..126G.

Gamow:1958:PMa

[GS58a] George Gamow and Marvin Stern. Puzzle-math. Viking Press, New York, NY, USA, 1958. 119 pp. LCCN GV1493 .G3.

Gamow:1958:PMb

[GS58b] George Gamow and Marvin Stern. Puzzle-math. Macmillan Publishing Company, New York, NY, USA, 1958. 119 pp. LCCN GV1493 .G3.

Gamow:1967:JMQ

[GS67] G. Gamow and M. Stern. Jeux mathématiques. Quelques cassetête. (French) [Puzzle-math]. Dunod, Paris, France, 1967.

Gamow:1936:SRD

[GT36] George Gamow and Edward Teller. Selection rules for the βdisintegration. Physical Review (2), 49(12):895–899, June 15, 1936. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (elec-

tronic). URL http://prola.aps.org/abstract/PR/v49/i12/
p895_1.

Gamow:1937:LES

[GT37] George Gamow and Edward Teller. Letter to editor: Some generalizations of the β-transformation theory. Physical Review, 51 (4):289, February 15, 1937. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://prola.aps.

org/abstract/PR/v51/i4/p289_1.

Gamow:1938:LER

[GT38a] George Gamow and Edward Teller. Letter to editor: The rate of selective thermonuclear reactions. Physical Review, 53(7):608– 609, April 1, 1938. CODEN PHRVAO. ISSN 0031-899X (print),

1536-6065 (electronic). URL http://prola.aps.org/abstract/
PR/v53/i7/p608_1.

Gamow:1938:NCS

[GT38b] George Gamow and Edward Teller. On the neutron core of stars. Physical Review, 53(11):929–930, June 1938. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://

prola.aps.org/vtoc/PR/v53/i11/p913_1.

Gamow:1939:EUOa

[GT39a] George Gamow and Edward Teller. The expanding universe and the origin of the Great Nebulæ. Nature, 143(3612):116–117, January 21, 1939. CODEN NATUAS. ISSN 0028-0836 (print),

1476-4687 (electronic). URL http://adsabs.harvard.edu/abs/
1939Natur.143..116G;   http://www.nature.com/nature/journal/
v143/n3612/pdf/143116a0.pdf.

Gamow:1939:EUOb

[GT39b] George Gamow and Edward Teller. The expanding universe and the origin of the Great Nebulæ. Nature, 143(3618):375, March 4, 1939. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (elec-

tronic). URL http://adsabs.harvard.edu/abs/1939Natur.
143..375G;   http://www.nature.com/nature/journal/v143/
n3618/pdf/143375a0.pdf.

Gamow:1939:LEE

[GT39c] George Gamow and Edward Teller. Letter to editor: Energy production in Red Giants. Physical Review, 55(8):791, April 15, 1939. CODEN PHRVAO. ISSN 0031-899X (print), 1536-

6065 (electronic). URL http://prola.aps.org/abstract/PR/
v55/i8/p791_1.

Gamow:1939:OGN

[GT39d] George Gamow and Edward Teller. On the origin of Great Nebulæ. Physical Review, 55(7):654–657, April 1, 1939. CO-DEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

URL http://adsabs.harvard.edu/abs/1939PhRv...55..654G;
 http://prola.aps.org/abstract/PR/v55/i7/p654_1.

Gamow:1956:EZD

[GT56] George Gamow and Walter Theimer. Eins, zwei, drei … Unendlichkeit: Grenzfragen d. modernen Wissenschaft verständlich gemacht. (German) [One, Two, Three, …, Infinity: Facts and Speculations of Science]. Fackelträger-Verlag Schmidt-Küster, Hannover, West Germany, 1956. 286 pp. LCCN ????

Gamow:1958:EZD

[GT58] George Gamow and Walter Theimer. Eins, zwei, drei … Unendlichkeit: Grenzfragen d. modernen Wissenschaft verständlich dargest. (German) [One, Two, Three, …, Infinity: Facts and Speculations of Science], volume 493/494 of Goldmanns gelbe Taschenbücher. Wilhelm Goldmann, München, West Germany, 1958. 318 + 16 pp. LCCN ????

Gamow:1955:SCP

[GY55] George Gamow and Martynas Yčas. Statistical correlation of protein and ribonucleic acid composition. Proceedings of the National Academy of Sciences of the United States of America, 41(12): 1011–1019, December 15, 1955. CODEN PNASA6. ISSN 0027-8424 (print), 1091-6490 (electronic). URL http://www.jstor.

org/stable/pdfplus/89096.pdf. See errata [GY56].

Gamow:1956:ESC

[GY56] George Gamow and Martynas Yčas. Errata: Statistical correlation of protein and ribonucleic acid composition. Proceedings of the National Academy of Sciences of the United States of America, 42(3):172, March 15, 1956. CODEN PNASA6. ISSN 0027-8424 (print), 1091-6490 (electronic). URL http://www.jstor.

org/stable/pdfplus/88924.pdf. See [GY55].

Gamow:1958:CAP

[GY58] George Gamow and Martynas Yças. The cryptographic approach to the problem of protein synthesis. In Das Universum. Unser Bild vom Weltall. (German) [The Universe. Our picture of the Universe], page ???? ????, Wiesbaden, Germany, 1958.

Gamow:1967:MTI

[GY67] George Gamow and Martynas Yčas. Mr. Tompkins Inside Himself: Adventures in the New Biology. Viking Press, New York, NY, USA, 1967. xiv + 274 pp. LCCN QP38 .G3 1967.

Hartree:1932:RBC

[Har32] D. R. Hartree. Review: The Constitution of Atomic Nuclei and Radioactivity by George Gamow. Mathematical Gazette, 16(220): 284–285, October 1932. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL http://www.jstor.org/

stable/pdfplus/3605938.pdf.

Harper:2001:AGG

[Har01] Eamon Harper. In appreciation: George Gamow: Scientific amateur and polymath. Physics in Perspective (PIP), 3(3):335–372, September 2001. CODEN PHPEF2. ISSN 1422-6944 (print),

1422-6960 (electronic). URL http://adsabs.harvard.edu/abs/
2001PhP.....3..335H;   http://link.springer.com/article/
10.1007/PL00000536;   http://www.springerlink.com/content/
m0htmt9ww9b9jjv1/;   http://www.springerlink.com/openurl.
asp?genre=journal&issn=1422-6944.

Harwit:2007:ORA

[Har07] Martin Harwit. Obituary: Ralph Asher Alpher. Physics Today, 60 (12):67–68, December 2007. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://www.physicstoday.

org/resource/1/phtoad/v60/i12/p67_s1.

Hawking:2011:DSM

[Haw11] Stephen Hawking, editor. The dreams that stuff is made of: the most astounding papers on quantum physics — and how they shook the scientific world. Running Press, Philadelphia, PA, USA, 2011. ISBN 0-7624-3434-1. xi + 1071 pp. LCCN QC173.98 .D74 2011.

Hayes:1998:CSI

[Hay98] Brian Hayes. Computing science: The invention of the genetic code. American Scientist, 86(1):8–14, January/February 1998. CODEN AMSCAC. ISSN 0003-0996 (print), 1545-2786 (elec-

tronic). URL http://www.jstor.org/stable/27856930. This ar-

ticle, by a former editor of this journal, describes George Gamow’s important contributions to the understanding of the genetic code in DNA.

Heisenberg:1934:CTG

[Hei34] Werner Heisenberg. Considérations théoriques générales sure la structure du noyau. (French) [General theoretical considerations of the structure of the nucleus]. In Cockcroft et al. [CCJ^{+}34], pages 289–335. LCCN ???? Publiés par la commission administrative de l’institut.

Heniser:1963:RBG

[Hen63] Virgil Heniser. Review: Gravity by George Gamow. The American Biology Teacher, 25(3):215, March 1963. ISSN 0002-7685 (print), 0002-8444 (electronic), 1938-4211, 1938-422X. URL http://www.

jstor.org/stable/pdfplus/4440302.pdf.

Herzfeld:1966:RQT

[Her66] Karl F. Herzfeld. Review: Quantum theory, 1900 to 1930: Thirty Years That Shook Physics by George Gamow. Science, 152(3724): 917–918, May 13, 1966. CODEN SCIEAS. ISSN 0036-8075 (print),

1095-9203 (electronic). URL http://www.jstor.org/stable/
pdfplus/1718546.pdf.

Harmon:2007:SLG

[HG07] Joseph E. Harmon and Alan G. Gross, editors. The scientific literature: a guided tour. University of Chicago Press, Chicago, IL, USA, 2007. ISBN 0-226-31655-6 (hardcover), 0-226-31656-4 (paperback). xxiv + 327 pp. LCCN Q225.5 .S35 2007. URL http:/

/catdir.loc.gov/catdir/enhancements/fy0701/2006016547-
t.html;   http://catdir.loc.gov/catdir/enhancements/fy0707/
2006016547-b.html;   http://catdir.loc.gov/catdir/enhancements/
fy0707/2006016547-d.html.

Hoyle:1972:CIP

[HN72] Fred Hoyle and J. V. Narlikar. Conformal invariance in physics and cosmology. In Reines [Rei72a], pages 15–28. ISBN 0-87081-

025-1. LCCN QC780 .C65. URL http://adsabs.harvard.edu/
abs/1972cht..conf...15H.

Hobson:2002:BRG

[Hob02] Art Hobson. Book review: George Gamow and Russell Stannard, The New World of Mr Tompkins. Cambridge, United Kingdom: Cambridge University Press, 1999, ix + 258 pages, $24.95 (cloth), $16.95 (paper). Physics in Perspective (PIP), 4(4):494–495, December 2002. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Hookham:1993:BRB

[Hoo93] J. Hookham. Books Received — Mr. Tompkins in Paperback, by George Gamow. Science, 261(??):501–??, July 1993. CODEN SCI-EAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http:

//adsabs.harvard.edu/abs/1993Sci...261..501G.

Houtermans:1930:NAQ

[Hou30] Fritz G. Houtermans. Neuere Arbeiten über Quantentheorie des Atomskerns. (German) [New work on the quantum theory of the atomic nucleus]. Ergebnisse der Exakten Naturwissenschaften, 9 (??):123–221, ???? 1930. CODEN EENAA3. ISSN 0367-0325.

Hoyle:1946:SEH

[Hoy46] Fred Hoyle. The synthesis of the elements from hydrogen. Monthly Notices of the Royal Astronomical Society, 106(5):343– 383, ???? 1946. CODEN MNRAA4. ISSN 0035-8711 (print),

1365-2966 (electronic). URL http://adsabs.harvard.edu/abs/
1946MNRAS.106..343H.

Hoyle:1954:NRO

[Hoy54] Fred Hoyle. On nuclear reactions occurring in very hot stars. I. The synthesis of elements from carbon to nickel. Astrophysical Journal. Supplement Series, 1(??):121–146, September 1954. CO-DEN APJSA2. ISSN 0067-0049 (print), 1538-4365 (electronic).

URL http://labs.adsabs.harvard.edu/ui/abs/1954ApJS....

1..121H. Reprinted in the centennial edition Astrophysical Journal 525, 571 (1999) with a modern commentary by David Arnett.

Hoyle:1990:AEA

[Hoy90] Fred Hoyle. An assessment of the evidence against the steadystate theory. In Bertotti et al. [BBBM90], pages 221–232. ISBN 0-521-37213-5. LCCN QB981 .M774 1990. URL http://www.loc.

gov/catdir/description/cam024/90041803.html;   http://
www.loc.gov/catdir/toc/cam028/90041803.html.

Harper:1997:EAG

[HPA97a] E. Harper, W. C. Parke, and G. D. Anderson. Epilogue / appendix: George Gamow’s publications. In Harper et al. [HPA97b], pages 140–151. ISBN 1-886733-49-X. LCCN QC16.G37 G46 1996. URL

http://adsabs.harvard.edu/abs/1997ASPC..129..140H.

Harper:1997:GGS

[HPA97b] Eamon Harper, William Carleton Parke, and G. D. (George David) Anderson, editors. The George Gamow Symposium: sponsored by the George Washington University and the Carnegie Institution of Washington, 12 April 1997 [i.e., 1996], volume 129 of Astronomical Society of the Pacific conference series. Astronomical Society of the Pacific, San Francisco, CA, USA, 1997. ISBN 1-886733-49-X.

LCCN QC16.G37 G46 1996. URL http://adsabs.harvard.edu/
abs/1997ASPC..129.....H.

Hahn:1939:NVB

[HS39] Otto Hahn and Fritz Strassmann. Über den Nachweis und das Verhalten der bei der Bestrahlung des Urans mittels Neutronen entstehenden Erdalkalimetalle. (German) [Concerning the existence of alkaline earth metals resulting from the neutron irradiation of uranium]. Naturwissenschaften, 27(1):11–15, January 1939. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic). A facsimile is also available in [Bey49, pages 87–91] and in [Gra64]. Abridged English translation in [GA71, pages 44–47].

Hayen:2019:BSG

[HS19] L. Hayen and N. Severijns. Beta spectrum generator: High precision allowed β spectrum shapes. Computer Physics Communications, 240(??):152–164, July 2019. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465519300645.

Hockey:2007:BEA

[HTB^{+}07] Thomas A. Hockey, Virginia Trimble, Katherine Bracher, et al., editors. The biographical encyclopedia of astronomers. Springer reference. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2007. ISBN 0-387-31022-3 (set), 0-387-30400-2 (set: electronic version), 0-387-33628-1 (set: print + electronic bundle). xlv + 1341 (two volumes) pp. LCCN QB35

.B565 2007. URL http://www.loc.gov/catdir/enhancements/
fy0814/2008270178-d.html;   http://www.loc.gov/catdir/
enhancements/fy0824/2008270178-t.html.

Hufbauer:2009:GG

[Huf09] Karl Hufbauer. George Gamow: March 4, 1904–August 19, 1968. Biographical memoirs — National Academy of Sciences of the United States of America, ??:1–39, 2009. CODEN BMNSAC. ISSN

0077-2933. URL http://books.nap.edu/html/biomems/ggamow.
pdf.

Hughes:1993:RCC

[Hug93] Jeffrey Alan Hughes. The radioactivists: community, controversy and the rise of nuclear physics. Ph.D. dissertation, University of Cambridge, Cambridge, UK, 1993. viii + 417 pp. URL https://

search.proquest.com/pqdtglobal/docview/301418262.

Hunaerts:1949:GGO

[Hun49] J. Hunaerts. G. Gamow — Het Ontstaan en Vergaan van de Zon. (Dutch) [G. Gamow — The Birth and Death of the Sun]. Ciel et Terre, 65:284–??, 1949. CODEN CIELAV. ISSN 0009-6709. URL

http://adsabs.harvard.edu/abs/1949C%26T....65..284H.

Iggers:1966:RBO

[Igg66] George G. Iggers. Review: he Origins of Modern Consciousness by John Higham, Albert William Levi, Eugen Weber, Roger Shattuck, Benjamin Nelson, Gerhard Masur, George Gamow, John Weiss. The Journal of Modern History, 38(1):105–106, March 1966. ISSN 0022-2801 (print), 1537-5358 (electronic). URL http://www.

jstor.org/stable/pdfplus/1877038.pdf.

Infeld:1948:RMB

[Inf48] Leopold Infeld. Review: Mathematics—and beyond: One Two Three… Infinity by George Gamow. The Scientific Monthly, 66 (5):440–441, May 1948. CODEN SCMOAA. ISSN 0096-3771 (print), 2327-7513 (electronic). URL http://www.jstor.org/

stable/pdfplus/19511.pdf.

Jones:2004:IGC

[JLA^{+}04] Mark H. (Mark Henry) Jones, Robert J. Lambourne, D. J. (David John) Adams, et al., editors. An introduction to galaxies and cosmology. Open University, Milton Keynes, UK, 2004. ISBN 0-521-83738-3, 0-521-54623-0 (paperback). vi + 442 pp. LCCN QB857 .I67 2004. URL http://catdir.loc.

gov/catdir/description/cam051/2004301655.html;   http:
//catdir.loc.gov/catdir/toc/cam051/2004301655.html.

Johnson:1954:BRB

[Joh54a] M. J. Johnson. Book review: Mr Tompkins Learns the Facts of Life, George Gamow. New York: Cambridge Univ. Press, 1953. 88 pp. Illus. $2.75. Science, 119(3080):76, January 8, 1954. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (elec-

tronic). URL http://www.sciencemag.org/content/119/3080/
76.2.citation.

Johnson:1954:RBT

[Joh54b] M. J. Johnson. Review: Mr Tompkins Learns the Facts of Life by George Gamow. Science, 119(3080):76–??, January 8, 1954. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

URL http://www.jstor.org/stable/pdfplus/1682717.pdf.

Judson:2001:BRH

[Jud01] Horace Freeland Judson. Book review: Honest Jim: the sequel: Further misadventures of one of the most influential scientists of our day.Genes, Girls and Gamow by James D. Watson. Nature, 413 (6858):775–776, October 25, 2001. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.

com/nature/journal/v413/n6858/full/413775a0.html. See ed-

itorial response [The01].

K:1962:RBE

[K.62] P. K. Review: Biography of the Earth by George Gamow. The American Biology Teacher, 24(1):55, January 1962. CODEN ???? ISSN 0002-7685 (print), 0002-8444 (electronic), 1938-4211, 1938-

422X. URL http://www.jstor.org/stable/pdfplus/4439798.
pdf.

Kavanagh:1972:RRP

[Kav72] Ralph W. Kavanagh. Reaction rates in the proton-proton chain. In Reines [Rei72a], pages 169–185. ISBN 0-87081-025-1. LCCN

QC780 .C65. URL http://adsabs.harvard.edu/abs/1972cht..
conf..169K.

Kox:2005:UGR

[KE05] Anne J. Kox and Jean Eisenstaedt, editors. The universe of General Relativity, volume 11 of Einstein studies. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 2005. ISBN 0-8176-4380-X, 0-8176-4454-7 (e-book). x + 383 pp. LCCN QC173.5 .U55 2005. URL http://www.loc.gov/

catdir/enhancements/fy0663/2005047817-d.html;   http:/
/www.loc.gov/catdir/toc/fy0602/2005047817.html.

Klein:1966:RBY

[Kle66] Martin J. Klein. Review: Thirty Years That Shook Physics by George Gamow. American Scientist, 54(3):308A, ???? 1966. CO-DEN AMSCAC. ISSN 0003-0996 (print), 1545-2786 (electronic).

URL http://www.jstor.org/stable/27836485.

Klein:2000:GGF

[Kle00] E. Klein. George Gamow, facetious knowledge. Recherche, ??(332): 60–61, ???? 2000. ISSN 0029-5671 (print), 1625-9955 (electronic).

Klein:2005:ESF

[Kle05] Etienne Klein. Il était sept fois la révolution: Albert Einstein et les autres . . .. (French) [The seven-times revolution: Albert Einstein and the others …]. Flammarion, Paris, France, 2005. ISBN 2-08-210343-9. 237 pp. LCCN QC16.E5.

Katzir:2013:TTH

[KLR13] Shaul Katzir, Christoph Lehner, and Jürgen Renn, editors. Traditions and transformations in the history of quantum physics: HQ-3, Third International Conference on the History of Quantum Physics, Berlin, June 28–July 2, 2010, volume 5 of Max Planck research library for the history and development of knowledge. Proceedings. Edition Open Access, Berlin, Germany, 2013. ISBN 3-8442-5134-0. LCCN QC173.98. URL http://www.edition-open-

access.de/proceedings/5/.

Kragh:1991:CEDb

[Kra91a] Helge S. Kragh. Cosmonumerology and empiricism — the Dirac/ Gamow dialogue. Astronomy Quarterly, 8:109–??, 1991. CODEN ASTQD2. ISSN 0364-9229 (print), 1878-0865 (electronic). URL

http://adsabs.harvard.edu/abs/1991AstQ....8..109K.

Kragh:1991:CEDa

[Kra91b] Helge S. Kragh. Cosmonumerology and empiricism: the Dirac– Gamow dialogue. Astronomy Quarterly, 8:107–??, 1991. CODEN ASTQD2. ISSN 0364-9229 (print), 1878-0865 (electronic). URL

http://adsabs.harvard.edu/abs/1991AstQ....8..107K.

Kragh:1996:CCH

[Kra96a] Helge Kragh. Cosmology and Controversy: the Historical Development of Two Theories of the Universe. Princeton University Press, Princeton, NJ, USA, 1996. ISBN 0-691-02623-8 (hardcover). xiii + 500 pp. LCCN QB981 .K73 1996. URL http://www.loc.

gov/catdir/description/prin031/96005612.html;   http:/
/www.loc.gov/catdir/toc/prin031/96005612.html.

Kragh:1996:GBB

[Kra96b] Helge Kragh. Gamow’s big bang. In Cosmology and Controversy: the Historical Development of Two Theories of the Universe [Kra96a], chapter 3, pages 80–141. ISBN 0-691-02623-8 (hardcover). LCCN QB981 .K73 1996. URL http://www.loc.

gov/catdir/description/prin031/96005612.html;   http:/
/www.loc.gov/catdir/toc/prin031/96005612.html.

Kragh:1996:GGR

[Kra96c] Helge Kragh. Gamow’s game: The road to the Hot Big Bang. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 38(4):335–361, December 1996. CODEN CENTA4. ISSN 0008-8994 (print),

1600-0498 (electronic). URL http://adsabs.harvard.edu/abs/
1996Cent...38..335K;   http://onlinelibrary.wiley.com/
doi/10.1111/j.1600-0498.1996.tb00020.x/abstract.

Kragh:2002:CCS

[Kra02] Helge Kragh. Cosmologies and cosmogonies of space and time. In Nye [Nye02], pages 522–537. ISBN 0-521-57244-4 (vol. 3 : hardback), 0-521-57243-6 (vol. 4), 0-521-57199-5 (vol. 5), 0-521-59442-1 (vol. 7). LCCN Q125 .C32 2003. URL http:

//www.h-net.org/review/hrev-a0d3l3-aa;   http://www.h-
net.org/review/hrev-a0d7k4-aa;   http://www.h-net.org/
review/hrev-a0f4e9-aa;   http://www.loc.gov/catdir/description/
cam021/2001025311.html;   http://www.loc.gov/catdir/enhancements/
fy0731/2001025311-b.html;   http://www.loc.gov/catdir/
samples/cam033/2001025311.html;   http://www.loc.gov/catdir/
toc/cam026/2001025311.html.

Kragh:2005:GGF

[Kra05] Helge Kragh. George Gamow and the ‘factual approach’ to relativistic cosmology. In Kox and Eisenstaedt [KE05], pages 175–

188. ISBN 0-8176-4380-X, 0-8176-4454-7 (e-book). LCCN QC173.5

.U55 2005. URL http://adsabs.harvard.edu/abs/2005ugr..
book..175K.

Kragh:2014:NBB

[Kra14] Helge Kragh. Naming the Big Bang. Historical Studies in the Natural Sciences, 44(1):3–36, February 2014. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic). URL http://www.jstor.

org/stable/10.1525/hsns.2014.44.1.3.

Kragh:2017:NPS

[Kra17] Helge Kragh. The Nobel Prize system and the astronomical sciences. Journal for the History of Astronomy, 48(3):257–280, August 2017. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556

(electronic). URL http://journals.sagepub.com/doi/full/
10.1177/0021828617721574.

Kragh:2018:GLP

[Kra18] Helge Kragh. Georges Lemaître, pioneer of modern theoretical cosmology. Foundations of Physics, 48(10):1333–1348, October 2018. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Kuhn:1967:RTR

[Kuh67] Thomas S. Kuhn. Review: The turn to recent science: The Questioners: Physicists and the Quantum Theory by Barbara Lovett Cline Thirty Years That Shook Physics: The Story of Quantum Theory by George Gamow; The Conceptual Development of Quantum Mechanics by Max Jammer; Korrespondenz, Individualitat, und Komplementaritat: Eine Studie zur Geistesgeschichte der Quantentheorie in den Beitragen Niels Bohrs by Klaus Michael Meyer-Abich; Niels Bohr: The Man, His Science, and the World They Changed by Ruth Moore; Sources of Quantum Mechanics by B. L. Van der Waerden. Isis, 58(3):409–419, ???? 1967. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/228002;   http:
//www.jstor.org/stable/pdfplus/228002.pdf.

LallenaRojo:2014:GBB

[Lal14] Antonio Lallena Rojo. Gamow, el Big Bang: el conocimiento en expansión. (Spanish) [Gamow, the Big Bang: expanding knowledge]. RBA, Barcelona, Spain, 2014. ISBN 84-473-7774-1. 151 pp. LCCN ????

Laster:1962:BRG

[Las62] Howard Laster. Book review: George Gamow, Gravity. American Journal of Physics, 30(10):777, October 1962. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). URL http://ajp.

aapt.org/resource/1/ajpias/v30/i10/p777_s1.

Leslie:1990:PCP

[Les90] John Leslie, editor. Physical Cosmology and Philosophy. Philosophical topics. Macmillan Publishing Company, New York, NY, USA, 1990. ISBN 0-02-370021-1. viii + 277 pp. LCCN QB981 .P563 1990.

Lewis:1934:GE

[Lew34] Gilbert N. Lewis. The genesis of the elements. Physical Review, 46(10):897–901, November 1934. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://link.aps.org/

doi/10.1103/PhysRev.46.897.

Linwood:2000:CGT

[Lin00] D. A. Linwood. The calculus and Gamow’s theory of gravitation. College Mathematics Journal, 31(4):281–285, September 2000. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.1080/
07468342.2000.11974159.

Lipkin:1986:GTA

[Lip86] Harry J. Lipkin. On Gamow’s theory of alpha-decay: A comment. In Ullmann-Margalit [UM86a], chapter 15, pages 187–192. ISBN 90-277-2159-9 (paperback), 90-277-2158-0 (hardcover), 94-009-5496-4 (e-book). ISSN 0068-0346. LCCN Q174 .B67 vol.

94 Q175. URL https://link.springer.com/chapter/10.1007/
978-94-009-5496-0_15. See [Stu86].

Longmire:1972:HCP

[Lon72] Conrad L. Longmire. Heating of charged particles by electric waves. In Reines [Rei72a], pages 232–240. ISBN 0-87081-025-1. LCCN QC780 .C65.

Lawrence:1956:ABM

[LT56] J. H. Lawrence and C. A. Tobias, editors. Advances in Biological and Medical Physics: Volume 4, volume 4. Academic Press, New York, NY, USA, 1956.

Lukas:1970:SPE

[Luk70] Jerome S. Lukas. Solarized de-polariz-ed [sound recording], 1970. 1 sound disc.

Lattes:1946:ANU

[LW46] C. Lattes and G. Wataghin. On the abundance of nuclei in the universe. Physical Review, 69(5–6):237, March 1946. CO-DEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

URL http://link.aps.org/doi/10.1103/PhysRev.69.237. See

remarks in [Dys93] about the relation of this work to [ABG48], and the subsequent incorrect neglect of Wataghin’s work. See also related papers [Wat46, DW48, Wat48].

M:1940:REP

[M.40a] F. R. M. Review: Evolution of the physical universe: The Birth and Death of the Sun by George Gamow. The Scientific Monthly, 51(4):373, October 1940. CODEN SCMOAA. ISSN 0096-3771 (print), 2327-7513 (electronic). URL http:/

/adsabs.harvard.edu/abs/1940SciMo..51..373G;   http://
www.jstor.org/stable/pdfplus/17358.pdf.

M:1940:BRBm

[M.40b] W. M. M. Book review: Mr. Tompkins in Wonderland by G. Gamow. Philosophy of Science, 7(3):386, July 1940. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http:

//www.jstor.org/stable/184854.

Marateck:2008:LEA

[Mar08] Samuel L. Marateck. Letter to the Editor: Alpher, Bethe, Gamow. Physics Today, 61(9):11–12, September 2008. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://www.

physicstoday.org/resource/1/phtoad/v61/i9/p11_s1.

Mather:1942:REB

[Mat42] Kirtley F. Mather. Review: The earth: Biography of the Earth by George Gamow. Science, 95(2455):71–72, January 16, 1942. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

URL http://adsabs.harvard.edu/abs/1942Sci....95...71M;
 http://www.jstor.org/stable/pdfplus/1669073.pdf.

Matthews:1966:RBT

[Mat66] Geoffrey Matthews. Review: Mr. Tompkins in Paperback by George Gamow. Mathematical Gazette, 50(374):405–406, December 1966. CODEN MAGAAS. ISSN 0025-5572 (print),

2056-6328 (electronic). URL http://www.jstor.org/stable/
pdfplus/3613951.pdf.

McCrea:1940:RBT

[McC40] W. H. McCrea. Review: Mr. Tompkins in Wonderland by George Gamow. Mathematical Gazette, 24(258):62–63, February 1940. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328

(electronic). URL http://www.jstor.org/stable/pdfplus/
3607110.pdf.

Meggers:1961:BRG

[Meg61] William F. Meggers. Book review: George Gamow, The Atom and Its Nucleus. Physics Today, 14(10):62–64, October 1961. CO-DEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

URL http://www.physicstoday.org/resource/1/phtoad/v14/
i10/p62_s2.

Meggers:1962:BRG

[Meg62] William F. Meggers. Book review: George Gamow, Biography of Physics. Physics Today, 15(3):78, March 1962. CODEN PHTOAD.

ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://www.

physicstoday.org/resource/1/phtoad/v15/i3/p78_s1.

Mehra:1975:SCP

[Meh75] Jagdish Mehra, editor. The Solvay conferences on physics: aspects of the development of physics since 1911. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1975. ISBN 90-277-0635-2. xxxii + 415 pp. LCCN QC1.S792 M43.

Mark:1969:PMU

[MF69] Hans Mark and Sidney Fernbach, editors. Properties of matter under unusual conditions: in honor of Edward Teller’s 60th birthday. Interscience Publishers, New York, NY, USA, 1969. ISBN 0-470-56990-5. ix + 389 pp. LCCN Q171 .P98.

Mather:1993:C

[MFS93] John C. Mather, D. J. Fixsen, and R. A. Shafer. COBE. Preprint 93-10, NASA Goddard Space Flight Center, Greenbelt,

MD, USA, 1993. URL http://lambda.gsfc.nasa.gov/product/
cobe/bibliography.cfm.

Mishra:2008:QMR

[Mis08] Subodha Mishra. A quantum mechanical relation connecting time, temperature, and cosmological constant of the universe: Gamow’s relation revisited as a special case. International Journal of Theoretical Physics, 47(10):2655–2662, October 2008. CO-DEN IJTPBM. ISSN 0020-7748 (print), 1572-9575 (electronic).

URL http://adsabs.harvard.edu/abs/2008IJTP...47.2655M;
 http://www.springerlink.com/content/u7l40761x0351734/.

Mladjenovic:1998:DYN

[Mla98] Milorad Mladjenović. The Defining Years in Nuclear Physics, 1932–1960s. IOP Publishing, Bristol, UK, 1998. ISBN 0-7503-0472-3 (hardcover). xx + 441 pp. LCCN QC773 .M54 1998.

Meitner:1931:STG

[MP31] Lise Meitner and Kurt Philipp. Das γ-Spektrum von ThC” und die gamowsche Theorie der α-Feinstruktur. (German) [The γ spectrum of ThC” and the Gamow theory of α fine structure]. Naturwissenschaften, 19(50):1007, December 1931. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Melchiorri:1986:GC

[MR86] F. Melchiorri and Remo Ruffini, editors. Gamow cosmology: Proceedings of the International School of Physics “Enrico-Fermi”, Course 86, held at Varenna, Italy, 13–23 July 1982, volume Course 86 of Proceedings of the International School of Physics “Enrico Fermi”. North-Holland Publishing Co., Amsterdam, The Netherlands, 1986. ISBN 0-444-87004-0. LCCN QB980 .I58 1982. URL

http://adsabs.harvard.edu/abs/1986gaco.conf.....M.

Mulders:1941:RBB

[Mul41] Gerard F. W. Mulders. Review: The Birth and Death of the Sun: Stellar Evolution and Subatomic Energy by George Gamow. Publications of the Astronomical Society of the Pacific, 53(311): 56–58, ???? 1941. CODEN PASPAU. ISSN 0004-6280 (print),

1538-3873 (electronic). URL http://www.jstor.org/stable/
pdfplus/40670486.pdf.

Mark:1988:EPW

[MW88] Hans Mark and Lowell Wood, editors. Energy in physics, war and peace: a festschrift celebrating Edward Teller’s 80th birthday. Fundamental theories of physics. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1988. ISBN 90-277-2775-9. vi + 403 pp. LCCN QC71 .E62 1988.

Nadyozhin:1995:GPE

[Nad95] D. K. Nadyozhin. Gamow and the physics and evolution of stars. Space Science Reviews, 74(3–4):455–461, ???? 1995. CODEN SP-SRA4. ISSN 0038-6308 (print), 1572-9672 (electronic). URL http:

//adsabs.harvard.edu/abs/1995SSRv...74..455N.

Nanjundiah:2004:GGG

[Nan04] Vidyanand Nanjundiah. George Gamow and the genetic code. Resonance, 9(7):44–49, July 2004. CODEN RESOFE. ISSN 0971-8044 (print), 0973-712X (electronic). URL http://www.springerlink.

com/content/8r6775x6l6659700/.

Nimtz:2011:TCS

[Nim11] Günter Nimtz. Tunneling confronts Special Relativity. Foundations of Physics, 41(7):1193–1199, July 2011. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://

springerlink3.metapress.com/content/n67286542k5r4186.

Novikov:2007:GGD

[Nov07] I. D. Novikov. George Gamow and the discovery of cosmic microwave background radiation. In Bisnovatyĭ-Kogan et al. [BKST^{+}07], pages 25–?? ISBN 1-904868-38-X. LCCN QB460

.G356. URL http://adsabs.harvard.edu/abs/2007acag.conf.
..25N.

Nugent:1954:RSU

[Nug54] Mark Nugent. Review: Science under the soft cover The Life of the Spider by John Crompton; The Meaning of Evolution by George Gaylord Simpson; Psychopathology of Everyday Life by Sigmund Freud; One Two Three… Infinity by George Gamow; Science and the Modern World by A. N. Whitehead. Chicago Review, 8(4):108–112, ???? 1954. CODEN ???? ISSN 0009-3696 (print), 2327-5804 (electronic). URL http://www.jstor.

org/stable/pdfplus/25293093.pdf.

Nye:2002:CHS

[Nye02] Mary Jo Nye, editor. The Cambridge history of science: The modern physical and mathematical sciences, volume 5. Cambridge University Press, Cambridge, UK, 2002. ISBN 0-521-57244-4 (vol. 3 : hardback), 0-521-57243-6 (vol. 4), 0-521-57199-5 (vol. 5), 0-521-59442-1 (vol. 7). xxvii + 678 pp. LCCN Q125 .C32

2003. URL http://www.h-net.org/review/hrev-a0d3l3-aa;
 http://www.h-net.org/review/hrev-a0d7k4-aa;   http://
www.h-net.org/review/hrev-a0f4e9-aa;   http://www.loc.
gov/catdir/description/cam021/2001025311.html;   http:
//www.loc.gov/catdir/enhancements/fy0731/2001025311-b.
html;   http://www.loc.gov/catdir/samples/cam033/2001025311.
html;   http://www.loc.gov/catdir/toc/cam026/2001025311.
html.

Okun:2002:KPF

[Oku02] L. B. Okun. Key problems in fundamental physics: On the article of George Gamow, D. Ivanenko, and L. Landau, “World Constants and Limiting Transition”. Physics of Atomic Nuclei, 65 (7):1370–1372, ???? 2002. CODEN PANUEO. ISSN 1063-7788 (print), 1562-692X (electronic). URL http://www.springerlink.

com/content/f27102121k72t261/. See [GIL28, GIL02b].

ORaifeartaigh:2018:ILE

[OM18] Cormac O’Raifeartaigh and Simon Mitton. Interrogating the legend of Einstein’s ”biggest blunder”. Physics in Perspective (PIP), 20(4):318–341, December 2018. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Opik:1969:GGD

[Öpi69] E. Öpik. G. Gamow, died 1968 August 19. Irish Astronomical Journal, 9:170–??, 1969. CODEN IRAJAW. ISSN 0021-1052. URL

http://adsabs.harvard.edu/abs/1969IrAJ....9R.170O.

Oppenheimer:1928:QTA

[Opp28] J. R. Oppenheimer. On the quantum theory of the autoelectric field currents. Proceedings of the National Academy of Sciences of the United States of America, 14(5):363–365, May 15, 1928. CODEN PNASA6. ISSN 0027-8424 (print), 1091-6490 (electronic). URL

http://www.jstor.org/stable/pdfplus/85208.pdf.

Pauli:1932:BRG

[Pau32] Wolfgang Pauli. Book review: G. Gamow, Der Bau des Atomkernes und die Radioaktivität. Naturwissenschaften, 20(??):582, ???? 1932. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Parijskij:2000:GMC

[PD00] Y. N. Parijskij and V. K. Dubrovich. Gamow memorial conference St. Petersburg session. Astronomical and Astrophysical Transactions, 19:201–202, 2000. CODEN AATREG. ISSN 1055-6796 (print), 1476-3540 (electronic). URL http://adsabs.harvard.

edu/abs/2000A%26AT...19..201P.

Pedder:2012:MTP

[Ped12] Chris Pedder. Mr Tompkins in Paperback, by George Gamow, Scope: general interest. Level: general readership, non-specialists, pre-university, undergraduate. Contemporary Physics, 53(6):508– 509, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Peebles:1971:PC

[Pee71] P. J. E. (Phillip James Edwin) Peebles. Physical Cosmology. Princeton series in physics. Princeton University Press, Princeton, NJ, USA, 1971. ISBN 0-691-08108-5. xvi + 282 pp. LCCN QB981 .P42.

Peebles:1993:PPC

[Pee93] P. J. E. (Phillip James Edwin) Peebles. Principles of Physical Cosmology. Princeton series in physics. Princeton University Press, Princeton, NJ, USA, 1993. ISBN 0-691-07428-3, 0-691-01933-9 (paperback). xviii + 718 pp. LCCN QB981 .P424

1993. US$79.50US$29.95. URL http://www.loc.gov/catdir/
description/prin021/92033370.html.

Penzias:1972:CMA

[Pen72] Arno A. Penzias. Cosmology and microwave astronomy. In Reines [Rei72a], pages 29–47. ISBN 0-87081-025-1. LCCN QC780 .C65.

URL http://adsabs.harvard.edu/abs/1972cht..conf...29P.

Penzias:1979:OE

[Pen79] Arno A. Penzias. The origin of the elements. Reviews of Modern Physics, 51(3):425–431, July 1979. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.51.425;
 http://rmp.aps.org/abstract/RMP/v51/i3/p425_1.

Perlick:2003:BRB

[Per03] Volker Perlick. Book review: The New World of Mr. Tompkins. By George Gamow and Russell Stannard. Cambridge University Press, 270 p. ISBN 0-521-63009-6, GBP16.95 (hardback 1999), ISBN 0-521-63992-1, GBP10.95 (paperback 2001). General Relativity and Gravitation, 35(1):143–145, January 2003. CO-DEN GRGVA8. ISSN 0001-7701 (print), 1572-9532 (electronic).

URL http://adsabs.harvard.edu/abs/2003GReGr..35..143P;
 http://www.springerlink.com/content/x82j5p752442k6l3/.

Prokofiew:1927:ADL

[PG27] W. Prokofiew and George Gamow. Anomale Dispersion an den Linien der Hauptserie des Kaliums (Verhältnis der Dispersionskonstanten des roten und violetten Dubletts). (German) [Anomalous dispersion of the lines of the principal series of potassium (the ratio of the dispersion constants of the red and violet doublets)]. Zeitschrift für Physik, 44(11–12):887–892, November 1927. CODEN ZEPYAA. ISSN 0044-3328. URL http://www.

springerlink.com/content/r1932n721m2mv828/.

Plavec:1966:RBE

[PG66] M. Plavec and J. Grygar. Review: Matter, Earth, and Sky by George Gamow. Publications of the Astronomical Society of the Pacific, 78(462):266–267, June 1966. CODEN PASPAU. ISSN 0004-6280 (print), 1538-3873 (electronic). URL http://www.jstor.

org/stable/pdfplus/40674357.pdf.

Pohl:1962:ED

[Poh62] Frederik Pohl. The expert dreamers. Doubleday, Garden City, NY, USA, 1962. 248 pp. LCCN PZ1 .P745 Ex. Reprinted in 1968 by Avon Books, New York.

Polya:1958:RBM

[Pol58] G. Polya. Review: Puzzle-Math by George Gamow, Marvin Stern. Science, 128(3317):195–??, July 25, 1958. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.

jstor.org/stable/pdfplus/1754692.pdf.

Polikarov:1972:LCG

[Pol72] A. Polikarov. L’hypothèse cosmologique de Gamov est-elle confirmée?. (French) [Is Gamow’s cosmological hypothesis confirmed?]. Izvestiya na Sektsiyata po Astronomiya, Bulgarska Akademiya na Naukite, 5:89–95, 1972. CODEN IBASBG. ISSN

0525-0897. URL http://adsabs.harvard.edu/abs/1972IzSAB.
..5...89P.

Pomerantz:1944:RAN

[Pom44] Jacob Pomerantz. Review: Atomic nightmares: Mr Tompkins Explores the Atom by George Gamow. The Scientific Monthly, 59 (3):239–240, September 1944. CODEN SCMOAA. ISSN 0096-3771 (print), 2327-7513 (electronic). URL http://www.jstor.

org/stable/pdfplus/18683.pdf.

Prasad:1993:RRR

[Pra93] Rajendra Prasad. Review: Re-reading classics A Mathematician’s Apology by G. H. Hardy; What Little I Remember by Otto Frisch; What Is Life?, with Mind and Matter and Autobiographical Sketches by Erwin Schrödinger; Mr. Tompkins in Paperback by George Gamow. Social Scientist, 21(3/4):98–101, April 1993. CODEN ???? ISSN 0970-0293. URL http://www.jstor.org/

stable/10.2307/3517634;   http://www.jstor.org/stable/
pdfplus/3517634.pdf.

Pustylnik:1996:GGU

[Pus96] I. Pustylnik. George Gamow’s unique style. Astronomical and Astrophysical Transactions, 10:167–173, 1996. CODEN AATREG.

ISSN 1055-6796 (print), 1476-3540 (electronic). URL http://

adsabs.harvard.edu/abs/1996A%26AT...10..167P.

Pustylnik:2007:RGG

[Pus07] I. B. Pustylnik. Reflections on George Gamow’s unique scientific legacy and personality. In Bisnovatyĭ-Kogan et al. [BKST^{+}07], pages 9–?? ISBN 1-904868-38-X. LCCN QB460 .G356. URL

http://adsabs.harvard.edu/abs/2007acag.conf....9P.

R:1953:RBT

[R.53] A. R. Review: Mr. Tompkins Learns the Facts of Life by George Gamow. Bios, 24(3):157–??, October 1953. CODEN ???? ISSN

???? URL http://www.jstor.org/stable/pdfplus/4605559.
pdf.

Rutherford:1929:DSA

[RAC^{+}29] Ernest Rutherford, F. W. Aston, James Chadwick, C. D. Ellis, George Gamow, Ralph Howard Fowler, O. W. Richardson, and Douglas R. Hartree. Discussion on the structure of atomic nuclei. Proceedings of the Royal Society of London. Series A, Containing Papers of a Mathematical and Physical Character, 123(792):373– 390, April 6, 1929. ISSN 0950-1207 (print), 2053-9150 (electronic).

URL http://www.jstor.org/stable/pdfplus/95202.pdf.

Racah:1935:BAH

[Rac35] Giulio Racah. Bemerkung zur Arbeit von Herrn Gamow: Empirische Stabilitätsgrenzen von Atomkernen. (German) [Note on the work of Mr. Gamow: Empirical stability limits of atomic nuclei]. Zeitschrift für Physik, 93(9–10):704, September 1935. CODEN ZEPYAA. ISSN 0044-3328. URL http://www.

springerlink.com/content/xv423381842u1571/. See [Gam34c].

Raychaudhury:2004:GSL

[Ray04] Somak Raychaudhury. And Gamow said, let there be a hot universe. Resonance, 9(7):32–43, July 2004. CODEN RESOFE. ISSN 0971-8044 (print), 0973-712X (electronic). URL http://

www.springerlink.com/content/q420320l5q4l4147/.

Raychaudhury:2005:GSL

[Ray05] Somak Raychaudhury. And Gamow said, let there be a hot universe. Resonance, 10(12):220–231, December 2005. CODEN RE-SOFE. ISSN 0971-8044 (print), 0973-712X (electronic). URL

http://www.springerlink.com/content/w68448111t66883q/.

Rutherford:1930:RRS

[RCE30] Sir Ernest Rutherford, James Chadwick, and C. D. (Charles Drummond) Ellis. Radiations from radioactive substances. Cambridge University Press, Cambridge, UK, 1930. xi + 1 + 588 pp. LCCN QC721 .R94.

Rutherford:1951:RRS

[RCE51] Sir Ernest Rutherford, Sir James Chadwick, and C. D. Ellis. Radiations from radioactive substances. Cambridge University Press, Cambridge, UK, 1951. xi + 588 pp. LCCN QC721 .R94 1951.

Reines:1972:CFOa

[Rei72a] Frederick Reines, editor. Cosmology, fusion and other matters: George Gamow memorial volume. Colorado Associated University Press, Boulder, CO, USA, 1972. ISBN 0-87081-025-1. xiv + 320 pp. LCCN QC780 .C65.

Reines:1972:CFOb

[Rei72b] Frederick Reines, editor. Cosmology, fusion and other matters: George Gamow memorial volume. Adam Hilger Ltd., Bristol, UK, 1972. ISBN 0-85274-223-1. xiv + 320 pp. LCCN QC780.

URL http://adsabs.harvard.edu/abs/1972cht..conf.....R;
 http://adsabs.harvard.edu/abs/1972cofu.book.....R.

Richtmyer:1971:RRB

[Ric71] R. D. Richtmyer. Review: Reminiscences: My World Line. An Informal Autobiography by George Gamow. Science, 171(3975): 997–998, March 12, 1971. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.jstor.org/

stable/pdfplus/1731302.pdf.

Rich:1997:GGC

[Ric97] Alexander Rich. Gamow and the genetic code. In Harper et al. [HPA97b], pages 114–122. ISBN 1-886733-49-X. LCCN QC16.G37

G46 1996. URL http://adsabs.harvard.edu/abs/1997ASPC..
129..114R.

Rigden:1995:JRO

[Rig95] John S. J. Rigden. J. Robert Oppenheimer: before the War. Scientific American, 273(1):76–81, July 1995. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (elec-

tronic). URL http://www.nature.com/scientificamerican/

journal/v273/n1/pdf/scientificamerican0795-76.pdf. Al-

though Oppenheimer is now best remembered for his influence during World War II, he made many important contributions to theoretical physics in the 1930s.

Righetti:2006:ABG

[Rig06] Pier Giorgio Righetti. The Alpher, Bethe, Gamow of isoelectric focusing, the alpha-Centaury of electrokinetic methodologies. Part I. Electrophoresis, 27(5–6):923–938, March 5–6, 2006. CODEN ELCTDN. ISSN 0173-0835 (print), 1522-2683 (electronic).

Righetti:2007:ABG

[Rig07] Pier Giorgio Righetti. The Alpher, Bethe and Gamow of IEF, the alpha-Centaury of electrokinetic methodologies. Part II: Immobilized pH gradients. Electrophoresis, 28(4):545–555, February 4, 2007. CODEN ELCTDN. ISSN 0173-0835 (print), 1522-2683 (electronic).

Rindler:2009:GEMa

[Rin09] Wolfgang Rindler. Gödel, Einstein, Mach, Gamow, and Lanczos: Gödel’s remarkable excursion into cosmology. American Journal of Physics, 77(6):498–510, June 2009. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). URL http:/

/adsabs.harvard.edu/abs/2009AmJPh..77..498R;   http://
ajp.aapt.org/resource/1/ajpias/v77/i6/p498_s1. Reprinted

in [BPP^{+}11, Chapter 9].

Rindler:2011:GEMb

[Rin11] Wolfgang Rindler. Gödel, Einstein, Mach, Gamow, and Lanczos: Gödel’s remarkable excursion into cosmology. In Baaz et al. [BPP^{+}11], pages 185–212. ISBN 0-521-76144-1 (print), 0-511-97423-X (e-book). LCCN QA9.65 .K87 2011.

Rogers:1962:RRS

[Rog62] Eric M. Rogers. Review: Responsibility and science writing: Spacecraft by James L. Haggerty, Jr.; Gravity by George Gamow; Artificial Satellites by Michael W. Ovenden. Science, 136(3515):499– 501, May 11, 1962. CODEN SCIEAS. ISSN 0036-8075 (print),

1095-9203 (electronic). URL http://www.jstor.org/stable/
pdfplus/1708868.pdf.

Rogers:2010:MIS

[Rog10] Kara Rogers, editor. The 100 Most Influential Scientists of All Time. The Britannica guide to the world’s most influential people. Britannica Educational Publishers, in association with Rosen Educational Services, New York, NY, USA, 2010. ISBN 1-61530-002-3 (library binding). 360 pp. LCCN Q162 .A15 2010.

Rosenfeld:1972:NR

[Ros72] Léon Rosenfeld. Nuclear reminiscences. In Reines [Rei72a], pages 289–299. ISBN 0-87081-025-1. LCCN QC780 .C65. URL http://

adsabs.harvard.edu/abs/1972cht..conf..289R.

Ranyuk:2007:GGN

[RSJ07] Y. Ranyuk, O. Shevchenko, and P. Josephson. George Gamow and nuclear physics in Ukraine. In Bisnovatyĭ-Kogan et al. [BKST^{+}07], pages 71–?? ISBN 1-904868-38-X. LCCN QB460 .G356. URL

http://adsabs.harvard.edu/abs/2007acag.conf...71R.

Rubin:1997:WGG

[Rub97] V. Rubin. What George Gamow did not know about the Universe. In Harper et al. [HPA97b], pages 95–?? ISBN 1-886733-49-X.

LCCN QC16.G37 G46 1996. URL http://adsabs.harvard.edu/
abs/1997ASPC..129...95R.

Rubin:2002:IIM

[Rub02] Vera C. Rubin. Intuition and inspiration made Gamow a star turn. Nature, 415(6867):13, January 3, 2002. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.

nature.com/nature/journal/v415/n6867/full/415013a.html.

Rutherford:1927:LSR

[Rut27] Sir Ernest Rutherford, O.M., P.R.S. LI. Structure of the radioactive atom and origin of the α-rays. Philosophical Magazine (7), 4(22):580–605, September 1927. CODEN PHMAA4. ISSN 1941-5982 (print), 1941-5990 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/14786440908564361. Cited

in [Wil83, page 441] as ‘a great paper’. Wilson (page 559) later notes that this paper inspired George Gamow to his prediction of the quantum tunneling effect in 1929 (credit also goes to Edward Condon and Ronald Gurney who wrote two papers in 1928 on that idea, and to Robert Oppenheimer, who published a paper on that topic five months before those of Condon and Gurney).

Rapport:1964:P

[RW64] Samuel Rapport and Helen Wright, editors. Physics. New York University Press, New York City, NY, USA, 1964. xiii + 333 pp.

Ryabov:2005:GYO

[Rya05] M. I. Ryabov. Gamow’s year in Odessa. Odessa Astronomical Publications = Izvestija Odesskoj Astronomičeskoj Observatorii, 17:5–7, 2005. CODEN ???? ISSN 1810-4215. URL http://

adsabs.harvard.edu/abs/2005OAP....17....5R.

Ryabov:2006:GSS

[Rya06] M. I. Ryabov. The 6th Gamow Summer School: Astronomy and Beyond: Astrophysics, Cosmology, Radioastronomy, Astrobiology. Astronomical and Astrophysical Transactions, 25:359–361, October 2006. CODEN AATREG. ISSN 1055-6796 (print), 1476-3540 (elec-

tronic). URL http://adsabs.harvard.edu/abs/2006A%26AT...
25..359R.

Sabadell:1996:GGS

[Sab96] M. A. Sabadell. George Gamow, science writer. A consideration of what science is and how it is done. Astronomical and Astrophysical Transactions, 10:9–19, 1996. CODEN AATREG. ISSN 1055-6796 (print), 1476-3540 (electronic). URL http://adsabs.harvard.

edu/abs/1996A%26AT...10....9S.

Salpeter:1952:NRS

[Sal52] Edwin E. Salpeter. Nuclear reactions in stars without hydrogen. Astrophysical Journal, 115(??):326–328, ???? 1952. CO-DEN ASJOAB. ISSN 0004-637X (print), 1538-4357 (electronic).

URL http://adsabs.harvard.edu/abs/1952ApJ...115..326S.

Reprinted in the centennial edition Ap. J. 525C, 450 (1999) with a modern commentary by André Maeder; also excerpted, with commentary, in Lang, Kenneth R. & Owen Gingerich, eds., A Source Book in Astronomy & Astrophysics, 1900–1975 (Harvard Univ. Press, Cambridge, MA, 1979), 349–352.

Salpeter:1996:RGG

[Sal96] Edwin Salpeter. Reminiscences of George Gamow and nuclear astrophysics. Astronomical and Astrophysical Transactions, 10(??): 3–37, ???? 1996. CODEN AATREG. ISSN 1055-6796 (print),

1476-3540 (electronic). URL http://adsabs.harvard.edu/abs/
1996A%26AT...10....3S.

Sneden:2008:NCE

[SCG08] Christopher Sneden, John J. Cowan, and Roberto Gallino. Neutron-capture elements in the early galaxy. Annual Review of Astronomy and Astrophysics, 46(??):241–288, September 2008. CODEN ARAAAJ. ISSN 0066-4146 (print), 1545-

4282 (electronic). URL http://www.annualreviews.org/doi/
abs/10.1146/annurev.astro.46.060407.145207.

Schatzman:1969:GG

[Sch69] E. Schatzman. G. Gamov, 1904–1968. Acta Astronautica, 14: 690–??, 1969. CODEN AASTCF. ISSN 0094-5765 (print),

1879-2030 (electronic). URL http://adsabs.harvard.edu/abs/
1969AcAau..14..690S.

Schils:2012:GG

[Sch12a] René Schils. George Gamow. In How James Watt invented the copier: forgotten inventions of our great scientists [Sch12b], pages 159–165. ISBN 1-4614-0859-8 (softcover), 1-4614-0860-1 (e-book).

LCCN T15 .S35513 2012. URL http://www.springerlink.com/
content/jv20604l06332535/.

Schils:2012:HJW

[Sch12b] René Schils. How James Watt invented the copier: forgotten inventions of our great scientists. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2012. ISBN 1-4614-0859-8 (softcover), 1-4614-0860-1 (e-book). vii + 170 pp. LCCN T15 .S35513 2012.

Scott:2007:GGG

[Sco07] Douglas Scott. Gamow, George [Georgiy] (Antonovich). In Hockey et al. [HTB^{+}07], pages 403–404 (part 7). ISBN 0-387-31022-3 (set), 0-387-30400-2 (set: electronic version), 0-387-33628-1 (set: print + electronic bundle). LCCN QB35 .B565 2007. URL http:

//www.loc.gov/catdir/enhancements/fy0814/2008270178-d.
html;   http://www.loc.gov/catdir/enhancements/fy0824/
2008270178-t.html.

Segre:2011:OGM

[Seg11] Gino Segrè. Ordinary geniuses: Max Delbrück, George Gamow, and the origins of genomics and Big Bang cosmology. Viking Press, New York, NY, USA, 2011. ISBN 0-670-02276-4. xxi + 330 pp. LCCN QH31.D434 S44 2011.

Stannard:2012:NMM

[SG12] Russell Stannard and George Gamow. Le nouveau monde de M. Tompkins. (French) [The new world of Mr. Tompkins]. Plumes de science Le Pommier! Le Pommier, Paris, France, 2012. ISBN 2-7465-0610-6 (paperback). 351 pp. LCCN QC24.5 .S73 2012; QC71 .S7314 2012.

Shapley:1953:RTM

[Sha53] Harlow Shapley. Review: Touring the moon: The Moon by George Gamow, A Guide to the Moon by Patrick Moore. The Scientific Monthly, 77(3):166, September 1953. CODEN SCMOAA. ISSN 0096-3771 (print), 2327-7513 (electronic).

URL http://adsabs.harvard.edu/abs/1953SciMo..77..166G;
 http://www.jstor.org/stable/pdfplus/20903.pdf.

Shapiro:1972:GGA

[Sha72] Maurice M. Shapiro. George Gamow — an appreciation. In Reines [Rei72a], pages 300–303. ISBN 0-87081-025-1. LCCN QC780 .C65.

URL http://adsabs.harvard.edu/abs/1972cht..conf..300S.

Shapiro:2007:GGG

[Sha07] M. M. Shapiro. George Gamow — a giant in 20th century science. In Bisnovatyĭ-Kogan et al. [BKST^{+}07], pages 5–?? ISBN 1-904868-

38-X. LCCN QB460 .G356. URL http://adsabs.harvard.edu/
abs/2007acag.conf....5S.

Singh:2004:BBO

[Sin04] Simon Singh. Big Bang: the origin of the universe. Fourth Estate, New York, NY, USA, 2004. ISBN 0-00-716220-0, 0-00-715251-5 (hardcover). 532 pp. LCCN QB991.B54 S56

2004.  URL http://www.simonsingh.com/;   http://www.
simonsingh.com/Big_Bang.html;   http://www.simonsingh.
com/Big_Bang_Reviews.html.

Sitterly:1964:SLB

[Sit64a] B. W. Sitterly. Science for the layman. (book reviews: A Star Called the Sun). Science, 145:476–477, July 1964. CODEN SCI-EAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http:

//adsabs.harvard.edu/abs/1964Sci...145..476S.

Sitterly:1964:BRB

[Sit64b] Bancroft W. Sitterly. Book review: A Star Called the Sun. George Gamow. Viking, New York, 1964. xiv + 208 pp. Illus. $5.75.

Science, 145(3631):476–477, July 31, 1964. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.

sciencemag.org/content/145/3631/476.3.extract.

Sitterly:1964:RSL

[Sit64c] Bancroft W. Sitterly. Review: Science for the layman: A Star Called the Sun by George Gamow. Science, 145(3631):476–477, July 31, 1964. CODEN SCIEAS. ISSN 0036-8075 (print),

1095-9203 (electronic). URL http://www.jstor.org/stable/
pdfplus/1714573.pdf.

Skewes:1954:RBH

[Ske54] George J. Skewes. Review: Man and His World Series: Your Blood and You by Nathan S. Washton, Sarah R. Riedman; Man and His World Series: The Way of Science: Its Growth and Method by Nathan S. Washton, John Somerville; Man and His World Series: The Moon by Nathan S. Washton, George Gamow; Man and His World Series: Life on the Earth by Nathan S. Washton, Rose Wyler, Gerald Ames. The School Review (Ithaca, NY), 62(1):58–60, January 1954. CODEN SCLRAE. ISSN 0036-6773 (print), 1940-1213 (electronic). URL http://www.jstor.org/

stable/pdfplus/1082161.pdf.

Smith:1961:BRB

[Smi61a] Malcolm K. Smith. Book review: The Atom and Its Nucleus. George Gamow. Prentice-Hall, New York, 1961, 153 pp. $1.95. Science, 134(3474):276, July 28, 1961. CO-DEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

URL http://www.sciencemag.org/content/134/3474/276.1.
citation?sid=199506b7-d99f-4510-8682-299948178e02.

Smith:1961:RMP

[Smi61b] Malcolm K. Smith. Review: Modern physics for laymen: The Atom and Its Nucleus by George Gamow; Atomic Physics Today by Otto R. Frisch. Science, 134(3474):276, July 28, 1961. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL

http://www.jstor.org/stable/pdfplus/1708183.pdf.

Shapiro:1972:DCR

[SST72] Maurice M. Shapiro, Rhein Silberberg, and Chen-Hsiang Tsao. Diffusion of cosmic rays and their source composition. In Reines [Rei72a], pages 124–149. ISBN 0-87081-025-1. LCCN QC780 .C65.

URL http://adsabs.harvard.edu/abs/1972cht..conf..124S.

Stannard:1999:NWM

[Sta99] Russell Stannard. The New World of Mr. Tompkins: George Gamow’s Classic Mr. Tompkins in Paperback. Cambridge University Press, Cambridge, UK, 1999. ISBN 0-521-63009-6 (hardcover), 0-521-63992-1 (paperback), 0-511-00800-7 (e-book), 0-511-58169-6 (e-book). xii + 258 pp. LCCN QC71.S775 1999. US$24.99. URL

http://www.loc.gov/catdir/description/cam029/98050379.
html;   http://www.loc.gov/catdir/enhancements/fy0732/
98050379-b.html;   http://www.loc.gov/catdir/toc/cam021/
98050379.html.

Stuewer:1971:BRBb

[Stu71] Roger H. Stuewer. Book review: My World Line: An Informal Autobiography by George Gamow. Isis, 62(2):266–267, Summer 1971. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/229269;   http:
//www.jstor.org/stable/pdfplus/229269.pdf.

Stuewer:1986:GTA

[Stu86] Roger H. Stuewer. Gamow’s theory of alpha-decay. In Ullmann-Margalit [UM86a], chapter 14, pages 147–186. ISBN 90-277-2159-9 (paperback), 90-277-2158-0 (hardcover), 94-009-5496-4 (e-book). ISSN 0068-0346. LCCN Q174 .B67 vol. 94

Q175. URL https://link.springer.com/chapter/10.1007/
978-94-009-5496-0_14. See comment [Lip86].

Stuewer:1994:OLD

[Stu94] Roger H. Stuewer. The origin of the liquid-drop model and the interpretation of nuclear fission. Perspectives on Science, 2(1):76– 129, Spring 1994. CODEN PRSIEU. ISSN 1063-6145 (print), 1530-

9274 (electronic). URL http://alsos.wlu.edu/information.

aspx?id=1736. See errata [Ano94].

Stuewer:1997:GAD

[Stu97] R. H. Stuewer. Gamow, alpha decay, and the liquid-drop model of the nucleus. In Harper et al. [HPA97b], pages 29–43. ISBN 1-886733-49-X. LCCN QC16.G37 G46 1996. URL http://adsabs.

harvard.edu/abs/1997ASPC..129...29S.

Stuewer:2013:ACM

[Stu13] Roger H. Stuewer. An act of creation: the Meitner–Frisch interpretation of nuclear fission. In Katzir et al. [KLR13], pages 231–245. ISBN 3-8442-5134-0. LCCN QC173.98. URL http://

www.edition-open-access.de/proceedings/5/.

Stuewer:2018:AIN

[Stu18] Roger H. Stuewer. The Age of Innocence: Nuclear Physics Between the First and Second World Wars. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2018. ISBN 0-19-186658-X, 0-19-882787-3 (hardback), 0-19-256290-8 (e-book). xv + 484 pp. LCCN QC773 .S78 2018.

Susman:1969:RBT

[Sus69] Millard Susman. Review: Mr. Tompkins Inside Himself. Adventures in the New Biology. by George Gamow, Martynas Yčas. Quarterly Review of Biology, 44(1):71–??, March 1969. CODEN QRBIAK. ISSN 0033-5770 (print), 1539-7718 (electronic). URL

http://www.jstor.org/stable/pdfplus/2818613.pdf.

Teller:1972:CC

[Tel72] Edwin Teller. Are the constants constant? In Reines [Rei72a], pages 60–66. ISBN 0-87081-025-1. LCCN QC780 .C65.

Teller:1997:SPM

[Tel97] Edward Teller. Some personal memories of George Gamow. In Harper et al. [HPA97b], pages 124–126. ISBN 1-886733-49-X.

LCCN QC16.G37 G46 1996. URL http://adsabs.harvard.edu/
abs/1997ASPC..129..123T.

Teller:1939:FWC

[TGF39] Edward Teller, George Gamow, and J. A. Fleming. The Fifth Washington Conference on Theoretical Physics. Science, 89(2304): 180–182, February 24, 1939. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

Teller:1941:SAW

[TGF41] Edward Teller, George Gamow, and J. A. Fleming. The Seventh Annual Washington Conference of Theoretical Physics, May 22–24, 1941. Science, 94(2430):92–94, July 25, 1941. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.

jstor.org/stable/pdfplus/1669027.pdf.

Editors:2001:CBG

[The01] The Editors. Clarification: Genes, Girls and Gamow. Nature, 414(6863):487, November 29, 2001. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.

nature.com/nature/journal/v414/n6863/full/414487b.html.

See [Jud01].

Trimble:2010:OAC

[Tri10] Virginia Trimble. The origins and abundances of the chemical elements before 1957: from Prout’s hypothesis to Pasadena. European Physical Journal H, 35(1):89–109, July 2010. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic). URL http://

link.springer.com/article/10.1140/epjh/e2010-00006-9.

Terlevich:2007:GLP

[TTL07] E. Terlevich, R. Terlevich, and V. Luridiana. Gamow legacy and the primordial abundance of light elements. In Bisnovatyĭ-Kogan et al. [BKST^{+}07], pages 37–?? ISBN 1-904868-38-X. LCCN QB460

.G356. URL http://adsabs.harvard.edu/abs/2007acag.conf.
..37T.

Tuck:1972:WER

[Tuc72] James L. Tuck. World energy reserves and some speculations on the future of nuclear fusion energy. In Reines [Rei72a], pages 192–231. ISBN 0-87081-025-1. LCCN QC780 .C65. URL http://adsabs.

harvard.edu/abs/1972cht..conf..192T.

Turner:2008:PCA

[Tur08] Michael S. Turner. From αβγ to precision cosmology: The amazing legacy of a wrong paper. Physics Today, 61(12):8–9, December 2008. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699

(electronic). URL http://www.physicstoday.org/resource/1/

PHTOAD/v61/i12. This is a retrospective of influence the famous paper [ABG48] on the origin of the chemical elements, and the Big Bang theory of the evolution of the Universe. See also comments [AWCT09]. The correct theory of the origin of the elements appears in [BBFH57].

Ulam:1972:GM

[Ula72] S. M. Ulam. Gamow — and mathematics. In Reines [Rei72a], pages 272–279. ISBN 0-87081-025-1. LCCN QC780 .C65. URL

http://adsabs.harvard.edu/abs/1972cht..conf..272U.

Ullmann-Margalit:1986:KSI

[UM86a] Edna Ullmann-Margalit, editor. The Kaleidoscope of Science: The Israel Colloquium: Studies in History, Philosophy, and Sociology of Science. Volume 1, volume 94 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1986. ISBN 90-277-2159-9 (paperback), 90-277-2158-0 (hardcover), 94-009-5496-4 (e-book). ISSN 0068-0346. xii + 250 pp. LCCN Q174 .B67 vol. 94 Q175.

URL https://link.springer.com/book/10.1007/978-94-009-
5496-0.

Ullmann-Margalit:1986:PSI

[UM86b] Edna Ullmann-Margalit, editor. The Prism of Science: The Israel Colloquium: Studies in History, Philosophy, and Sociology of Science. Volume 2, volume 95(2) of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1986. ISBN 90-277-2160-2, 90-277-2161-0 (paperback), 94-009-4566-3 (e-book). ISSN 0068-0346. ix + 250 pp. LCCN Q174 .B67 vol. 95 Q175. URL http://www.

springerlink.com/content/978-94-009-4566-1;   https://
link.springer.com/book/10.1007/978-94-009-4566-1.

Unsold:1960:BGA

[Uns60] A. Unsöld. Buchbesprechung: George Gamow, Die Geburt des Alls (Übersetz von W. Petri). H. Reich Verlag, München, 158 s. Geb. DM 9.80. (German) [Book review: George Gamow, The Birth of Everything [The Creation of the Universe]]. Zeitschrift f”ur Astrophysik, 49:229, 1960. CODEN ZEASAJ. ISSN 0372-8331. URL

http://adsabs.harvard.edu/abs/1960ZA.....49..229G.

Ulam:1986:GMP

[URR86a] Stanisław M. Ulam, Mark C. Reynolds, and Gian-Carlo Rota. Gamow and mathematics: Personal reminiscences. In Science, computers, and people: from the tree of mathematics [URR86b], pages 231–240. ISBN 0-8176-3276-X. LCCN QA7 .U431 1986; QA7 .U43 1986. From the tree of mathematics, With a preface by Martin Gardner, With an introduction by Françoise Ulam.

Ulam:1986:SCP

[URR86b] Stanisław M. Ulam, Mark C. Reynolds, and Gian-Carlo Rota. Science, computers, and people: from the tree of mathematics. Birkhäuser Boston Inc., Cambridge, MA, USA, 1986. ISBN 0-8176-3276-X. xxii + 264 pp. LCCN QA7 .U431 1986; QA7 .U43 1986. From the tree of mathematics, With a preface by Martin Gardner, With an introduction by Françoise Ulam.

VanAmringe:1953:RBM

[Van53] Edwin V. Van Amringe. Review: The Moon by George Gamow. The Clearing House, 28(3):185–??, November 1953. CODEN ???? ISSN 0009-8655 (print), 1939-912X (electronic). URL http://

www.jstor.org/stable/pdfplus/30176228.pdf.

VanName:1962:BRG

[Van62] F. W. Van Name, Jr. Book review: George Gamow, Biography of Physics. American Journal of Physics, 30(2):147–148, February 1962. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (elec-

tronic). URL http://ajp.aapt.org/resource/1/ajpias/v30/
i2/p147_s2.

vandenBerg:2012:GVA

[vdBS12] Robert Erwin van den Berg and Ansfried Scheifes. Gamow. Van atoomkern tot kosmos. (Dutch) [Gamow, from the atomic nucleus to the cosmos], volume 42 of Wetenschappelijke biografie. Natuurwetenschap and Techniek, Amsterdam, The Netherlands, 2012. ISBN 90-8571-378-1. 160 pp. LCCN Q124.6-127.2.

vonNeumann:1996:PJN

[vN96] John von Neumann. Papers of John von Neumann, 1912–1996 (bulk 1935–1957). US Library of Congress archival manuscript material (collection)., 1996. 11,660 items. 34 containers plus 1 vault container. 13.4 linear feet. Manuscript number MSS44180. Correspondence, memoranda, journals, speeches, article and book drafts, notes, charts, graphs, patent, biographical material, family papers, printed materials, newspaper clippings, photographs, and other materials pertaining primarily to von Neumann’s career as professor of mathematics at the Institute for Advanced Study including his directorship of the Electronic Computer Project; adviser and commissioner on the U.S. Atomic Energy Commission; scientific consultant to government and private concerns, including the Los Alamos Scientific Laboratory, Los Alamos, New Mexico, and the U.S. Army Ballistic Research Laboratory, Aberdeen, Maryland; and author of works on ballistic research, computers, continuous geometries, logic, operator theory, quantum mechanics, and the theory of games. Includes evaluations of his work written after his death by colleagues including Herman Heine Goldstine, Paul R. Halmos, and Abraham H. Taub. Of special interest are an Albert Einstein letter and report on theoretical physics (1937). Also includes a small amount of material pertaining to Eva and Peter Aldor. Correspondents include Eva Aldor, Frank Aydelotte, Hans Albrecht Bethe, Garrett Birkhoff, S. Chandrasekhar, George Bernard Dantzig, P. A. M. Dirac, Carl Eckart, Enrico Fermi, Abraham Flexner, George Gamow, Kurt Gödel, Herman Heine Goldstine, Werner Heisenberg, L. van Hove, Cuthbert Corwin Hurd, Pascual Jordan, R. H. Kent, George B. Kistiakowsky, Oskar Morgenstern, J. Robert Oppenheimer, Rudolf Ortvay, Wolfgang Pauli, Marshall H. Stone, Lewis L. Strauss, Abraham Haskel Taub, Edward Teller, Stanislaw M. Ulam, Oswald Veblen, Klara Dan Von Neumann, Warren Weaver, Hermann Weyl, Norbert Wiener, and Eugene Paul Wigner. Gift, Marina Von Neumann Whitman, 1974– 1975. Gift, Nicholas A. Vonneuman, 1993.

vonWeizsacker:1935:TKG

[vW35] C. F. von Weizsäcker. Zur Theorie der Kernmassen. (German) [On the theory of nuclear mass]. Zeitschrift für Physik, 96(??):431–458, ???? 1935. CODEN ZEPYAA. ISSN 0044-3328.

Wataghin:1946:ANU

[Wat46] G. Wataghin. On the abundance of nuclei in the universe. Physical Review, 70:430–431, September 1946. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://

link.aps.org/doi/10.1103/PhysRev.70.430.2. See remarks in

[Dys93] about the relation of this work to [ABG48], and the subsequent incorrect neglect of Wataghin’s work. See also related papers [LW46, DW48, Wat48].

Wataghin:1948:FCE

[Wat48] Gleb Wataghin. On the formation of chemical elements inside the stars. Physical Review, 73(1):79, January 1948. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http:/

/link.aps.org/doi/10.1103/PhysRev.73.79. See remarks in

[Dys93] about the relation of this work to [ABG48], and the subsequent incorrect neglect of Wataghin’s work. See also related papers [LW46, Wat46, DW48].

Wataghin:1972:MEU

[Wat72] Gleb Wataghin. On a model of the expanding universe. In Reines [Rei72a], pages 48–55. ISBN 0-87081-025-1. LCCN QC780 .C65.

URL http://adsabs.harvard.edu/abs/1972cht..conf...48W.

Watson:2001:GGG

[Wat01] James D. Watson. Genes, girls, and Gamow. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2001. ISBN 0-19-850976-6. xxvii + 275 pp. LCCN QD435.W373 W38 2001.

Watson:2002:GGG

[Wat02] James D. Watson. Genes, Girls, and Gamow: After the Double Helix. Alfred A. Knopf, New York, NY, USA, 2002. ISBN 0-375-41283-2. xxix + 259 pp. LCCN QH435.W373

W38 2002. URL http://www.loc.gov/catdir/bios/random053/
2001038543.html;   http://www.loc.gov/catdir/description/
random042/2001038543.html;   http://www.loc.gov/catdir/
samples/random041/2001038543.html.

Weber:1973:RWS

[Web73] Robert L. Weber, editor. A Random Walk In Science. The Institute of Physics, Bristol and London, 1973. ISBN 0-85498-027-X. xvii + 206 pp. LCCN Q167.W42. Compiled by R. L. Weber, edited by E. Mendoza, with a foreword by William Cooper.

Weiner:1968:IGG

[Wei68] Charles Weiner. Interview with George Gamow at Professor Gamow’s home in Boulder, Colorado April 25, 1968. Niels Bohr Library & Archives, American Institute of Physics, College Park,

MD, USA., April 25, 1968. URL http://www.aip.org/history/
ohilist/4325.html.

Weinberg:1972:GCP

[Wei72a] Steven Weinberg. Gravitation and cosmology: principles and applications of the General Theory of Relativity. Wiley, New York, NY, USA, 1972. ISBN 0-471-92567-5. xxviii + 657 pp. LCCN

QC6 .W47. URL http://www.loc.gov/catdir/toc/onix04/
78037175.html.

Weiner:1972:MNP

[Wei72b] Charles Weiner. 1932 — moving into the new physics. Physics Today, 25(5):40–49, May 1972. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://scitation.aip.

org/content/aip/magazine/physicstoday/article/25/5/10.

1063/1.3070853. Reprinted in [Wei85].

Weinberg:1977:FTM

[Wei77] Steven Weinberg. The First Three Minutes: a Modern View of the Origin of the Universe. Basic Books, New York, NY, USA, 1977. ISBN 0-465-02435-1. x + 188 + 5 pp. LCCN QB981 .W48.

Weiner:1985:MNP

[Wei85] Charles Weiner. 1932 — moving into the new physics. In Weart and Phillips [WP85], pages 332–339. ISBN 0-88318-468-0 (paperback). LCCN QC7 .H694 1985. Reprint of [Wei72b].

Weinberg:1993:FTM

[Wei93] Steven Weinberg. The First Three Minutes: a Modern View of the Origin of the Universe. Basic Books, New York, NY, USA, second edition, 1993. ISBN 0-465-02437-8 (paperback). x + 203 pp.

LCCN QB981 .W48 1993. URL http://www.loc.gov/catdir/
enhancements/fy0830/93232406-b.html;   http://www.loc.
gov/catdir/enhancements/fy0830/93232406-d.html.

Weinberg:2008:C

[Wei08] Steven Weinberg. Cosmology. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2008. ISBN 0-19-852682-2 (hardcover). xvii + 593 pp. LCCN QB981 .W475 2008.

Weinstein:2013:GGA

[Wei13] Galina Weinstein. George Gamow and Albert Einstein: Did Einstein say the cosmological constant was the “biggest blunder” he ever made in his life? arxiv.org, ??(??):??, October 3, 2013. URL

http://arxiv.org/abs/1310.1033.

Whitfield:2007:WMR

[WH07] Donald Whitfield and James L. Hicks, editors. What’s the matter?: readings in physics. Great Books Foundation, Chicago, IL, USA, 2007. ISBN 1-880323-91-5. xx + 540 pp. LCCN QC7.5 .W53 2007.

URL http://store.greatbooks.org/adu-wtm.html. Foreword

by Alan Lightman.

Wilson:1971:BRS

[Wil71] Jane Wilson. Book reviews: Several lives and more: My World Line, by George Gamow, The Viking Press, New York, 1970, 178 pages; Enrico Fermi: Physicist, by Emilio Segrè, The University of Chicago Press, Chicago, 1970, 276 pages; My Several Lives, Memoirs of a Social Inventor, by James B. Conant, Harper & Row, New York, 1970, 701 pages. Bulletin of the Atomic Scientists, 27(2):47–48, February 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wilson:1979:CMB

[Wil79] R. W. Wilson. The cosmic microwave background radiation. Reviews of Modern Physics, 51(3):433–445, July 1979. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.51.433;
 http://rmp.aps.org/abstract/RMP/v51/i3/p433_1.

Wilson:1983:RSG

[Wil83] David Wilson. Rutherford, simple genius. Hodder and Stoughton, London, UK, 1983. ISBN 0-340-23805-4. 638 + 8 pp. LCCN QC16.R8 W54 1983. US$14.95. URL http://catalog.

hathitrust.org/api/volumes/oclc/10560773.html.

Woese:1967:GCM

[Woe67] Carl R. Woese. The genetic code: the molecular basis for genetic expression. Modern perspectives in biology. Harper & Row, New York, NY, USA, 1967. vii + 200 pp. LCCN ????

Weart:1985:HP

[WP85] Spencer R. Weart and Melba Phillips, editors. History of physics, volume 2 of Readings from Physics Today. American Institute of Physics, Woodbury, NY, USA, 1985. ISBN 0-88318-468-0 (paperback). 375 pp. LCCN QC7 .H694 1985.

Yourgrau:1972:EPN

[YvdM72] Wolfgang Yourgrau and Alwyn van der Merwe. Entropy (positive and negative), information and statistical thermodynamics. In Reines [Rei72a], pages 241–271. ISBN 0-87081-025-1. LCCN

QC780 .C65. URL http://adsabs.harvard.edu/abs/1972cht..
conf..241Y.

Zeldovich:1973:SEV

[ZN73] Ya. B. Zel’dovich and I. D. Novikov. Stroenie i Evolyutsiya Vselennoi (Russian) [Structure and Evolution of the Universe]. Nauka, Moscow, USSR, 1973. ISBN ???? ???? pp. LCCN ???? US$.